### v1

# Terminal 3-folds that are not Cohen-Macaulay

Terminal 3-folds that are not Cohen-Macaulay
Burt Totaro
An important ingredient of the minimal model program is that Kawamata log terminal singularities in characteristic zero are rational, and in particular Cohen-Macaulay. In the special case of cone singularities, this fact is related to the Kodaira vanishing theorem restricted to Fano varieties. It turns out that Kodaira vanishing fails even for Fano varieties, in every characteristic $p>0$ [26]. This led to examples of klt, and even terminal, singularities that are not Cohen-Macaulay [26, 28]. (Terminal singularities are the smallest class of singularities that can be allowed on minimal models.)
The most notable example was a terminal singularity of dimension 3 that is not Cohen-Macaulay. Namely, let $X$ be the quotient $(A^{1}-0)^{3}/G$ over the field $\text{\bf F}_{2}$, where the generator $\sigma$ of the group $G=\text{\bf Z}/2$ acts by
$\sigma(x_{1},x_{2},x_{3})=\bigg{(}\frac{1}{x_{1}},\;\frac{1}{x_{2}},\;\frac{1}{x_{3}}\bigg{)}.$
Then $X$ is terminal but not Cohen-Macaulay [26, Theorem 5.1]. This is the lowest possible dimension, because every terminal (or just normal) surface is Cohen-Macaulay. Cohen-Macaulayness and stronger properties such as $F$-regularity help to construct contractions of varieties. Partly for this reason, the MMP for 3-folds is known only in characteristics at least 5 [13]. By Arvidsson–Bernasconi–Lacini, klt singularities in characteristic greater than 5 are Cohen-Macaulay, whereas there are klt singularities that are not Cohen-Macaulay in characteristics 2, 3, and 5 [4, 6, 9].
In this paper, we construct terminal 3-fold singularities that are not Cohen-Macaulay in five new cases: mixed characteristic $(0,2)$, characteristic 3, mixed characteristic $(0,3)$, characteristic 5, and mixed characteristic $(0,5)$ (Theorems 0.1, 6.1, 7.1, 8.1, and 9.1. This is optimal, in view of the result of Arvidsson–Bernasconi–Lacini. Indeed, the MMP for schemes of dimension 3 was developed in mixed characteristic when the residue characteristic is greater than 5 [7, 25]. This raised the question of whether vanishing theorems for 3-folds hold in mixed characteristic. Given our counterexample over $\text{\bf F}_{2}$, one might expect an example of dimension 4, flat over the 2-adic integers $\text{\bf Z}_{2}$, with fiber over $\text{\bf F}_{2}$ being the 3-fold singularity above. In fact, each of our examples has dimension 3 as a scheme. For example, over $\text{\bf Z}_{2}$ we have:
Theorem 0.1.
Let $Y=\{(x,y,i)\in A^{3}_{\text{\bf Z}_{2}}:x\neq 0,y\neq 0,i^{2}=-1\}$. Let the group $G=\text{\bf Z}/2=\{1,\sigma\}$ act on $Y$ by
$\sigma(x,y,i)=(1/x,1/y,-i).$
Then the scheme $Y/G$ is terminal, not Cohen-Macaulay, of dimension 3, and flat over $\text{\bf Z}_{2}$.
Note that an action of a $p$-group on a smooth variety in characteristic $p$ with an isolated fixed point is never formally isomorphic to a linear action, since a nonzero representation $V$ of a $p$-group $G$ in characteristic $p$ has nonzero $G$-fixed subspace. In fact, there are continuous families of non-equivalent actions on smooth varieties in characteristic $p$, and likewise on regular schemes of mixed characteristic. For an action of $G=\text{\bf Z}/p$ on a regular scheme $W$ of dimension 3 with an isolated fixed point in characteristic $p$ (as in Theorem 0.1), it is common for $W/G$ not to be Cohen-Macaulay, essentially because the cohomology of $G$ contributes to the local cohomology of $W/G$. The difficulty is to construct an example with $W/G$ terminal. For a more complicated action of $G$, the quotient scheme would usually not be terminal or even log canonical. To find the examples in this paper, the idea was to look for the simplest possible actions of $G=\text{\bf Z}/p$ on a regular 3-dimensional scheme with an isolated fixed point of residue characteristic $p$.
Our examples build on Artin’s examples of the simplest $\text{\bf Z}/p$-actions on smooth surfaces in characteristic $p$ with isolated fixed points [3]. Namely, he constructed a $\text{\bf Z}/2$-action in characteristic 2 with quotient a du Val singularity of type $D_{4}$, a $\text{\bf Z}/3$-action in characteristic 3 with quotient an $E_{6}$ singularity, and a $\text{\bf Z}/5$-action in characteristic 5 with quotient an $E_{8}$ singularity. These special group actions arise globally from actions on del Pezzo surfaces, for example $\text{\bf Z}/5$ acting on the quintic del Pezzo surface (as discussed in section 8).
To show that our 3-dimensional quotients $W/G$ are terminal, the obvious approach would be to resolve the singularities of $W/G$ and make a calculation. Resolving these singularities is hard, however. We can greatly simplify the work by stopping at a partial resolution of $W/G$ that has toric singularities (specifically, $\mu_{p}$-quotient singularities, which we call tame quotient singularities); those are easy to analyze in combinatorial terms. (Recent advances suggest that an efficient substitute for resolving singularities in any characteristic would be to seek a resolution by a tame stack, rather than by a regular scheme [2, 21, 1].) Our key technical tool is Theorem 2.2, which gives a sufficient condition for a quotient scheme $U/G$ (where $G=\text{\bf Z}/p$, in positive or mixed characteristic) to have toric singularities.
These examples should lead to other failures of vanishing theorems. In particular, by Baudin, Bernasconi, and Kawakami, these examples imply that Frobenius-stable Grauert-Riemenschneider vanishing fails in characteristic 2, 3, and 5 [5, Remark 1.2].
This work was supported by NSF grant DMS-2054553, Simons Foundation grant SFI-MPS-SFM-00005512, and the Charles Simonyi Endowment at the Institute for Advanced Study.
1 Notation
We use the notation “$x=y+I$”, for elements $x$ and $y$ of a ring and an ideal $I$, to mean that there is an $i\in I$ such that $x=y+i$. We also use variants of this notation such as “$x=(y+I)(z+J)$”. Another variant (modeled on big-O notation in analysis) is to write “$x=y+O(z)$” for “$x=y+(z)$”.
We write $R\{x_{1},\ldots,x_{n}\}$ to mean the free module over a ring $R$ with basis elements $x_{1},\ldots,x_{n}$.
For a closed point $P$ in a regular scheme $U$ with residue field $k_{U}$, we say that $f_{1},\ldots,f_{n}$ are coordinates for $U$ at $P$ (or a regular system of parameters) if $f_{1},\ldots,f_{n}$ are elements of the maximal ideal $\operatorname{\mathfrak{m}}$ of $O(U)$ (the regular functions vanishing at $P$) that map to a basis for the $k_{U}$-vector space $\operatorname{\mathfrak{m}}/\operatorname{\mathfrak{m}}^{2}$.
For a group $G$ acting on a scheme $X$, $G$ acts on the ring of regular functions $O(X)$ by $(g(f))(x)=f(g^{-1}x)$, or equivalently $g(f)=(g^{-1})^{*}(f)$. The inverse is needed because of our convention of writing group actions on the left. Throughout the paper, we write $G=\text{\bf Z}/p=\langle\sigma:\sigma^{p}=1\rangle$ for the cyclic group of prime order $p$. We fix the name $\tau=\sigma^{-1}$, because of the inverse that comes up in writing the $G$-action on functions. Write $I(f)=\sigma(f)-f$ for a function $f$ on a $G$-scheme.
See section 5 for the definition of terminal singularities on general schemes, following [18, section 2.1].
For a positive integer $r$, let $\mu_{r}$ be the group scheme (over any base scheme) of $r$th roots of unity. The Reid-Tai criterion is the following description of which cyclic quotient singularities are canonical or terminal [23, Theorem 4.11]. This is often stated over a field, but it works even in mixed characteristic for $\mu_{r}$-quotient singularities. The point is that Kato’s theory of log regular schemes provides a mixed-characteristic analog of toric singularities, which includes the case of $\mu_{r}$-quotient singularities [15]. For such schemes, resolutions of singularities and the canonical divisor can be described in purely combinatorial terms. (The proof of Theorem 5.1 explains how to interpret the canonical divisor and terminal singularities in mixed characteristic.) In the following criterion, for $a$ an integer and $b$ a positive integer, consider $a\bmod b$ as an integer in the set $\{0,\ldots,b-1\}$.
Theorem 1.1.
For a positive integer $r$, let $\mu_{r}$ act on a regular scheme $X$, fixing a closed point $P$ with maximal ideal $\operatorname{\mathfrak{m}}$. Suppose that $\mu_{r}$ acts on a basis for $\operatorname{\mathfrak{m}}/\operatorname{\mathfrak{m}}^{2}$ by $\zeta(t_{1},\ldots,t_{n})=(\zeta^{b_{1}}t_{1},\ldots,\zeta^{b_{n}}t_{n})$, for some $b_{1},\ldots,b_{n}\in\text{\bf Z}/r$. (The quotient $X/\mu_{r}$ is said to be a $\mu_{r}$-quotient singularity of type $\frac{1}{r}(b_{1},\ldots,b_{n})$.) Assume that the action is well-formed in the sense that $\gcd(r,b_{1},\ldots,\widehat{b_{j}},\ldots,b_{n})=1$ for all $j=1,\ldots,n$. Then $X/\mu_{r}$ is canonical (resp. terminal) near the image of $P$ if and only if
$\sum_{j=1}^{n}ib_{j}\bmod r\geq r$
(resp. $>r$) for all $i=1,\ldots,r-1$.
We sometimes need the extension of the Reid-Tai criterion that describes when a toric pair is terminal, as follows. The proof is the same as Reid’s: $X/\mu_{r}$ has a toric resolution of singularities, and so it suffices to compute discrepancies for toric divisors over $X/\mu_{r}$.
Theorem 1.2.
Under the assumptions of Theorem 1.1, let $t_{1},\ldots,t_{n}$ be coordinates for $X$ at $P$ that are $\mu_{r}$-eigenfunctions with weights $b_{1},\ldots,b_{n}$. For $i=1,\ldots,n$, let $D_{i}$ be the irreducible divisor on $X/\mu_{r}$ that is the image of $\{t_{i}=0\}$ in $X$. Let $c_{1},\ldots,c_{n}$ be real numbers. Then the pair $(X,\sum c_{j}D_{j})$ is terminal near the image of $P$ if and only if $c_{j}<1$ for each $j$, $c_{j}+c_{k}<1$ for each $j\neq k$, and
$\sum_{j=1}^{n}(1-c_{j})(ib_{j}\bmod r)>r$
for all $i=1,\ldots,r-1$.
2 Recognizing $\mu_{p}$-quotient singularities
Király and Lütkebohmert analyzed when a quotient scheme by $\text{\bf Z}/p$ is regular, the hard case being where the residue characteristic is $p$ [16]. More generally, we now give a sufficient condition for a quotient by $\text{\bf Z}/p$ to be a toric singularity, or in particular when it is isomorphic to a $\mu_{p}$-quotient singularity. (Actions of $\mu_{p}$ are far simpler than actions of $\text{\bf Z}/p$: they are linearizable near a fixed point, like finite group actions in characteristic zero, because $\mu_{p}$ is linearly reductive. Also, $\mu_{p}$-quotient singularities are always Cohen-Macaulay.)
It would be very appealing to find a broader sufficient condition for a quotient by $\text{\bf Z}/p$ to have toric singularities, perhaps necessary and sufficient. Theorem 2.2 is adapted to the situation where the fixed point scheme is a Cartier divisor $E_{1}=\{e=0\}$ except on a closed subset of $E_{1}$. When it applies, the theorem can be described as a $\text{\bf Z}/p$$\mu_{p}$ switch.
Here is Király and Lütkebohmert’s main result [16, Theorem 2]. (See Theorem 2.6 for a more detailed statement.)
Theorem 2.1.
Let $G$ be a cyclic group of prime order which acts on a regular scheme $X$. If the fixed point scheme $X^{G}$ is a Cartier divisor in $X$, then the quotient space $X/G$ is regular.
Here is our sufficient condition for toric singularities. Recall that $I(f)$ means $\sigma(f)-f$, for a function $f$ on a scheme with an action of the group $G=\text{\bf Z}/p=\langle\sigma:\sigma^{p}=1\rangle$.
Theorem 2.2.
Let $U$ be a regular scheme with an action of the group $G=\text{\bf Z}/p=\langle\sigma:\sigma^{p}=1\rangle$, for a prime number $p$. Suppose that $G$ fixes a closed point $P$ with perfect residue field $k_{U}$ of characteristic $p$. Write $\operatorname{\mathfrak{m}}$ for the maximal ideal in the local ring $O_{U,P}$. Suppose that there are $e,s\in\operatorname{\mathfrak{m}}-\{0\}$ and coordinates $x_{1},\ldots,x_{n}$ for $U$ at $P$ such that
$I(s)=es(\text{unit})$
and
$I(x_{i})\in(e)$
for $i=1,\ldots,n$. Suppose that $p\in e^{p-1}\operatorname{\mathfrak{m}}$. (For example, that holds if $p=0$ on $U$.) Then $U/G$ is regular or has a $\mu_{p}$-quotient singularity at the image of $P$.
More precisely, if $I(x_{i})/e$ is nonzero at $P$ for some $i$, then $U/G$ is regular near $P$. Otherwise, let $\varphi(y)=I(y)/e$, which gives a linear map from $\operatorname{\mathfrak{m}}/\operatorname{\mathfrak{m}}^{2}$ to itself. Then $\varphi$ is diagonalizable, and after multiplying it by some nonzero scalar, its eigenvalues $b_{1},\ldots,b_{n}$ are in $\text{\bf F}_{p}\subset k_{U}$. In this case, $U/G$ has a singularity of the form $\frac{1}{p}(b_{1},\ldots,b_{n})$. That is, near the image of $P$, $U/G$ is the quotient of a regular scheme by an action of $\mu_{p}$ with weights $b_{1},\ldots,b_{n}$ at a fixed point.
Theorem 4.1 is a refinement of Theorem 2.2, showing that certain divisors in $U$ can be viewed as toric divisors in $U/G$.
We will often use the special case of Theorem 2.2 where $e=s=x_{1}$ for one of the coordinates $x_{1}$; in that case, we are assuming that $I(x_{1})=x_{1}^{2}(\text{unit})$ and $I(x_{i})\in(x_{1})$ for $i=1,\ldots,n$. Then the fixed point scheme $U^{G}$ is the regular divisor $\{x_{1}=0\}$ except on a closed subset of $E_{1}$. For other applications in this paper, we need the greater generality of Theorem 2.2, where $(U^{G})_{\operatorname{red}}$ may be a singular divisor.
One might ask whether the assumption that $p\in e^{p-1}\operatorname{\mathfrak{m}}$ in Theorem 2.2 can be omitted. Fortunately, this assumption is automatic in characteristic $p$, and it will be easy to check in our mixed characteristic examples.
Proof.
The assumption that $P$ is fixed by $G$ means that $G$ preserves $\operatorname{\mathfrak{m}}=\operatorname{\mathfrak{m}}_{U}$ and acts as the identity on the residue field $k_{U}:=O_{U,P}/\operatorname{\mathfrak{m}}$, which has characteristic $p$. For each regular function $f$ in the local ring $O_{U,P}$, its norm $N(f):=\prod_{i=0}^{p-1}\sigma^{i}(f)$ is $G$-invariant, and it maps to $f^{p}$ in $k_{U}$ by triviality of the $G$-action there. So the residue field $k_{U/G}\subset k_{U}$ at the image of the point $P$ contains $k_{U}^{p}$. Since we assume that $k_{U}$ is perfect, $k_{U/G}$ is equal to $k_{U}$. In what follows, we replace $U$ by a $G$-invariant open neighborhood of $P$ as needed, since we are only trying to describe $U/G$ near the image of $P$ (which we also call $P$). In particular, we can assume that $U$ is affine.
The fixed point scheme $U^{G}$ is defined as the closed subscheme of $U$ cut out by the ideal generated by $I(O(U))$. The following formulas hold for any action of $G$ on a commutative ring [16, Remark 3]:
Lemma 2.3.
(1) $I(xy)=I(x)\sigma(y)+xI(y).$
(2) For $m\geq 0$, $I(x^{m})=I(x)\sum_{i=1}^{m}\sigma(x)^{i-1}x^{m-i}.$
We are given coordinates $x_{1},\ldots,x_{n}$ for $U$ near $P$. Under our assumptions ($U$ regular and $k_{U/G}=k_{U}$), the ideal generated by $I(O_{U,P})$ in $O_{U,P}$ is generated by $I(x_{1}),\ldots,I(x_{n})$ [16, Proposition 6]. That is, after shrinking $U$ around $P$ if necessary, the fixed point scheme $U_{G}$ is the closed subscheme defined by $I(x_{1}),\ldots,I(x_{n})$.
In particular, if $I(x_{i})/e$ is a unit for some $i=1,\ldots,n$, then the fixed point scheme $U^{G}$ is the Cartier divisor $\{e=0\}$, and then Theorem 2.1 gives that $U/G$ is regular. So we can assume from now on that $I(x_{i})$ is in $e\operatorname{\mathfrak{m}}$ for each $i$. Equivalently, after shrinking $U$ around $P$, $I(O(U))$ is contained in $e\operatorname{\mathfrak{m}}$. In this case, we will show that $U/G$ has a $\mu_{p}$-quotient singularity at $P$.
The following lemma is implicit in the statement of Theorem 2.2.
Lemma 2.4.
Let $U$ be a regular scheme with an action of the group $G=\text{\bf Z}/p=\langle\sigma:\sigma^{p}=1\rangle$. Suppose that $G$ fixes a closed point $P$ with perfect residue field $k_{U}$ of characteristic $p$. Write $\operatorname{\mathfrak{m}}$ for the maximal ideal in the local ring $O_{U,P}$. Let $e\in\operatorname{\mathfrak{m}}$, $e\neq 0$, such that $I:=\sigma-1$ satisfies $I(\operatorname{\mathfrak{m}})\subset e\operatorname{\mathfrak{m}}$. Then $\varphi(y):=I(y)/e$ is a well-defined $k_{U}$-linear map from $\operatorname{\mathfrak{m}}/\operatorname{\mathfrak{m}}^{2}$ to itself.
Proof.
Since the ring $O_{U,P}$ is regular, it is a domain, and so $I(y)/e$ is well-defined for each element $y$ in $\operatorname{\mathfrak{m}}$. Since $G$ fixes the point $P$, $G$ maps $\operatorname{\mathfrak{m}}$ into itself. Since $I(\operatorname{\mathfrak{m}})\subset e\operatorname{\mathfrak{m}}$, Lemma 2.3 gives that $I(\operatorname{\mathfrak{m}}^{2})\subset e\operatorname{\mathfrak{m}}^{2}$. Since $I$ is additive, it follows that $\varphi$ is a well-defined additive function from $\operatorname{\mathfrak{m}}/\operatorname{\mathfrak{m}}^{2}$ to itself. By Lemma 2.3, $\varphi$ is linear over the ring of invariants $(O_{U,P})^{G}$. Since that ring has residue field $k_{U/G}=k_{U}$, $\varphi$ is $k_{U}$-linear. ∎
We are given a function $s\in\operatorname{\mathfrak{m}}-\{0\}$ such that $I(s)=es(\text{unit})$. After multiplying $e$ by a unit, we can assume that $I(s)=es$; this does not change the other assumption that $I(x_{i})\in(e)$ for $i=1,\ldots,n$. Thus, from now on, we have $I(s)=es$. This changes the endomorphism $\varphi$ of $\operatorname{\mathfrak{m}}/\operatorname{\mathfrak{m}}^{2}$ (which is defined in terms of $e$) by a nonzero scalar. Having made this change, we will show that $\varphi$ is diagonalizable, with eigenvalues $b_{1},\ldots,b_{n}$ in $\text{\bf F}_{p}\subset k_{U}$, and that $U/G$ has a $\mu_{p}$-quotient singularity of the form $\frac{1}{p}(b_{1},\ldots,b_{n})$.
Let $v=\sigma(s)/s$. By our assumptions, $v$ is a unit on $U$, and $v=1+e$. Since $v=\sigma(s)/s$, $v$ has norm 1 for the action of $G$. Write $(\sigma/\operatorname{id})(x)$ for $\sigma(x)/x$; this is the multiplicative action of $\sigma-1\in\text{\bf Z}G$ on a commutative ring. In these terms, define $f=(\sigma/\operatorname{id})^{p-2}(v)$. In the group ring $\text{\bf Z}G$, we have $(\sigma-1)^{p-1}\equiv\sigma^{p-1}+\cdots+\sigma+1\pmod{p}$, and so we can define an element $\alpha\in\text{\bf Z}G$ by
$(\sigma-1)^{p-1}=\sigma^{p-1}+\cdots+\sigma+1-p\alpha.$ ($\ast$)
It follows that $\sigma(f)/f=N(v)/g^{p}$, with $g:=\alpha(v)$ (where $\alpha$ acts multiplicatively). Since $v$ has norm 1, we have $\sigma(f)/f=1/g^{p}$. This formula will be exactly what we need to construct a $\mu_{p}$-torsor $W\to U$ with a commuting action of $G$.
We first analyze the function $g$ in more detail. By equation (*), the sum of the coefficients of $\alpha$ in $\text{\bf Z}G$ is 1. Therefore, $g=\alpha(v)$ is a product of integer powers of the functions $\sigma^{j}(v)=1+e(1+\operatorname{\mathfrak{m}})$ with total exponent 1, and so $g=1+e(1+\operatorname{\mathfrak{m}})$.
Lemma 2.5.
For any $p$th root of unity $\zeta$ in $O(U)$, $1-\zeta$ is in $e\operatorname{\mathfrak{m}}$.
Proof.
The statement is clear if $\zeta=1$. So assume that $\zeta\neq 1$. Since $O(U)$ is a domain and $(1-\zeta)(1+\zeta+\cdots+\zeta^{p-1})=1-\zeta^{p}=0$, we must have $01+\zeta+\cdots+\zeta^{p-1}=$ 0. That is, we have a homomorphism from the ring of integers $\text{\bf Z}[\mu_{p}]$ in $\text{\bf Q}(\mu_{p})$ to $O(U)$, taking a primitive $p$th root of unity to $\zeta$. In $\text{\bf Z}[\mu_{p}]$, $p$ is $(1-\zeta)^{p-1}$ times a unit [19, section IV.1], and so the same is true in $O(U)$. We are given that $p$ is in $e^{p-1}\operatorname{\mathfrak{m}}$. Since $U$ is regular, $O_{U,P}$ is a unique factorization domain, and so $1-\zeta$ must be a multiple of $e$; write $1-\zeta=ea$ for some $a\in O_{U,P}$. If $a$ is a unit, then $p$ would be $e^{p-1}$ times a unit, a contradiction. So, as an element of $O_{U,P}$, $1-\zeta$ is in $e\operatorname{\mathfrak{m}}$. After shrinking $U$ if necessary, this gives the same conclusion in $O(U)$. ∎
The reason for constructing units $f$ and $g$ with $\sigma(f)/f=1/g^{p}$ is to define a $\mu_{p}$-torsor over $U$ with a commuting action of $G$. Namely, define a $\mu_{p}$-torsor $W\to U$ by $w^{p}=f$. Here $\mu_{p}$ acts on $W\subset A^{1}\times U$ by $\zeta(w,x)=(\zeta w,x)$, for $\zeta\in\mu_{p}$. Write $\tau=\sigma^{-1}$ in $G$. Since $\sigma(f)/f=1/g^{p}$, $W$ has an action of $G$ that commutes with the action of $\mu_{p}$, by $\tau(w,x)=(w/g(x),\tau(x))$. (We check these properties in the next paragraph.) In particular, $\sigma(w)=\tau^{*}(w)=w/g$, by definition of the $G$-action on functions. We will show that the scheme $Q:=W/G$ is regular. Then $U/G=Q/\mu_{p}$ will be a quotient of a regular scheme by $\mu_{p}$, as we want.
[…]
For convenience, we write $P$ for the closed point of interest in each of these schemes. (There is a unique closed point in $W$ over $P$ in $U$, and it maps to a closed point in $Q$ and in $U/G=Q/\mu_{p}$.) We have seen that the points $P\in U$ and $P\in U/G$ have the same residue field $k_{U}$. Since $P\in W$ is given by the equations $w=1$ and $x_{1}=\cdots=x_{n}=0$, we see that $P\in W$ has the same residue field $k_{U}$. Since $k_{U/G}\subset k_{Q}\subset k_{W}$, $P\in Q$ also has the same residue field.
For clarity, let us first check that the formulas above give an action of $G$ on $W$ that commutes with the action of $\mu_{p}$. First, to show that $\sigma$ as above maps $W=\{w^{p}=f\}$ into itself, we have to show that if $w^{p}=f(x)$, then $(w/g(x))^{p}=f(\sigma^{-1}(x))$, or equivalently that $w^{p}/g^{p}=\sigma(f)$; this follows from the fact that $\sigma(f)/f=1/g^{p}$. Next, let us show that $\sigma^{p}=1$ on $W$. By induction, we have $\sigma^{-i}(w,x)=(w/(g(x)g(\sigma^{-1}x)\cdots g(\sigma^{1-i}x)),\sigma^{-i}x)$ for each natural number $i$. Therefore, to show that $\sigma^{p}$ is the identity on $W$ (hence that we have an action of $G$), it suffices to show that $g$ has norm 1. But that is true, because $v$ has norm 1 and $g$ is a product of integer powers of $\sigma^{j}(v)$ for integers $j$. So we have an action of $G$ on $W$. Finally, to show that $G$ and $\mu_{p}$ commute on $W$: for $\zeta\in\mu_{p}$, we have $\zeta\sigma^{-1}(w,x)=\zeta(w/g(x),\sigma^{-1}(x))=(\zeta w/g(x),\sigma^{-1}(x))$ while $\sigma^{-1}\zeta(w,x)=\sigma^{-1}(\zeta w,x)=(\zeta w/g(x),\sigma^{-1}(x))$. These are equal as regular functions on the scheme $\mu_{p}\times W$. So we have shown that $G$ and $\mu_{p}$ commute on $W$.
We need to show that $f$ is not a $p$th power in $O(U)$. Suppose it is, say $f=u^{p}$ for a regular function $u$ on $U$. Since $f$ is a unit, so is $u$. Then $1/g^{p}=\sigma(f)/f=(\sigma(u)/u)^{p}$, and so $\zeta/g=\sigma(u)/u$ for some $p$th root of unity $\zeta$ in $O(U)$. Here $\zeta=1+(\zeta-1)=1+e\operatorname{\mathfrak{m}}$, by Lemma 2.5, and $1/g=1-e(1+\operatorname{\mathfrak{m}})$, so $\sigma(u)/u=1-e(1+\operatorname{\mathfrak{m}})$. Here $\sigma(u)/u=1+I(u)/u$, so $I(u)=ue(1+\operatorname{\mathfrak{m}})=e(\text{unit})$. This contradicts our assumption that $I(O(U))$ is contained in $e\operatorname{\mathfrak{m}}$. So in fact $f$ is not a $p$th power.
From there, we can show that the scheme $W$ is integral (after shrinking $U$ near $P$, if necessary). Namely, since $f$ is not a $p$th power in $O(U)$, $f$ is also not a $p$th power in the function field $k(U)$, and so $k(U)[f^{1/p}]$ is a degree-$p$ field extension of $k(U)$. Write $\alpha$ for the $\mu_{p}$-torsor $W\to U$. Since $W=\{w^{p}=f\}$, there is a nonempty open subset $V\subset U$ with $\alpha^{-1}(V)$ integral. Since $W\to U$ is finite and flat, $W$ is Cohen-Macaulay and equidimensional. By equidimensionality, every irreducible component of $W$ must dominate $U$. Since $\alpha^{-1}(V)$ is irreducible, it follows that $W$ is irreducible. Since $\alpha^{-1}(V)$ is reduced, $W$ is reduced in codimension 0; since $W$ is Cohen-Macaulay, it follows that $W$ is reduced [24, Tag 031R]. Since $W$ is reduced and irreducible, it is integral.
It is not needed for what follows, but for clarity, let us analyze the singularities of $W$ in the special case where $s=e(1+\operatorname{\mathfrak{m}})$ (for example when $s=e$), so that $I(e)=e^{2}(1+\operatorname{\mathfrak{m}})$. The equation of $W$ is $w^{p}=f$. Since $v=1+e$, one can show by induction from the formula for $I(e)$ that $f=(\sigma/\operatorname{id})^{p-2}(v)=1+e^{p-1}q$ for some unit $q$ on $U$, and so we can rewrite the equation of $W$ as $w^{p}=1+e^{p-1}q$. Let $w_{0}=w-1$, so that $w_{0}$ vanishes at the point $P$ of interest in $W$. In terms of $w_{0}$, the equation of $W$ becomes $(1+w_{0})^{p}=1+e^{p-1}q$, that is,
$w_{0}^{p}=e^{p-1}q-\binom{p}{1}w_{0}-\cdots-\binom{p}{p-1}w_{0}^{p-1}.$
We are given that $p$ is in $e^{p-1}\operatorname{\mathfrak{m}}$, and so this equation has the form $w_{0}^{p}=e^{p-1}s$ for some unit $s$ on $W$. If $p=2$ and $e\not\in\operatorname{\mathfrak{m}}^{2}$, it follows that $W$ is regular. However, if $p>2$, then $W$ is not normal. For example, if $e\not\in\operatorname{\mathfrak{m}}^{2}$, then the singularity of $W$ looks roughly like the cuspidal curve $\{w_{0}^{p}=x_{1}^{p-1}\}$ times a smooth variety.
We will need the following version of Király and Lütkebohmert’s results.
Theorem 2.6.
  1. (1)
    Let $B$ be a local domain with residue field $k_{B}$. Let $p$ be a prime number, and let $G=\text{\bf Z}/p$ act nontrivially on $B$. Suppose that the ideal $B\cdot I(B)$ that defines the fixed point scheme in $\operatorname{Spec}B$ is generated by one element. Then $B$ is free of rank $p$ over the ring of invariants $B^{G}$. More precisely, for any element $t$ such that $I(t)$ generates the ideal $B\cdot I(B)$, we have $B=A\{1,t,\ldots,t^{p-1}\}$.
  2. (2)
    In addition to the assumptions of (1), assume that $B$ is regular. Then $B^{G}$ is regular.
  3. (3)
    In addition to the assumptions of (1), assume that $B$ is noetherian and the inclusion $k_{B^{G}}\subset k_{B}$ is an equality. Then there is a minimal set of generators $y_{1},\ldots,y_{r}$ for $\operatorname{\mathfrak{m}}_{B}$ such that $I(y_{1})$ generates the ideal $B\cdot I(B)$ and $y_{2},\ldots,y_{r}$ are $G$-invariants.
Proof.
Statement (1) is due to Király and Lütkebohmert for $B$ normal, but their proof works without change for $B$ a domain [16, Theorem 2 and Proposition 5]. They also prove statement (2). They prove statement (3) only when $B$ is regular. We now extend the proof of (3) for $B$ only a domain.
Since the inclusion $k_{B^{G}}\subset k_{B}$ is an equality, we have $B=B^{G}+\operatorname{\mathfrak{m}}_{B}$, and so $I(B)=I(\operatorname{\mathfrak{m}}_{B})$. Since the ideal $B\cdot I(B)$ is generated by one element, there is an element $y_{1}\in\operatorname{\mathfrak{m}}_{B}$ such that $I(y_{1})$ generates this ideal. By Lemma 2.3, we have $I(\operatorname{\mathfrak{m}}_{B}^{2})\subset\operatorname{\mathfrak{m}}_{B}I(\operatorname{\mathfrak{m}}_{B})$. Here $I(\operatorname{\mathfrak{m}}_{B})$ is not zero (because the $G$-action is nontrivial), and so $y_{1}$ is not in $\operatorname{\mathfrak{m}}_{B}^{2}$.
Since $B$ is noetherian, the ideal $\operatorname{\mathfrak{m}}_{B}$ is finitely generated. Choose elements $z_{2},\ldots,z_{r}$ in $\operatorname{\mathfrak{m}}_{B}$ such that $y_{1},z_{2},\ldots,z_{r}$ form a basis for $\operatorname{\mathfrak{m}}_{B}/\operatorname{\mathfrak{m}}_{B}^{2}$. By part (1), we know that $B=B^{G}\{1,y_{1},\ldots,y_{1}^{p-1}\}$. For $2\leq i\leq r$, let $y_{i}$ be the projection of $z_{i}$ to $B^{G}$ with respect to this decomposition. Then $y_{i}\equiv z_{i}\pmod{(y_{1})+\operatorname{\mathfrak{m}}_{B}^{2}}$, and so $y_{1},\ldots y_{r}$ map to a basis for $\operatorname{\mathfrak{m}}_{B}/\operatorname{\mathfrak{m}}_{B}^{2}$. Thus $y_{1},\ldots,y_{r}$ are a minimal set of generators for $\operatorname{\mathfrak{m}}_{B}$ (by Nakayama’s lemma), and $y_{2},\ldots,y_{r}$ are $G$-invariant. ∎
Let us write out the action of $G$ on $W$. The maximal ideal of $P$ in $W$ is generated by $w_{0},x_{1},\ldots,x_{n}$. We have $I(w_{0})=I(w)=w(\frac{1}{g}-1)=(1+w_{0})(\frac{1}{g}-1)$. Since $1/g=1+eu$ for some unit $u$ on $U$, we have $I(w_{0})=eu(1+w_{0})=e(\text{unit})$. We also have $I(x_{i})$ in the ideal $(e)$ for $i=1,\ldots,n$; so the fixed point scheme $W^{G}$ is defined by the single equation $e=0$ in $W$. As a result, even though $W$ is typically not normal, Theorem 2.6 gives that the morphism $W\to W/G=Q$ is finite and faithfully flat of degree $p$. It follows that $Q$ is noetherian [24, Tag 033E]. (Beware that for a general noetherian scheme $X$ with an action of a finite group $G$, $X/G$ need not be noetherian, and the morphism $X\to X/G$ need not be finite [22, Proposition 0.10]. These properties do hold if $X$ is of finite type over a noetherian ring $A$ and $G$ acts $A$-linearly [11, Theorem and Corollary 4].)
The action of $\mu_{p}$ on the affine scheme $Q$ gives a grading of $O(Q)$ by $\text{\bf Z}/p$. For each $j\in\text{\bf Z}/p$, since $O(Q)$ is noetherian, the ideal in $O(Q)$ generated by the $j$th graded piece $O(Q)_{j}$ is finitely generated, and so $O(Q)_{j}$ is a finitely generated module over $O(Q)_{0}=O(Q/\mu_{p})$. So the whole ring $O(Q)$ is finite over $O(Q/\mu_{p})$; that is, $Q\to Q/\mu_{p}$ is finite. Also, $O(Q/\mu_{p})$ is a pure subring (because it is a summand) of the noetherian ring $O(Q)$, so it is noetherian [14, Proposition 6.15]; that is, $Q/\mu_{p}=U/G$ is noetherian. Finally, the composition $W\to Q\to Q/\mu_{p}=U/G$ is finite, and $O(U)$ is a sub-$O(U/G)$-module of $O(W)$, so $O(U)$ is a finitely generated $O(U/G)$-module; that is, $U\to U/G$ is finite.
Let $h_{1},\ldots,h_{r}$ be a minimal set of generators for the maximal ideal at $P$ of $O(Q)$. (So $r$ is at least the dimension $n$ of $Q$.)
Lemma 2.7.
The ideals $(h_{1},\ldots,h_{r})$ and $(x_{1},\ldots,x_{n})$ in $O(W)$ are equal. (That is: the fiber in $W$ over the closed point $P\in Q$ is equal to the fiber in $W$ over the closed point $P\in U$, as a closed subscheme.)
Proof.
We have seen that the degree-$p$ morphism $W\to W/G=Q$ is finite and flat. So the fiber in $W$ over the point $P$ in $Q$ has degree $p$ over the residue field of $P\in Q$, which we have seen is $k_{U}$. As a set, this fiber is one point $P\in W$, with the same residue field $k_{U}$. So the quotient ring $O(W)/(h_{1},\ldots,h_{r})$ is an artinian local ring of length $p$.
The ideal $(x_{1},\ldots,x_{n})$ in $O(W)$ defines the fiber in $W$ over the point $P$ in $U$. Since $W\to U$ is a $\mu_{p}$-torsor, this fiber has degree $p$ over the residue field of $P\in U$, which is the same field $k_{U}$. Again, this fiber is one point $P\in W$ as a set, with the same residue field; so $O(W)/(x_{1},\ldots,x_{n})$ is an artinian local ring of length $p$. So if we can show that the ideal $(h_{1},\ldots,h_{r})$ in $O(W)$ is contained in $(x_{1},\ldots,x_{n})$ in $O(W)$, then they are equal.
It suffices to show (*) that every function $y$ on $W$ that vanishes at the point $P$ in $W$ but has nonzero image in $O(W)/(x_{1},\ldots,x_{n})$ has $I(y)\neq 0$ (that is, it is not $G$-invariant). (Namely, this would imply that the $G$-invariant functions $h_{1},\ldots,h_{n}$ on $W$ lie in the ideal $(x_{1},\ldots,x_{n})$, as we want.) By the formula for the $G$-action on $W$, in particular that $\sigma(w)=w/g$ where $g=1+e(1+\operatorname{\mathfrak{m}})$, we see that $G$ fixes the closed subscheme $\{e=0\}$ in $W$. That is, $I$ maps $O(W)$ into the ideal $(e)$ in $O(W)$. Also, we know that $I(x_{i})\in e\operatorname{\mathfrak{m}}_{U}=e(x_{1},\ldots,x_{n})\subset O(U)$ for $i=1,\ldots,n$. So $\varphi(y):=I(y)/e$ is a well-defined linear map from $O(W)/(x_{1},\ldots,x_{n})$ to itself. Explicitly, by the equation of $W$, $O(W)/(x_{1},\ldots,x_{n})$ is a $k_{U}$-vector space with basis $1,w,\ldots,w^{p-1}$. Equivalently, in terms of $w_{0}=w-1$ (which vanishes at $P$ in $W$), a basis for $O(W)/(x_{1},\ldots,x_{n})$ is given by $1,w_{0},\ldots,w_{0}^{p-1}$.
The claim (*) will follow if the map $\varphi$ restricted to $k\{w_{0},w_{0}^{2},\ldots,w_{0}^{p-1}\}$ is injective. Since $g=1+e(1+\operatorname{\mathfrak{m}}_{U})$, we have $1/g=1-e(1+\operatorname{\mathfrak{m}}_{U})$, and hence $I(w_{0})=I(w)=(w/g)-w=-ew(1+\operatorname{\mathfrak{m}}_{U})=-e(1+w_{0})(1+\operatorname{\mathfrak{m}}_{U})$. So $\varphi(w_{0})=-(1+w_{0})$. By Lemma 2.3, for $m\geq 0$,
$I(w_{0}^{m})$ $=I(w_{0})\sum_{j=1}^{m}\sigma(w_{0})^{j-1}w_{0}^{m-j}$
$=-e(1+w_{0})(1+\operatorname{\mathfrak{m}}_{U})\sum_{j=1}^{m}(w_{0}-(1+w_{0})e(1+\operatorname{\mathfrak{m}}_{U}))^{j-1}w_{0}^{m-j}.$
Since $\varphi$ takes values in $O(W)/(x_{1},\ldots,x_{n})$ (where $e$ is zero), it follows that $\varphi(w_{0}^{m})=-mw_{0}^{m}(1+w_{0})$. It is clear that these elements are linearly independent over $k_{U}$ for $m=1,\ldots,p-2$; to show that they are linearly independent for $m=1,\ldots,p-1$, it will suffice to show that $w_{0}^{p}$ is zero in $O(W)/(x_{1},\ldots,x_{n})=k_{U}\{1,w_{0},\ldots,w_{0}^{p-1}\}$. (This comes up because $w_{0}^{p}$ appears in our formula for $I(w_{0}^{p-1})$.)
Namely, we have $w_{0}^{p}=(w-1)^{p}=w^{p}-1$ plus a multiple of $p$ in $O(W)$. Here $w^{p}-1=f-1=e^{p-1}(1+\operatorname{\mathfrak{m}}_{U})$, and $p$ is in $e^{p-1}\operatorname{\mathfrak{m}}_{U}$, as we assumed; so $w_{0}^{p}=e^{p-1}(1+\operatorname{\mathfrak{m}}_{U})$, which is zero in $O(W)/(x_{1},\ldots,x_{n})$, as we want. Lemma 2.7 is proved. ∎
The number $r$ of generators $h_{1},\ldots,h_{r}$ for the maximal ideal $\operatorname{\mathfrak{m}}_{Q}$ in the local ring $O_{Q,P}$ is at least $n=\dim(Q)$. On the other hand, Lemma 2.7 implies that the extended ideal $(h_{1},\ldots,h_{r})$ in $O_{W,P}$ can be generated by only $n$ elements, so the vector space $(h_{1},\ldots,h_{r})\otimes_{O_{W,P}}k_{W}$ has dimension at most $n$. So this vector space is spanned by $n$ of the $h_{i}$’s, which we can assume are $h_{1},\ldots,h_{n}$. By Nakayama’s lemma, it follows that the extended ideal $(h_{1},\ldots,h_{r})$ is equal to the extended ideal $(h_{1},\ldots,h_{n})$ in $O_{W,P}$. Since $W\to Q$ is faithfully flat, extending and contracting an ideal in $O_{Q,P}$ gives the same ideal [20, Theorem 7.5]. As a result, we have $(h_{1},\ldots,h_{n})=(h_{1},\ldots,h_{r})$ in $O_{Q,P}$. That is, the maximal ideal $\operatorname{\mathfrak{m}}_{Q}$ can be generated by $n=\dim(Q)$ elements, which means that $Q$ is regular. (This is somewhat surprising, since $W$ is typically not regular or even normal.)
It remains to show that the point $P$ in $Q$ is a fixed point for $\mu_{p}$, with weights given by the eigenvalues of $\varphi$. First, let us show that $\mu_{p}$ fixes the point $P$ in $Q$. (This does not seem obvious, since $\mu_{p}$ does not fix the point $P$ in $W$; in fact, $\mu_{p}$ acts freely on $W$.) The functions $x_{1},\ldots,x_{n}$ on $W$ are pulled back from $U$, hence fixed by $\mu_{p}$. As a result, the ideal $(x_{1},\ldots,x_{n})$ in $O(W)$ is preserved by $\mu_{p}$. Equivalently, by Lemma 2.7, the ideal $(h_{1},\ldots,h_{n})$ in $O(W)$ is preserved by $\mu_{p}$. We have seen that the morphism $W\to W/G=Q$ is faithfully flat. As a result, the ideal $(h_{1},\ldots,h_{n})$ in $O(Q)$ is equal to the intersection of the extended ideal $(h_{1},\ldots,h_{n})$ in $O(W)$ with $O(Q)$. Since $W\to Q$ is $\mu_{p}$-equivariant, it follows that the ideal $(h_{1},\ldots,h_{n})=\operatorname{\mathfrak{m}}_{Q}$ in $O(Q)$ is preserved by $\mu_{p}$. Also, the residue field of $Q/\mu_{p}$ at $P$ maps isomorphically to the residue field of $Q$ at $P$, and so $\mu_{p}$ acts trivially on the latter field. That is, $\mu_{p}$ fixes the point $P$ in $Q$, as we want.
We now change our choice of the functions $h_{1},\ldots,h_{n}$. Since $\mu_{p}$ is linearly reductive, we can choose coordinates $h_{1},\ldots,h_{n}$ for $Q$ near $P$ that are $\mu_{p}$-eigenfunctions. That is, each $h_{i}$ has some weight $b_{i}\in\text{\bf Z}/p$ for the action of $\mu_{p}$. In these terms, $Q/\mu_{p}=U/G$ is a toric singularity of type $\frac{1}{p}(b_{1},\ldots,b_{n})$. It remains to show that the endomorphism $\varphi$ of $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$ is diagonalizable, with eigenvalues in $\text{\bf F}_{p}\subset k_{U}$, and that these eigenvalues are equal to $b_{1},\ldots,b_{n}$.
Consider $h_{1},\ldots,h_{n}$ as $G$-invariant functions on $W$. Here $W$ is a $\mu_{p}$-torsor over $U$ defined by $w^{p}=f$; so $O(W)=O(U)\{1,w,\ldots,w^{p-1}\}$, and this grading by $\text{\bf Z}/p$ describes the action of $\mu_{p}$ on $O(W)$. For $i=1,\ldots,n$, $h_{i}$ has weight $b_{i}$ for the action of $\mu_{p}$, and so we can write $h_{i}=g_{i}w^{b_{i}}$ for some regular function $g_{i}$ on $U$. (Here we think of $b_{i}$ as an integer in $\{0,\ldots,p-1\}$.) Clearly the functions $g_{1},\ldots,g_{n}$ vanish at $P$ (using that $w$ is a unit). Also, $(g_{1},\ldots,g_{n})$ is equal to $(h_{1},\ldots,h_{n})$ as an ideal in $O(W)$, and we showed that the latter ideal is equal to $(x_{1},\ldots,x_{n})$ in $O(W)$. Since $W\to U$ is faithfully flat, it follows that $(g_{1},\ldots,g_{n})$ is equal to $(x_{1},\ldots,x_{n})$ as an ideal in $O(U)$. That is, $g_{1},\ldots,g_{n}$ form coordinates on $U$ near $P$.
For $1\leq i\leq n$, we have $I(g_{i})\equiv e\varphi(g_{i})\in e(\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2})$, by definition of the endomorphism $\varphi$ of $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$. We showed above that $I(w_{0})=-e(1+w_{0})(1+\operatorname{\mathfrak{m}}_{U})$, and so $I(w)=I(w_{0})=-ew(1+\operatorname{\mathfrak{m}}_{U})$. For each $b\geq 0$, Lemma 2.3 gives that
$I(w^{b})$ $=I(w)\sum_{m=1}^{b}\sigma(w)^{m-1}w^{b-m}$
$=-bew^{b}(1+\operatorname{\mathfrak{m}}_{U}),$
using that $e$ is in $\operatorname{\mathfrak{m}}_{U}$. Since the function $g_{i}w^{b_{i}}$ is $G$-invariant on $W$, we have $0=I(g_{i}w^{b_{i}})=\sigma(g_{i})I(w^{b_{i}})+I(g_{i})w^{b_{i}}=(g_{i}+I(g_{i}))I(w^{b_{i}})+I(g_{i})w^{b_{i}}$. We want to consider this equality modulo $e\operatorname{\mathfrak{m}}_{U}^{2}O(W)$, and so the term $I(g_{i})I(w^{b_{i}})$ can be omitted. Namely, we have
00 0 $\equiv e(-b_{i}g_{i}w^{g_{i}}(1+\operatorname{\mathfrak{m}})+\varphi(g_{i})w^{g_{i}})\pmod{e\operatorname{\mathfrak{m}}_{U}^{2}O(W)}$
$\equiv ew^{g_{i}}(-b_{i}g_{i}+\varphi(g_{i}))\pmod{e\operatorname{\mathfrak{m}}_{U}^{2}O(W)}.$
Since $w$ is a unit, it follows that $0\equiv e(-b_{i}g_{i}+\varphi(g_{i}))\pmod{e\operatorname{\mathfrak{m}}_{U}^{2}O(W)}$. Since $W\to U$ is faithfully flat, $e\operatorname{\mathfrak{m}}_{U}/e\operatorname{\mathfrak{m}}_{U}^{2}\to e\operatorname{\mathfrak{m}}_{U}O(W)/e\operatorname{\mathfrak{m}}_{U}^{2}O(W)$ is injective, and so $0\equiv e(-b_{i}g_{i}+\varphi(g_{i}))\pmod{e\operatorname{\mathfrak{m}}_{U}^{2}}$. So $\varphi(g_{i})=b_{i}g_{i}$ in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$ for each $i=1,\ldots,n$. Also, $g_{1},\ldots,g_{n}$ form a basis for $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$. So $\varphi$ is diagonalizable, its eigenvalues $b_{1},\ldots,b_{n}$ are in $\text{\bf F}_{p}$, and $U/G$ is a $\mu_{p}$-quotient singularity of the form $\frac{1}{p}(b_{1},\ldots,b_{n})$, as we want. ∎
3 Review of ramification theory
We recall here how to compute the ramification behavior of a $\text{\bf Z}/p$-covering in characteristic $p$ or mixed characteristic, following Xiao and Zhukov [27].
Let $G=\text{\bf Z}/p=\langle\sigma:\sigma^{p}=1\rangle$ act nontrivially on a normal noetherian integral scheme $Y$. Assume that $Y$ is of finite type over a field or over $\text{\bf Z}_{p}$, so that we can talk about the canonical class $K_{Y}$. Write $f$ for the quotient map $Y\rightarrow Y/G$. For each irreducible divisor $E$ in $Y$ that is mapped into itself by $G$, let $F$ be its image (as an irreducible divisor) in $Y/G$. Assume that $p=0$ on $E$. There are several invariants we want to compute in this situation: the ramification index of the divisor $E$ in $Y$ (the positive integer $e$ such that $f^{*}F=eE$), and the coefficient $c$ of $E$ in the ramification divisor (meaning that $K_{Y}=f^{*}K_{Y/G}+cE$ near the generic point of $E$). Another name for $c$ is the valuation of the different $v_{E}(\mathcal{D}_{L/K})$, where $L$ is the function field $k(Y)$, $K=k(Y/G)$, and the valuation $v_{E}$ is the order of vanishing along $E$. Here $ef=p$, where $f$ is the degree of the field extension $k(E)$ over $k(F)$.
An easy case is where $G$ does not fix $E$ (in other words, where $G$ acts nontrivially on $E$). We say that $f$ is unramified along $E$. In this case, $f^{*}F=E$, and $K_{Y}=f^{*}K_{Y/G}$ near the generic point of $E$.
In the ramified case, the relative canonical class (that is, the valuation of the different) can be computed as follows. Define the Artin ramification number $i(E)$ of $Y$ over $Y/G$ along $E$ to be the coefficient of $E$ in the fixed point scheme $Y^{G}$. Equivalently, in terms of $I(a)=\sigma(a)-a$:
$i(E)=\min_{a\in O_{Y,E}}v_{E}(I(a)).$
Then the valuation of the different is $(p-1)i(E)$ [27, section 2.1]. When $Y$ is regular, so that $Y/G$ is Q-factorial, we can equivalently say that
$K_{Y}=f^{*}K_{Y/G}+(p-1)[Y^{G}],$
where $[Y^{G}]$ denotes the Weil divisor associated to the fixed point scheme.
In the ramified case, the field $k(E)$ is purely inseparable over $k(F)$, with degree $f$ equal to 1 or $p$. We can distinguish the two cases as follows. Since $ef=p$, either $e=p$ and $f=1$ (called wild ramification) or $e=1$ and $f=p$ (called fierce ramification). Let $t$ be a defining function of $E$, that is, a rational function on $Y$ with valuation $v_{E}(t)=1$. It is clear that $v_{E}(I(t))\geq i(E)$. A very convenient criterion is: $Y\to Y/G$ is wildly ramified along $E$ if and only if equality holds [27, section 2.1]. Otherwise, it is fiercely ramified.
4 Toric divisors
In addition to recognizing when a quotient by $G=\text{\bf Z}/p$ has a $\mu_{p}$-quotient singularity (as in Theorem 2.2), Theorem 4.1 analyzes when a $G$-invariant divisor is pulled back from a divisor on the quotient. Using this, we can view certain $G$-invariant divisors as toric divisors on the quotient scheme, which will be convenient for applications.
Theorem 4.1.
  1. (1)
    Let $G=\text{\bf Z}/p$ act on a regular scheme $U$, with the assumptions of Theorem 2.2. Assume moreover that the function $e$ is the greatest common divisor of the functions $I(x_{i})$ for $i=1,\ldots,n$ in the local ring $O_{U,P}$. Then, for each $y\in\operatorname{\mathfrak{m}}_{U}$ such that $\{y=0\}$ is an irreducible divisor $D\subset U$ and $I(y)\in(ey)$, $D$ is the pullback of a Weil divisor $F$ in $U/G$.
  2. (2)
    If in addition $I(x_{i})\in e\operatorname{\mathfrak{m}}_{U}$ for $i=1,\ldots,n$, so that $U/G$ is the quotient of a regular scheme $Q$ by $\mu_{p}$ (by Theorem 2.2), then the pullback of $F$ to $Q$ is the divisor $\{h=0\}$ for some $\mu_{p}$-eigenfunction $h$ on $Q$.
  3. (3)
    Continue to assume that $I(x_{i})\in e\operatorname{\mathfrak{m}}_{U}$ for $i=1,\ldots,n$, so that $U/G$ is the quotient of a regular scheme $Q$ by $\mu_{p}$. Let $y_{1},\ldots,y_{r}$ be functions on $U$, vanishing at $P$, that are linearly independent in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$. Suppose that $I(y_{j})\in(ey_{j})$ for $j=1,\ldots,r$. Then the corresponding $\mu_{p}$-eigenfunctions $h_{1},\ldots,h_{r}$ on $Q$ (from (2)) are linearly independent in $\operatorname{\mathfrak{m}}_{Q}/\operatorname{\mathfrak{m}}_{Q}^{2}$. That is, these functions are part of a toric coordinate system on $Q$. Finally, if $I(s)=es$ in Theorem 2.2 (as we can assume), then the $\mu_{p}$-weight of $h_{j}$ is equal to the eigenvalue of $\varphi$ on $y_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$, which is in $\text{\bf F}_{p}$.
Proof.
(1) Let $E$ be the irreducible divisor $\{y=0\}$ on $U$. The assumptions imply that $E$ is mapped into itself by $G$, and that there is an $i\in\{1,\ldots,n\}$ such that $v_{E}(I(y))>v_{E}(I(x_{i}))$. By section 3, $f\colon U\to U/G$ is either unramified (if $E$ is not fixed by $G$) or fiercely ramified along $E$. In either case, there is an irreducible divisor $F$ on $U/G$ such that $E=f^{*}F$, as we want. (The divisor $F$ need not be Cartier, but the pullback of a Weil divisor is still a Weil divisor (with integer coefficients). Indeed, $F$ is a Cartier divisor outside a codimension-2 subset of $U/G$, by normality of $U/G$.)
(2) Since $Q$ is regular, the pullback of $F$ to $Q$ is a Cartier divisor, hence (after possibly shrinking $U$ and $Q$) of the form $\{t=0\}$ for some function $t\in\operatorname{\mathfrak{m}}_{Q}-\{0\}$. Clearly the divisor $\{t=0\}$ is $\mu_{p}$-invariant. I claim that $t$ times some unit is a $\mu_{p}$-eigenfunction on $Q$. Indeed, in algebraic terms, the action of $\mu_{p}$ on $Q$ makes $O(Q)$ a comodule over $O(\mu_{p})$, and the ideal $(t)$ is an sub-$O(\mu_{p})$-comodule. Every $O(\mu_{p})$-comodule (with no finiteness assumption needed) is the direct sum of its weight spaces, indexed by $\text{\bf Z}/p$. So we can write $t=t_{0}+\cdots+t_{p-1}$ with $t_{i}\in(t)$ and $t_{i}$ of weight $i$. Since $t_{i}\in(t)$, we can write $t_{i}=a_{i}t$ for some $a_{i}\in O_{Q,P}$. Since $O_{Q,P}$ is regular, it is a domain, and hence $1=a_{0}+\cdots+a_{p-1}$. So at least one $a_{i}$ is not in $\operatorname{\mathfrak{m}}_{Q}$, hence is a unit. Then $h:=t_{i}=a_{i}t$ is a unit times $t$ and also a $\mu_{p}$-eigenfunction (of weight $i$), as we want.
(3) After multiplying $e$ by a unit, we can assume that $I(s)=es$, in the terminology of Theorem 2.2. The assumption that $I(y_{j})\in(ey_{j})$ for $j=1,\ldots,r$ implies that $y_{1},\ldots,y_{r}$ in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$ are eigenfunctions of the map $\varphi$. By the proof of Theorem 2.2, the corresponding eigenvalues are in $\text{\bf F}_{p}\subset k_{U}$.
For $j=1,\ldots,r$, we know from (1) that the divisor $\{y_{j}=0\}$ on $U$ is pulled back from a divisor $F$ on $U/G$. (Here $F$ is a Weil divisor, but it is a Cartier divisor outside a codimension-2 subset of $U/G$, since $U/G$ is normal.) By (2), $F$ pulls back to a divisor $\{h_{j}=0\}$ on $Q$ with $h_{j}$ a $\mu_{p}$-eigenfunction. It follows that the Cartier divisors $\{y_{j}=0\}$ on $U$ and $\{h_{j}=0\}$ on $Q$ have the same pullback to $W$; that is, $h_{j}=y_{j}(\text{unit})$ on $W$. Since $h_{j}$ is a $\mu_{p}$-eigenfunction of some weight $b_{j}\in\{0,\ldots,p-1\}$, we have $h_{j}=g_{j}w^{b_{j}}$ for some function $g_{j}$ on $U$, in the notation of the proof of Theorem 2.2. Therefore, $g_{j}=y_{j}(\text{unit})$ on $U$. Since $y_{1},\ldots,y_{r}$ are linearly independent in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$, the same is true for $g_{1},\ldots,g_{r}$. By the proof of Theorem 2.2, the $\mu_{p}$-weight of $h_{j}$ is equal to the eigenvalue of $\varphi$ on the eigenvector $g_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$. Since $g_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$ is a nonzero multiple of $y_{j}$, this is the same as the eigenvalue of $\varphi$ on $y_{j}$.
The ring $O_{W,P}$ is faithfully flat over $O_{U,P}$ and over $O_{Q,P}$. By Lemma 2.7, the maximal ideals $\operatorname{\mathfrak{m}}_{U}$ and $\operatorname{\mathfrak{m}}_{Q}$ generate the same ideal in $O_{W,P}$. Since $g_{1},\ldots,g_{r}$ are linearly independent in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$, it follows that $h_{1},\ldots,h_{r}$ are linearly independent in $\operatorname{\mathfrak{m}}_{Q}/\operatorname{\mathfrak{m}}_{Q}^{2}$, as we want. ∎
5 The example over the 2-adic integers
Theorem 5.1.
Let $Y=\{(x,y,i)\in A^{3}_{\text{\bf Z}_{2}}:x\neq 0,y\neq 0,i^{2}=-1\}$. Let the group $G=\text{\bf Z}/2=\{1,\sigma\}$ act on $Y$ by
$\sigma(x,y,i)=(1/x,1/y,-i).$
Then the scheme $Y/G$ is terminal, not Cohen-Macaulay, of dimension 3, and flat over $\text{\bf Z}_{2}$. Also, the canonical class of $Y/G$ over $\text{\bf Z}_{2}$ is Cartier.
Proof.
The scheme $Y$ is regular, being an open subset of the affine plane over the discrete valuation ring $\text{\bf Z}_{2}[\mu_{4}]=\text{\bf Z}_{2}[i]/(i^{2}+1)$. Since $Y$ is a normal integral affine scheme of dimension 3, so is $Y/G$. The ring $O(Y/G)$ of regular functions is a torsion-free $\text{\bf Z}_{2}$-module, since it is a subring of the torsion-free $\text{\bf Z}_{2}$-module $O(Y)$; so $Y/G$ is flat over $\text{\bf Z}_{2}$. The fixed point scheme of $G$ on $Y$ is defined by: $I(x)=(1-x^{2})/x$, and $I(y)=(1-y^{2})/y$, and $I(i)=\sigma(i)-i=-2i$, hence by $\{x^{2}=1,y^{2}=1,2i=0\}$. Together with the equation $i^{2}=-1$ on $Y$, these equations imply set-theoretically that $2=0$, $x=1$, $y=1$, and $i=1$; so the fixed point set of $G$ is a single closed point $P$ in $Y$, with residue field $\text{\bf F}_{2}$. Since $Y$ is regular, it follows that $Y/G$ is regular outside the image of $P$, which we also call $P$.
For $Y/G$ to be Cohen-Macaulay at $P$ would mean that the local cohomology $H^{i}_{P}(Y/G,O)$ was zero for $i<\dim(Y/G)=3$. Consider the exact sequence $H^{1}(Y/G,O)\to H^{1}(Y/G-P,O)\to H^{2}_{P}(Y/G,O)$. Since $Y/G$ is affine, we have $H^{1}(Y/G,O)=0$. So Cohen-Macaulayness of $Y/G$ would imply that $H^{1}(Y/G-P,O)=0$.
Fogarty showed that for $G=\text{\bf Z}/p$ acting with an isolated fixed point on a normal scheme $W$ over $\text{\bf F}_{p}$ of dimension at least 3, $W/G$ is not Cohen-Macaulay [12, Proposition 4]. When $W$ has mixed characteristic $(0,p)$, he needed $\dim(W)\geq 4$ to get the same conclusion. Nonetheless, we can build on his ideas to study the 3-dimensional scheme $Y$ in mixed characteristic.
We first show that $H^{1}(G,O(Y))$ is not zero. This cohomology group is $\ker(\operatorname{tr})/\operatorname{im}(1-\sigma)$ on $O(Y)$, where the trace is $1+\sigma$. Since $\operatorname{tr}(i)=0$, $i$ defines an element of $H^{1}(G,O(Y))$. Note that $i$ restricts to $1\in O(P)=\text{\bf F}_{2}$ on the fixed point $P$. Therefore, $i$ has nonzero image under the restriction map $H^{1}(G,O(Y))\to H^{1}(G,O(P))\cong\text{\bf F}_{2}$. So $i$ is nonzero in $H^{1}(G,O(Y))$, as we want.
Since $G$ acts freely on $Y$ outside $P$, we have a spectral sequence (as discussed in [12]):
$E_{2}^{pq}=H^{p}(G,H^{q}(Y-P,O))\Rightarrow H^{p+q}(Y/G-P,O).$
Here $H^{0}(Y-P,O)=O(Y-P)$ is equal to $O(Y)$, since $Y$ is normal and $P$ has codimension 3 in $Y$ (at least 2 would suffice). So $H^{1}(G,H^{0}(Y-P,O))=H^{1}(G,O(Y))\neq 0$. The spectral sequence shows that this group injects into $H^{1}(Y/G-P,O)$, and so $H^{1}(Y/G-P,O)\neq 0$. As discussed above, it follows that $Y/G$ is not Cohen-Macaulay.
It remains to show that $Y/G$ is terminal. Let us recall the definition. For a normal quasi-projective scheme $X$ over a regular base scheme $S$, Hartshorne defined the canonical sheaf $\omega_{X/S}$ [18, Definition 1.6]. It is a reflexive sheaf of rank 1, or equivalently the sheaf associated to a Weil divisor. In this paper, $S$ will be Spec of the $p$-adic integers or of a field, and we write $K_{X}$ for $\omega_{X/S}$. Toward the end of the proof of Theorem 5.1, we compute $K_{X}$ directly from the definition in our example.
A normal scheme $X$ is terminal if $K_{X}$ is Q-Cartier and, for every normal scheme $Z$ with a proper birational morphism $\pi\colon Z\to X$, we have
$K_{Z}=\pi^{*}(K_{X})+\sum_{j}a_{j}E_{j}$
with $a_{j}>0$ for every exceptional divisor $E_{j}$ of $\pi$. If $X$ has a resolution of singularities, terminality of $X$ is equivalent to positivity of the discrepancies $a_{j}$ on this one resolution [18, Corollary 2.12].
Let $Y_{0}=Y$. To prove that our example $Y_{0}/G$ is terminal, one approach would be to construct an explicit resolution of singularities. As with the analogous example in characteristic 2 [26, Theorem 5.1], this can be done by making $G$-equivariant blow-ups of $Y_{0}$ along regular closed subschemes. Namely, we can make $G$-equivariant blow-ups $Y_{2}\to Y_{1}\to Y_{0}$ such that $Y_{2}/G$ is regular. That resolution of $Y_{0}/G$ has dual complex a star, with one edge from a vertex $F_{0}$ to each of 7 other vertices $F_{1},\ldots,F_{7}$.
However, we can simplify the proof that $Y_{0}/G$ is terminal by stopping with a partial resolution with toric singularities. Namely, after only one blow-up $Y_{1}\to Y_{0}$, we can recognize the seven singularities of $Y_{1}/G$ as $\mu_{2}$-quotient singularities of the form $\frac{1}{2}(1,1,1)$, thanks to Theorem 2.2. So $Y_{1}/G$ is terminal, and it is then straighforward to compute that $Y_{0}/G$ is terminal. This method would also simplify the proof of terminality for the example in characteristic 2 that we are imitating [26]. The simplification is more striking for our more complicated examples in characteristic 3 or mixed characteristic $(0,3)$ (Theorems 6.1 and 7.1), and even more significant for our even more complicated examples in characteristic 5 or mixed characteristic $(0,5)$ (Theorems 8.1 and 9.1).
We now begin to blow up. To simplify the equations, change coordinates by $x_{0}:=x-1$, $x_{1}:=y-1$, and $e_{2}:=1+i$, so that the $G$-fixed point is defined by $0=x_{0}=x_{1}=e_{2}$. Then
$Y_{0}=\{(x_{0},x_{1},e_{2})\in A^{3}_{\text{\bf Z}_{2}}:0=e_{2}^{2}-2e_{2}+2,\;1+x_{0}\neq 0,\;1+x_{1}\neq 0\},$
and $G$ acts by
[…]
The blow-up at the $G$-fixed point is:
$Y_{1}=\{((x_{0},x_{1},e_{2}),[y_{0},y_{1},y_{2}])\in A^{3}_{\text{\bf Z}_{2}}\times_{\text{\bf Z}_{2}}\text{\bf P}^{2}_{\text{\bf Z}_{2}}:e_{2}^{2}-2e_{2}+2=0,\\ x_{0}y_{1}=x_{1}y_{0},\;x_{0}y_{2}=e_{2}y_{0},\;x_{1}y_{2}=e_{2}y_{1},\;1+x_{0}\neq 0,\;1+x_{1}\neq 0\}.$
The exceptional divisor $E_{0}\subset Y_{1}$ is isomorphic to $\text{\bf P}^{2}_{\text{\bf F}_{2}}$. Here $G$ acts on $Y_{1}$ by
$\sigma((x_{0},x_{1},e_{2}),[y_{0},y_{1},y_{2}])=\bigg{(}\bigg{(}\frac{-x_{0}}{1+x_{0}},\frac{-x_{1}}{1+x_{1}},2-e_{2}\bigg{)},\bigg{[}\frac{-y_{0}}{1+x_{0}},\frac{-y_{1}}{1+x_{1}},y_{2}(1-e_{2})\bigg{]}\bigg{)}.$
First consider the open subset $U_{0}=\{y_{0}=1\}$ in $Y_{1}$. Then $x_{1}=x_{0}y_{1}$ and $e_{2}=x_{0}y_{2}$, so
$U_{0}=\{(x_{0},y_{1},y_{2})\in A^{3}_{\text{\bf Z}_{2}}:0=(x_{0}y_{2})^{2}-2(x_{0}y_{2})+2,1+x_{0}\neq 0,1+x_{0}y_{1}\neq 0\}.$
Here $E_{0}=\{x_{0}=0\}$. The group $G=\text{\bf Z}/2$ acts by
$\sigma(x_{0},y_{1},y_{2})=\bigg{(}\frac{-x_{0}}{1+x_{0}},\;\frac{y_{1}(1+x_{0})}{1+x_{0}y_{1}},\;-y_{2}(1-x_{0}y_{2})(1+x_{0}).\bigg{)}.$
The fixed point scheme $Y_{1}^{G}$ is defined by: $I(x_{0})=\sigma(x_{0})-x_{0}=x_{0}^{2}(-1-x_{0}y_{2}^{2}+x_{0}^{2}y_{2}^{3})/(1+x_{0})$, $I(y_{1})=x_{0}y_{1}(1-y_{1})$, and $I(y_{2})=x_{0}y_{2}(-1-y_{2}+x_{0}y_{2}+x_{0}y_{2}^{2})$. We know that $Y_{1}^{G}$ (as a set) is contained in $E_{0}$. To focus on the fixed point scheme near $E_{0}$, we can say (more simply): $I(x_{0})=x_{0}^{2}(1+O(x_{0}))$, $I(y_{1})=x_{0}y_{1}(1+y_{1}+O(x_{0}))$, and $I(y_{2})=x_{0}y_{2}(1+y_{2}+O(x_{0}))$. We see that the fixed point scheme $Y_{1}^{G}$ is generically the Cartier divisor $E_{0}=\{x_{1}=0\}$. The bad locus in $E_{0}$ (where that fails) is given by removing a factor of $x_{0}$ from the equations and setting $x_{0}=0$, so we have: $0=x_{0}$, $0=y_{1}(1+y_{1})$, and $0=y_{2}(1+y_{2})$. (Note that $x_{1}=0$ implies $2=0$, by the equation for $U_{0}$.) So the fixed point scheme $Y_{1}^{G}$ (in this chart) is $E_{0}$ as a Cartier divisor except at the 4 points $(x_{0},y_{1},y_{2})$ equal to $(0,0,0)$, $(0,1,0)$, $(0,0,1)$, or $(0,1,1)$.
The open set $U_{1}=\{y_{1}=1\}$ works the same way, by the symmetry switching $x_{0}$ and $x_{1}$, hence also switching $y_{0}$ and $y_{1}$. Together, that gives 6 bad points in $E_{0}\cong\text{\bf P}^{2}_{\text{\bf F}_{2}}$ so far, namely $[y_{0},y_{1},y_{2}]$ equal to $[1,0,0]$, $[1,1,0]$, $[1,0,1]$, $[1,1,1]$, $[0,1,0]$, and $[0,1,1]$.
Finally, look at the open set $U_{2}=\{y_{2}=1\}$ in $Y_{1}$. Then $x_{0}=e_{2}y_{0}$, $x_{1}=e_{2}y_{1}$, and so
$U_{2}=\{(y_{0},y_{1},e_{2})\in A^{3}_{\text{\bf Z}_{2}}:e_{2}^{2}-2e_{2}+2=0,1+y_{0}e_{2}\neq 0,1+y_{1}e_{2}\neq 0\}.$
Here $E_{0}=\{e_{2}=0\}$. On $U_{2}$, $G$ acts by
[…]
We know that $Y_{1}^{G}$ is contained as a set in $E_{0}$. The fixed point scheme $Y_{1}^{G}$ (near $E_{0}$) is defined by: $I(y_{0})=y_{0}e_{2}(1+y_{0}+O(e_{2}))$, $I(y_{1})=y_{1}e_{2}(1+y_{1}+O(e_{2}))$, and $I(e_{2})=e_{2}^{2}(1+O(e_{2}))$. We see that the fixed point scheme $Y_{1}^{G}$ is generically $E_{0}=\{e_{2}=0\}$ as a Cartier divisor. The bad locus on $E_{0}$ (where this fails) is given by removing a factor of $e_{2}$ from the equations and setting $e_{2}=0$, so we have: $0=e_{2}$, $0=y_{1}(1+y_{1})$, and $0=y_{2}(1+y_{2})$. So there are 4 bad points on $E_{0}=\text{\bf P}^{2}_{\text{\bf F}_{2}}$ in this open set: $(y_{0},y_{1},e_{2})$ equal to $(0,0,0)$, $(1,0,0)$, $(0,1,0)$, or $(1,1,0)$.
We conclude that the fixed point scheme in $Y_{1}$ is $E_{0}\cong\text{\bf P}^{2}_{\text{\bf F}_{2}}$ with multiplicity 1 except at the 7 points:
$[y_{0},y_{1},y_{2}]=[1,0,0],[0,1,0],[0,0,1],[1,1,0],[1,0,1],[0,1,1],[1,1,1].$
(The same thing happens for the first blow-up of the analogous example in characteristic 2 [26].) By Theorem 2.1, $Y_{1}/G$ is regular outside the images of these 7 points.
[…]
Figure 1: The fixed point schemes in $Y_{0}$ and in the blow-up $Y_{1}$. Here $E_{0}\cong\text{\bf P}^{2}$, which we view as a toric variety; so the three edges of the triangle denote the coordinate lines in $\text{\bf P}^{2}$. We consider three coordinate charts on $Y_{1}$, each containing one vertex of the triangle.
One further $G$-equivariant blow-up at each of these 7 points suffices to resolve $Y_{1}/G$, but the equations for these blow-ups are a bit messy. Instead, we will use Theorem 2.2 to show that the 7 singular points of $Y_{1}/G$ are all mixed characteristic analogs of the singularity $A^{3}/\mu_{2}$, the simplest terminal singularity whose canonical class is not Cartier. More precisely, each of these 7 singular points is of the form $\frac{1}{2}(1,1,1)$, meaning that it can be written as $Q/\mu_{2}$ for some regular scheme $Q$ of dimension 3 (at an isolated fixed point of $\mu_{2}$). As a result, $Y_{1}/G$ is terminal. With one last calculation, we will deduce that $Y_{0}/G$ is terminal.
We first consider the singularities of $Y_{1}/G$ in the chart $U_{0}=\{y_{0}=1\}$, as above. Here $U_{0}$ has coordinates $(x_{0},y_{1},y_{2})$ with $20=(x_{0}y_{2})^{2}+2(x_{0}y_{2})-$ 2, and $E_{0}=\{x_{0}=0\}$. We saw that the fixed point scheme $U_{0}^{G}$ is the Cartier divisor $E_{0}=\{x_{0}=0\}$ except at the 4 points $(x_{0},y_{1},y_{2})$ equal to $(0,0,0),(0,0,1),(0,1,0)$, or $(0,1,1)$. At each of these points $P$, the $G$-action has the form required for Theorem 2.2 with $e=s:=x_{0}$, namely that $I(x_{0})=x_{0}^{2}(1+\operatorname{\mathfrak{m}})$ and $I(z_{j})=x_{0}(z_{j}+\operatorname{\mathfrak{m}}^{2})$ for $j=1,2$, for some coordinates $x_{0},z_{1},z_{2}$ at $P$, with $\operatorname{\mathfrak{m}}$ the maximal ideal at $P$. Also, since $x_{0}^{2}y_{2}^{2}=2(\text{unit})$, 2 is in the ideal $(x_{0}^{2})$, hence in $x_{0}\operatorname{\mathfrak{m}}$, which is another assumption in Theorem 2.2. So the theorem gives that these 4 singular points of $Y_{1}/G$ are of the form $\frac{1}{2}(1,1,1)$.
The calculations are identical in the chart $\{y_{1}=1\}$ in $Y_{1}$. They are slightly different in the chart $\{y_{2}=1\}$, but the conclusion is the same: the singularities of $Y_{1}/G$ in this chart are again of the form $\frac{1}{2}(1,1,1)$. Namely, this chart has coordinates $(y_{0},y_{1},e_{2})$ with $20=e_{2}^{2}-2e_{2}+$ 2, and $E_{0}=\{e_{2}=0\}$. The fixed point scheme $Y_{1}^{G}$ is the Cartier divisor $E_{0}$ except at the 4 points $(y_{0},y_{1},e_{2})$ equal to $(0,0,0),(1,0,0),(0,1,0)$, or $(1,1,0)$. At each of these points $P$, the $G$-action has the form required for Theorem 2.2 with $e=s:=e_{2}$, namely that $I(e_{2})=e_{2}^{2}(1+\operatorname{\mathfrak{m}})$ and $I(z_{j})=e_{2}(z_{j}+\operatorname{\mathfrak{m}}^{2})$ for $j=0,1$, for some coordinates $(z_{0},z_{1},e_{2})$ at $P$, where $\operatorname{\mathfrak{m}}$ is the maximal ideal at $P$. Also, the equation for $e_{2}$ implies that 2 is in the ideal $(e_{2}^{2})$, hence in $e_{2}\operatorname{\mathfrak{m}}$, which is another assumption in Theorem 2.2. So the theorem gives that these 4 singular points of $Y_{1}/G$ are again of the form $\frac{1}{2}(1,1,1)$.
Thus all 7 singular points of $Y_{1}/G$ are of the form $\frac{1}{2}(1,1,1)$. By the Reid-Tai criterion (Theorem 1.1), they are terminal. (To check that by hand: each singular point has a resolution $Z\to U_{1}/G$ whose exceptional divisor is $E_{j}\cong\text{\bf P}^{2}_{\text{\bf F}_{2}}$ with normal bundle $O(-2)$. As a result, the singularities of $Y_{1}/G$ are terminal, with $K_{Z}=\pi^{*}(K_{Y_{1}/G})+\frac{1}{2}E_{j}$ near each $E_{j}$.)
Recall that $Y_{0}=Y$ is the regular scheme of dimension 3 that we started with. (Thus $Y_{1}$ is the blow-up of $Y_{0}$ at the $G$-fixed point.) We now go on to show that $X=Y_{0}/G$ is terminal. Write $F_{0}$ for the image in $Y_{1}/G$ of the exceptional divisor $E_{0}$. Note that although $G$ fixes $E_{0}$ in $Y_{1}$, the morphism $E_{0}\rightarrow F_{0}$ is a finite purely inseparable morphism, not necessarily an isomorphism. (Indeed, $G=\text{\bf Z}/2$ is not linearly reductive over $\text{\bf Z}_{2}$. So if $G$ acts on an affine scheme $T$ preserving a closed subscheme $S$, the morphism $S/G\rightarrow T/G$ need not be a closed immersion. Equivalently, the $G$-equivariant surjection $O(T)\rightarrow O(S)$ need not yield a surjection $O(T)^{G}\rightarrow O(S)^{G}$.)
Write $K_{Y_{0}}$ for the canonical sheaf $\omega_{Y_{0}/\text{\bf Z}_{2}}$. Since $Y_{0}$ is regular, $K_{Y_{0}}$ is a line bundle, described as follows [18, Definition 1.6]. First, let $R=\text{\bf Z}_{2}[i]/(i^{2}+1)$. Then we have an embedding $D=\operatorname{Spec}R\subset A^{1}_{\text{\bf Z}_{2}}$; write $I$ for the ideal $(i^{2}+1)\subset\text{\bf Z}_{2}[i]$ defining this subscheme. Then the adjunction formula $K_{D}=(K_{A^{1}}+D)|_{D}$ is made into a definition:
$\omega_{R/\text{\bf Z}_{2}}=\Omega^{1}_{\text{\bf Z}_{2}[i]/\text{\bf Z}_{2}}\otimes_{\text{\bf Z}_{2}[i]}R\otimes_{R}(I/I^{2})^{*}.$
In these terms, one trivializing section of $\omega_{R/\text{\bf Z}_{2}}$ is $\alpha:=\frac{di}{i}\cdot f$, where $f\colon I/I^{2}\to R=\text{\bf Z}_{2}[i]/I$ is the map sending $i^{2}+1$ to 1. (Formally, one could think of this section of $\omega_{R/\text{\bf Z}_{2}}$ as $\frac{1}{i}\frac{di}{d(i^{2}+1)}$.) Next, since $\pi\colon Y_{0}\to\operatorname{Spec}R$ is smooth of relative dimension 2, we have $K_{Y_{0}}=\omega_{Y_{0}/\text{\bf Z}_{2}}=\Omega^{2}_{Y_{0}/R}\otimes\pi^{*}\omega_{R/\text{\bf Z}_{2}}$. So one trivializing section of $K_{Y_{0}}$ is $\beta:=\frac{dx}{x}\wedge\frac{dy}{y}\wedge\frac{di}{i}\cdot f$. I claim that this section is fixed by $G$. Indeed, if we extend the action of $G$ on $R$ to $A^{1}_{\text{\bf Z}_{2}}$ by $\sigma(i)=-i$, then $\sigma(f)=f$ and $\sigma(di)=-di$, so $\sigma(\frac{di}{i}f)=\frac{di}{i}f$. Also, $\sigma(dx/x)=-dx/x$ and $\sigma(dy/y)=-dy/y$, from which we see that $\sigma(\beta)=\beta$ as claimed. It follows that the divisor class $K_{Y_{0}/G}$ is linearly equivalent to zero, in particular Cartier. Here $K_{Y_{0}/G}$ is the canonical sheaf in the sense of [18, Definition 1.6]; $Y_{0}/G$ is not Gorenstein, since (as we have shown) it is not Cohen-Macaulay.
Since $K_{X}$ is Cartier, we can write
$K_{Y_{1}/G}=\pi^{*}K_{X}+a_{0}F_{0}$
for an integer $a_{0}$. Since $Y_{1}/G$ is terminal, $X$ is terminal if and only if the discrepancy $a_{0}$ is positive. Here and below, we write $\pi$ for all the relevant contractions, which in the formula above means $\pi\colon Y_{1}/G\rightarrow Y_{0}/G=X$.
The analogous formula for $Y_{1}$ is easy. Since $Y_{1}$ is the blow-up of the regular 3-dimensional scheme $Y_{0}$ at a closed point,
$K_{Y_{1}}=\pi^{*}K_{Y_{0}}+2E_{0}.$
Write $f$ for the quotient map $Y_{0}\rightarrow Y_{0}/G$ or $Y_{1}\rightarrow Y_{1}/G$. The ramification of $f$ along $E_{0}$ can be computed as follows.
Corollary 5.2.
Let $U$ and $G$ be as in Theorem 2.2. So $U$ is a regular scheme with an action of the group $G=\text{\bf Z}/p=\langle\sigma:\sigma^{p}=1\rangle$, for a prime number $p$, and assume that $U^{G}$ is generically a regular divisor $E_{1}=\{x_{1}=0\}$ and that $I(x_{1})=x_{1}^{2}(\text{unit})$. Assume that $U$ is of finite type over a regular base scheme $S$ and that $G$ acts on $U$ over $S$. Write $K_{U}$ and $K_{U/G}$ for the canonical classes over $S$. Then $U\to U/G$ is fiercely ramified along $E_{1}$. In particular, the image $F_{1}$ of $E_{1}$ in $U/G$ is Q-Cartier, and $f\colon U\to U/G$ satisfies $f^{*}F_{1}=E_{1}$ and $K_{U}=f^{*}K_{U/G}+(p-1)E_{1}$.
Proof.
The norm $N(x_{1})$ is a function on $U/G$ that defines a positive multiple of the divisor $F_{1}$, and so $F_{1}$ is Q-Cartier. Since the fixed point scheme $U^{G}$ is generically the Cartier divisor $E_{1}$, the ramification divisor of $f$ is $(p-1)E_{1}$, as mentioned above. That is, $K_{U}=f^{*}K_{U/G}+(p-1)E_{1}$. Also, the fixed point scheme is generically $E_{1}=\{x_{1}=0\}$ with coefficient 1, whereas $I(x_{1})$ vanishes to order 2 along $E_{1}$; so section 3 gives that the ramification of $f$ along $E_{1}$ is fierce. In particular, the ramification index $e$ is 1, meaning that $f^{*}F_{1}=E_{1}$. ∎
In particular, returning to our example with $p=2$, we have seen that the divisor $E_{0}$ has multiplicity 1 in the fixed point scheme $(Y_{1})^{G}$. Also, Corollary 5.2 gives that $f\colon Y_{1}\to Y_{1}/G$ is fiercely ramified along $E_{0}$. So we have
$K_{Y_{1}}=f^{*}K_{Y_{1}/G}+E_{0}$
and $f^{*}F_{0}=E_{0}$. (The same is true for the example in characteristic 2 that we are imitating [26].)
Since $f\colon Y_{0}\rightarrow Y_{0}/G$ is étale in codimension 1, we have $K_{Y_{0}}=f^{*}K_{Y_{0}/G}$. It follows that
$f^{*}K_{Y_{1}/G}$ $=K_{Y_{1}}-E_{0}$
$=(\pi^{*}K_{Y_{0}}+2E_{0})-E_{0}$
$=\pi^{*}f^{*}K_{Y_{0}/G}+E_{0}$
$=f^{*}(\pi^{*}K_{Y_{0}/G}+F_{0}).$
Therefore,
$K_{Y_{1}/G}=\pi^{*}K_{Y_{0}/G}+F_{0}.$
Because the coefficient of the exceptional divisor $F_{0}$ is positive, and $Y_{1}/G$ is terminal as shown above, $X=Y_{0}/G$ is terminal. ∎
6 Characteristic 3
Theorem 6.1.
Let the group $G=\text{\bf Z}/3$ with generator $\tau$ act on $\text{\bf P}^{2}$ over $\text{\bf F}_{3}$ by
$\tau([u_{0},u_{1},u_{2}])=[u_{1},u_{2},u_{0}]$
and on $\text{\bf P}^{1}$ by
$\tau([y_{0},y_{1}])=[y_{0},y_{0}+y_{1}].$
Then $(\text{\bf P}^{2}\times\text{\bf P}^{1})/G$ is terminal, not Cohen-Macaulay, and of dimension 3 over $\text{\bf F}_{3}$.
Proof.
We work throughout over $k=\text{\bf F}_{3}$. Write $G=\text{\bf Z}/3=\langle\sigma:\sigma^{3}=1\rangle$, with $\tau:=\sigma^{-1}$. Let $Y_{0}=\text{\bf P}^{2}\times\text{\bf P}^{1}$ and $X=Y_{0}/G$. The only fixed point of $G$ on $Y_{0}$ is $A=([1,1,1],[0,1])$. So $X$ is normal of dimension 3, and $X$ is smooth over $k$ outside the image $P$ of $A$. Also, $3K_{X}$ is Cartier. By Fogarty, since $A$ is an isolated fixed point of $G=\text{\bf Z}/p$ on a smooth 3-fold in characteristic $p$, $X$ is not Cohen-Macaulay at $P$ [12, Proposition 4].
It remains to show that $X$ is terminal. One can resolve the singularities of $X$ by performing $G$-equivariant blow-ups of $Y_{0}$. However, as in section 5, we will shorten the proof by recognizing that, after two $G$-equivariant blow-ups $Y_{2}\to Y_{1}\to Y_{0}$, the singularities of $Y_{2}/G$ become toric, namely quotients of a regular scheme by $\mu_{3}$. That makes it easy to check that $Y_{0}/G$ is terminal, without having to continue making $G$-equivariant blow-ups.
Before this approach, I found a $G$-equivariant blow-up $Y_{18}\to\cdots\to Y_{0}$ with $Y_{18}/G$ regular, but the construction involved 18 blow-ups along points or curves. The approach here, looking for toric singularities instead of regularity, saves a lot of work.
To put the fixed point at the origin, we change coordinates on $Y_{0}$ by: $x_{0}=(u_{0}+u_{1}+u_{2})/u_{1}$, $x_{1}=(-u_{1}+u_{2})/u_{1}$, and $x_{2}=y_{0}/y_{1}$. Then $G$ acts on the open subset $U$ of $Y_{0}$ given by
[…]
The $G$-action on $U$ is given by
$\tau(x_{0},x_{1},x_{2})=\bigg{(}\frac{x_{0}}{1+x_{1}},\;\frac{x_{1}+x_{0}}{1+x_{1}},\;\frac{x_{2}}{1+x_{2}}\bigg{)}.$
As we blow up, we will not need to keep track of the precise affine open set on which $G$ acts, since we are only concerned with the action near the fixed point set.
Let $Y_{1}$ be the blow-up of $Y_{0}$ at the $G$-fixed point, which is the origin in these coordinates. Then the open subset of $Y_{1}$ over $U\subset Y_{0}$ is
$\{((x_{0},x_{1},x_{2}),[y_{0},y_{1},y_{2}])\in U\times\text{\bf P}^{2}:x_{0}y_{1}=x_{1}y_{0},\;x_{0}y_{2}=x_{2}y_{0},x_{1}y_{2}=x_{2}y_{1}\}.$
Clearly the fixed point set $Y_{1}^{G}$ is contained in the exceptional divisor $E_{0}\cong\text{\bf P}^{2}$. It turns out to be a curve isomorphic to $\text{\bf P}^{1}$. We need three coordinate charts to cover $E_{0}$. First consider the open subset $\{y_{0}=1\}$ in $Y_{1}$. Here $(x_{0},x_{1},x_{2})=(x_{0},x_{0}y_{1},x_{0}y_{2})$, and $G$ acts by
[…]
By the action on the $y_{1}$ coordinate, there are no fixed points in this open set.
Next, work in the open set $\{y_{1}=1\}\subset Y_{1}$. Here $(x_{0},x_{1},x_{2})=(y_{0}x_{1},x_{1},y_{2}x_{1})$, $E_{0}=\{x_{1}=0\}$, and $G$ acts by
$\tau(y_{0},x_{1},y_{2})=\bigg{(}\frac{y_{0}}{1+y_{0}},\;\frac{x_{1}(1+y_{0})}{1+x_{1}},\;\frac{y_{2}(1+x_{1})}{(1+y_{0})(1+x_{1}y_{2})}\bigg{)}.$
The fixed point scheme $Y_{1}^{G}$ is defined by: $I(y_{0})=-y_{0}^{2}/(1+y_{0})$, $I(x_{1})=x_{1}(y_{0}+O(x_{1}))$, and $I(y_{2})=y_{2}(-y_{0}+O(x_{1}))/(1+y_{0}+O(x_{1}))$. Since we know that the fixed point set $Y_{1}^{G}$ is contained in $E_{0}=\{x_{1}=0\}$, we read off that the fixed point set is the line $\{0=y_{0}=x_{1}\}$ in $E_{0}$.
Finally, consider the open set $\{y_{2}=1\}$. Then $(x_{0},x_{1},x_{2})=(x_{2}y_{0},x_{2}y_{1},x_{2})$, and $G$ acts by
$\tau(y_{0},y_{1},x_{2})=\bigg{(}\frac{y_{0}(1+x_{2})}{1+y_{1}x_{2}},\;\frac{(y_{0}+y_{1})(1+x_{2})}{1+y_{1}x_{2}},\;\frac{x_{2}}{1+x_{2}}\bigg{)}.$
In these coordinates, the exceptional divisor $E_{0}\cong\text{\bf P}^{2}$ in $Y_{1}$ is $\{x_{2}=0\}$. The fixed point scheme of $G$ is defined by: $I(y_{0})=y_{0}x_{2}(1-y_{1}+O(x_{2}))$, $I(y_{1})=y_{0}+O(x_{2})$, and $I(x_{2})=x_{2}^{2}(-1+O(x_{2}))$. We know that the fixed point set is contained in $E_{0}$, and these equations imply that the fixed point set is the line $\{0=y_{0}=x_{2}\}$ in $E_{0}$, the same line as in the previous chart. Thus we have shown that the fixed point set in all of $Y_{1}$ is a curve isomorphic to $\text{\bf P}^{1}$.
Let $Y_{2}$ be the blow-up of $Y_{1}$ along the (reduced) $G$-fixed curve, with exceptional divisor $E_{1}\subset Y_{2}$. It is clear that $Y_{2}^{G}$ is contained in $E_{1}$ as a set. Write $E_{0}$ for the strict transform of $E_{0}\subset Y_{1}$ in $Y_{2}$. It turns out that the fixed point scheme $Y_{2}^{G}$ is equal to the Cartier divisor $E_{1}$ except at six points in $E_{1}$, three over the point $[y_{0},y_{1},y_{2}]=[0,1,0]$ in $E_{0}\subset Y_{1}$ and one each over $[y_{0},y_{1},y_{2}]$ equal to $[0,1,1]$, $[0,1,-1]$, or $[0,0,1]$. Since $E_{1}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}$, we will need to look at four affine charts to see all of it. (See Figure 2. Each affine chart we consider in $Y_{2}$ contains exactly one of the four vertices of the square.)
First work over the chart $\{y_{1}=1\}$ in $Y_{1}$. We wrote out the $G$-action on this chart above, with coordinates $(y_{0},x_{1},y_{2})$. We defined $Y_{2}$ as the blow-up of $Y_{1}$ along the $G$-fixed curve $\{0=y_{0}=x_{1}\}$. So, over this open subset of $Y_{1}$, $Y_{2}$ is an open subset of
$\{((y_{0},x_{1},y_{2}),[w_{0},w_{1}])\in A^{3}\times\text{\bf P}^{1}:y_{0}w_{1}=x_{1}w_{0}\}.$
[…]
Figure 2: For $j=0,1,2$, $G$ acts freely on $Y_{j}$ outside the shaded or marked loci; $Y_{j}/G$ is regular outside the marked loci; and $Y_{j}/G$ has toric singularities outside the red loci.
First look at the chart $\{w_{0}=1\}$ in $Y_{2}$ over $\{y_{1}=1\}$ in $Y_{1}$. (This chart contains the upper left vertex of $E_{1}$, in Figure 2.) Then $(y_{0},x_{1},y_{2})=(y_{0},y_{0}w_{1},y_{2})$, and $G=\text{\bf Z}/3$ acts by
$\tau(y_{0},w_{1},y_{2})=\bigg{(}\frac{y_{0}}{1+y_{0}},\;\frac{w_{1}(1+y_{0})^{2}}{1+y_{0}w_{1}},\;\frac{y_{2}(1+y_{0}w_{1})}{(1+y_{0}w_{1}y_{2})(1+y_{0})}\bigg{)}.$
Here $E_{0}=\{w_{1}=0\}$ and $E_{1}=\{y_{0}=0\}$. The fixed point scheme is defined by: $I(y_{0})=y_{0}^{2}(-1+O(y_{0}))$, $I(w_{1})=y_{0}w_{1}(-1-w_{1}+O(y_{0}))$, and $I(y_{2})=y_{0}y_{2}(-1+w_{1}-w_{1}y_{2}+O(y_{0}))$. So the fixed point scheme (near $E_{1}$) is $E_{1}$ with multiplicity 1 except at points with $0=y_{0}$, $0=w_{1}(w_{1}+1)$, and $0=y_{2}(1-w_{1}+w_{1}y_{2})$. Thus the bad points are $(y_{0},w_{1},y_{2})=(0,0,0)$ in $E_{0}\cap E_{1}$, $(0,-1,0)$, and $(0,-1,-1)$. We have found three bad points in $E_{1}$, with the first one in $E_{0}\cap E_{1}$.
The other chart over $\{y_{1}=1\}$ in $Y_{1}$ is $\{w_{1}=1\}$ in $Y_{2}$ (the chart containing the upper right vertex of $E_{1}$ in Figure 2). Then $(y_{0},x_{1},y_{2})=(w_{0}x_{1},x_{1},y_{2})$, and $G$ acts by
$\tau(w_{0},x_{1},y_{2})=\bigg{(}\frac{w_{0}(1+x_{1})}{(1+w_{0}x_{1})^{2}},\;\frac{x_{1}(1+w_{0}x_{1})}{1+x_{1}},\;\frac{y_{2}(1+x_{1})}{(1+x_{1}y_{2})(1+w_{0}x_{1})}\bigg{)}.$
Here $E_{0}$ does not appear, and $E_{1}=\{x_{1}=0\}$. The fixed point scheme is defined (near $E_{1}$) by: $I(w_{0})=w_{0}x_{1}(1+w_{0}+O(x_{1}))$, $I(x_{1})=x_{1}^{2}(-1+w_{0}+O(x_{1}))$, and $I(y_{2})=x_{1}y_{2}(1-w_{0}-y_{2}+O(x_{1}))$. So $Y_{2}^{G}$ is $E_{1}$ with multiplicity 1 except at points where $0=x_{1}$, $0=w_{0}(1+w_{0})$, and $0=y_{2}(1-w_{0}-y_{2})$. Thus the bad points in this chart are $(w_{0},x_{1},y_{2})$ equal to $(0,0,0)$, $(0,0,1)$, $(-1,0,0)$, and $(-1,0,-1)$. The points with $w_{0}\neq 0$ appeared in the previous chart, $\{w_{0}=1\}$. So we have two new bad points, $(w_{0},x_{1},y_{2})$ equal to $(0,0,0)$ or $(0,0,1)$, for a total of five so far.
To see all of $E_{1}$ in $Y_{2}$, we also need to work over $\{y_{2}=1\}\subset Y_{1}$, the chart containing the lower left vertex of $E_{1}$ in Figure 2. The corresponding open subset of $Y_{2}$ is an open subset of
$\{((y_{0},y_{1},x_{2}),[z_{0},z_{2}])\in A^{3}\times\text{\bf P}^{1}:y_{0}z_{2}=x_{2}z_{0}\}.$
First consider the chart $\{z_{0}=1\}$ in $Y_{2}$. Then $(y_{0},y_{1},x_{2})=(y_{0},y_{1},y_{0}z_{2})$, and $G$ acts by
$\tau(y_{0},y_{1},z_{2})=\bigg{(}\frac{y_{0}(1+y_{0}z_{2})}{1+y_{0}y_{1}z_{2}},\;\frac{(y_{0}+y_{1})(1+y_{0}z_{2})}{1+y_{0}y_{1}z_{2}},\;\frac{z_{2}(1+y_{0}y_{1}z_{2})}{(1+y_{0}z_{2})^{2}}\bigg{)}.$
Here $E_{0}=\{z_{2}=0\}$ and $E_{1}=\{y_{0}=0\}$. The fixed point scheme $Y_{2}^{G}$ is defined (near $E_{1}$) by: $I(y_{0})=y_{0}^{2}z_{2}(1-y_{1}+O(y_{0}))$, $I(y_{1})=y_{0}(1+y_{1}z_{2}-y_{1}^{2}z_{2}+O(y_{0}))$, and $I(z_{2})=y_{0}z_{2}^{2}(1+y_{1}+O(y_{0}))$. This is equal to $E_{1}$ with multiplicity 1 except at the point $(y_{0},y_{1},z_{2})=(0,-1,-1)$. This appeared in the chart $\{w_{0}=1\}$ over $\{y_{1}=1\}$ (as the point $(y_{0},w_{1},y_{2})=(0,-1,-1)$).
The other chart is $\{z_{2}=1\}$ in $Y_{2}$, which contains the lower right vertex of $E_{1}$ in Figure 2. Then $(y_{0},y_{1},x_{2})=(z_{0}x_{2},y_{1},x_{2})$, and $G$ acts by
$\tau(z_{0},y_{1},x_{2})=\bigg{(}\frac{z_{0}(1+x_{2})^{2}}{1+y_{1}x_{2}},\;\frac{(y_{1}+z_{0}x_{2})(1+x_{2})}{1+y_{1}x_{2}},\;\frac{x_{2}}{1+x_{2}}\bigg{)}.$
In this open set, $E_{0}$ does not appear, and $E_{1}=\{x_{2}=0\}$. The fixed point scheme $Y_{2}^{G}$ is defined by: $I(z_{0})=z_{0}x_{2}(-1-y_{1}+O(x_{2}))$, $I(y_{1})=x_{2}(z_{0}+y_{1}-y_{1}^{2}+O(x_{2}))$, and $I(x_{2})=x_{2}^{2}(-1+O(x_{2}))$. We know that $Y_{2}^{G}$ is contained in $E_{1}=\{x_{2}=0\}$ as a set. We read off that the fixed point scheme $Y_{2}^{G}$ is generically the Cartier divisor $E_{1}$. The bad locus (where that fails) is given, in $E_{1}$, by: $0=x_{2}$, $0=z_{0}(1+y_{1})$, and $0=z_{0}+y_{1}-y_{1}^{2}$. So we see three bad points in $E_{1}$: $(z_{0},y_{1},x_{2})=(0,1,0)$, $(-1,-1,0)$, or $(0,0,0)$. The first two points appeared in the chart $\{w_{1}=1\}$ in $Y_{2}$ over $\{y_{1}=1\}$ in $Y_{1}$ (as $(w_{0},x_{1},y_{2})=(0,0,1)$ or $(-1,0,-1)$, respectively). On the other hand, the origin in this chart is new; so we have a total of six bad points.
Thus the fixed point scheme $Y_{2}^{G}$ is $E_{1}$ with multiplicity 1 except at six points on $E_{1}$. We will use Theorem 2.2 to recognize the resulting six singularities of $Y_{2}/G$ as toric; so we have no further need for $G$-equivariant blow-ups.
First consider three of the six bad points of $E_{1}\subset Y_{2}$, the ones with $[y_{0},y_{1},y_{2}]$ equal to $[0,0,1]$, $[0,1,1]$, or $[0,1,-1]$. In these cases, our calculation of the action of $G$ shows (by Theorem 2.2) that the singularity of $Y_{2}/G$ at these points is of the form $\frac{1}{3}(1,1,2)$. (For example, for the point $[y_{0},y_{1},y_{2}]=[0,0,1]$, work in the chart $\{z_{2}=1\}$ in $Y_{2}$ over $\{y_{2}=1\}$ in $Y_{1}$. Here we have coordinates $(z_{0},y_{1},x_{2})$, $(Y_{2})^{G}_{\operatorname{red}}$ is $\{x_{2}=0\}$, and the point we are considering is the origin. As above, we have $I(z_{0})=z_{0}x_{2}(-1-y_{1}+O(x_{2}))$, $I(y_{1})=x_{2}(z_{0}+y_{1}-y_{1}^{2}+O(x_{2}))$, and $I(x_{2})=x_{2}^{2}(-1+O(x_{2}))$. So Theorem 2.2 applies, with $e=s:=x_{2}$. The linear map $\varphi$ in the theorem is given by: $\varphi(z_{0})=-z_{0}$, $\varphi(y_{1})=z_{0}+y_{1}$, and $\varphi(x_{2})=-x_{2}$. This map has eigenvalues $-1,1,-1$ in $\text{\bf F}_{3}$, and so Theorem 2.2 gives that the singularity of $Y_{2}/G$ at this point is of the form $\frac{1}{3}(-1,1,-1)\cong\frac{1}{3}(1,1,2)$.) By the Reid-Tai criterion (Theorem 1.1), this singularity is terminal.
Next, consider the chart $\{w_{1}=1\}$ in $Y_{2}$ over $\{y_{1}=1\}$ in $Y_{1}$. As shown above, in coordinates $(w_{0},x_{1},y_{2})$, the points $(0,0,0)$ and $(-1,0,0)$ are bad for the action of $G$ on $Y_{2}$, and $(Y_{2})_{\operatorname{red}}=\{x_{1}=0\}$. Here again, the singularity of $Y_{2}/G$ at these points is of the form $\frac{1}{3}(1,1,2)$. (For example, for the point $(w_{0},x_{1},y_{2})=(0,0,0)$, we saw that $I(w_{0})=w_{0}x_{1}(1+w_{0}+O(x_{1}))$, $I(x_{1})=x_{1}^{2}(-1+w_{0}+O(x_{1}))$, and $I(y_{2})=x_{1}y_{2}(1-w_{0}-y_{2}+O(x_{1}))$. So Theorem 2.2 applies with $e=s:=x_{1}$. The linear map $\varphi$ in the theorem is given by $\varphi(w_{0})=w_{0}$, $\varphi(x_{1})=-x_{1}$, and $\varphi(y_{2})=y_{2}$. This has eigenvalues $1,-1,1$, and so the theorem gives that the singularity of $Y_{2}/G$ at this point is of the form $\frac{1}{3}(1,-1,1)\cong\frac{1}{3}(1,1,2)$.) In particular, these two points are terminal. Thus $Y_{2}/G$ is terminal at five of its six singular points.
Finally, we consider the last singular point of $Y_{2}/G$. In the chart $\{w_{0}=1\}$ in $Y_{2}$ over $\{y_{1}=1\}$ in $Y_{1}$, the point $P$ is $(y_{0},w_{1},y_{2})=(0,0,0)$, and $(Y_{2})^{G}_{\operatorname{red}}=\{y_{0}=0\}$. In this case, our calculation of the action of $G$ shows (by Theorem 2.2) that the singularity of $Y_{2}/G$ is of the form $\frac{1}{3}(1,1,1)$. (As above, we have $I(y_{0})=y_{0}^{2}(-1+O(y_{0}))$, $I(w_{1})=y_{0}w_{1}(-1-w_{1}+O(y_{0}))$, and $I(y_{2})=y_{0}y_{2}(-1+w_{1}-w_{1}y_{2}+O(y_{0}))$. So Theorem 2.2 applies with $e=s:=y_{0}$. The linear map $\varphi$ in the theorem is given by $\varphi(y_{0})=-y_{0}$, $\varphi(w_{1})=-w_{1}$, and $\varphi(y_{2})=-y_{2}$. So Theorem 2.2 gives that the singularity of $Y_{2}/G$ at this point is of the form $\frac{1}{3}(-1,-1,-1)\cong\frac{1}{3}(1,1,1)$.) By the Reid-Tai criterion (Theorem 1.1), $Y_{2}/G$ is canonical at this point, but not terminal.
We can now begin the proof that $Y_{0}/G$ is terminal. Write $f$ for any of the quotient maps $Y_{j}\to Y_{j}/G$. The fixed point scheme $Y_{2}^{G}$ is the Cartier divisor $E_{1}$ except at six points on $E_{1}$. Clearly $3K_{Y_{2}/G}$ is Cartier. Write $E_{0}\subset Y_{2}$ for the strict transform of $E_{0}\subset Y_{1}$. For $j=0,1$ in $Y_{2}$, let $F_{j}$ be the image (as an irreducible divisor) of $E_{j}$ in $Y_{2}/G$. Since $G$ acts nontrivially on $E_{0}$, we have $f^{*}F_{0}=E_{0}$. The divisor $E_{1}$ is fixed by $G$. By Corollary 5.2, $f$ is fiercely ramified along $E_{1}$. So the ramification divisor is $p-1$ times $E_{1}$, meaning that $K_{Y_{2}}=f^{*}(K_{Y_{2}/G})+2E_{1}$, and we have $f^{*}F_{1}=E_{1}$.
Write $\pi_{ij}$ for the birational morphism $Y_{i}\to Y_{j}$ or $Y_{i}/G\to Y_{j}/G$ (with $i>j$). Since $\pi_{20}\colon Y_{2}\to Y_{0}$ is defined by blowing up points and smooth curves on a smooth 3-fold, we have:
$K_{Y_{1}}=\pi_{10}^{*}(K_{Y_{0}})+2E_{0}$
and
$K_{Y_{2}}$ $=\pi_{21}^{*}(K_{Y_{1}})+E_{1}$
$=\pi_{20}^{*}(K_{Y_{0}})+2(E_{0}+E_{1})+E_{1}$
$=\pi_{20}^{*}(K_{Y_{0}})+2E_{0}+3E_{1}.$
Therefore,
$f^{*}\pi_{20}^{*}K_{Y_{0}/G}={}$ $\pi_{20}^{*}f^{*}K_{Y_{0}/G}$
[…]
So
$K_{Y_{2}/G}=\pi_{20}^{*}(K_{Y_{0}/G})+2F_{0}+F_{1}.$
Here every exceptional divisor of the birational morphism $Y_{2}/G\to Y_{0}/G$ has positive coefficient. Also, we have shown that $Y_{2}/G$ is terminal outside one point which is canonical, and that point lies on $F_{1}$. Therefore, $Y_{0}/G$ is terminal. We showed earlier that it is not Cohen-Macaulay. Theorem 6.1 is proved. ∎
Remark 6.2.
The divisor class $\pi^{*}K_{Y_{0}/G}=K_{Y_{2}/G}-2F_{0}-F_{1}$ has some non-integer discrepancies, for example over the origin in the chart $\{z_{2}=1\}$ in $Y_{2}$ over $\{y_{2}=1\}$ in $Y_{1}$. As a result, $K_{Y_{0}/G}$ is not Cartier (as one can also check directly), in contrast to our examples in residue characteristic 2 ([26, Theorem 5.1] and Theorem 5.1). I expect that there is also a 3-fold $X$ over $\text{\bf F}_{3}$ that is terminal and non-Cohen-Macaulay with $K_{X}$ Cartier. Namely, one should replace $\text{\bf P}^{1}$ in Theorem 6.1 by the “Harbater-Katz-Gabber” curve $C$$\{0=y^{p}-yz^{p-1}-x^{p-1}z\}$ in $\text{\bf P}^{2}_{\text{\bf F}_{p}}$ [8], here with $p=3$, which has a $\text{\bf Z}/p$-action by $\tau([x,y,z])=[x+z,y,z]$ that preserves a nonzero 1-form near the fixed point $[1,0,0]$. For $p=3$, $C$ is a supersingular elliptic curve.
7 The example over the 3-adic integers
Theorem 7.1.
Let $G$ be the group $G=\text{\bf Z}/3$ with generator $\tau$. Let $R=\text{\bf Z}_{3}[e]/(e^{3}-3e^{2}+3)$, which is the ring of integers in a Galois extension of $\text{\bf Q}_{3}$ with group $G=\text{\bf Z}/3$. Let $G$ act on the scheme $\text{\bf P}^{2}_{R}$ by
$\tau([u_{0},u_{1},u_{2}],e)=([u_{1},u_{2},u_{0}],3+e-e^{2}).$
Then the scheme $\text{\bf P}^{2}_{R}/G$ is terminal, not Cohen-Macaulay, of dimension 3, and flat over $\text{\bf Z}_{3}$.
This example behaves much like the example over $\text{\bf F}_{3}$, Theorem 6.1. Theorem 8.1. In particular, Figure 2 accurately depicts the blow-ups we make in mixed characteristic $(0,3)$, just as in characteristic 3. We can view $R$ as the subring $\text{\bf Z}_{3}[\zeta_{9}]^{+}$ of the cyclotomic ring $\text{\bf Z}_{3}[\zeta_{9}]$ fixed by $\zeta_{9}\mapsto\zeta_{9}^{-1}$, with $e=\zeta_{9}+1+\zeta_{9}^{-1}$. Informally, $R$ is the simplest ramified $\text{\bf Z}/3$-extension of $\text{\bf Z}_{3}$. More broadly, this action of $G$ on $\text{\bf P}^{2}_{R}$ was chosen as possibly the simplest action of $\text{\bf Z}/3$ on a 3-fold in mixed characteristic $(0,3)$ with an isolated fixed point. The simplicity helps to ensure that the quotient scheme is terminal.
Proof.
We write $G=\text{\bf Z}/3=\langle\sigma:\sigma^{3}=1\rangle$, with $\tau:=\sigma^{-1}$. Let $Y_{0}=\text{\bf P}^{2}_{R}$ with $G$ acting diagonally on $\text{\bf P}^{2}$ and on $R$, and let $X=Y_{0}/G$. Write $e_{2}$ for the generator $e$ of $R$, to fit better with our numbering of coordinates on $Y_{0}$; so we have
$30=e_{2}^{3}-3e_{2}^{2}+$ 3 .
The only fixed point of $G$ on $Y_{0}$ is the closed point $P\cong\operatorname{Spec}\text{\bf F}_{3}$ given by $([u_{0},u_{1},u_{2}],e_{2})=([1,1,1],0)$. So $X$ is normal of dimension 3, and $X$ is regular outside the image of $P$, which we also call $P$. Clearly $3K_{X}$ is Cartier.
It is not automatic from Fogarty’s results [12], but we can use his methods to show that $X$ is not Cohen-Macaulay at $P$. As in the proof of Theorem 5, using that $G$ has an isolated fixed point on the 3-fold $Y_{0}$, it suffices to show that $H^{1}(G,O(Y_{0}))$ is not zero. This cohomology group is $\ker(\operatorname{tr})/\operatorname{im}(1-\sigma)$ on $O(Y_{0})$, where the trace is $1+\sigma+\sigma^{2}$. The equation $30=e_{2}^{3}-3e_{2}^{2}+$ 3 (specifically, the coefficient of $e_{2}^{2}$) implies that $e_{2}$ has trace 3. So $\operatorname{tr}(1-e_{2})=0$, and hence $1-e_{2}$ defines an element of $H^{1}(G,O(Y_{0}))$. Note that $1-e_{2}$ restricts to $1\in O(P)=\text{\bf F}_{3}$ on the fixed point $P$. Therefore, $1-e_{2}$ has nonzero image under the restriction map $H^{1}(G,O(Y_{0}))\to H^{1}(G,O(P))\cong\text{\bf F}_{3}$. So $H^{1}(G,O(Y_{0}))$ is not zero, and hence $Y_{0}/G$ is not Cohen-Macaulay.
It remains to show that $X$ is terminal. One can resolve the singularities of $X$ by performing $G$-equivariant blow-ups of $Y_{0}$. However, as in sections 5 and 6, we will shorten the proof by recognizing that, after two $G$-equivariant blow-ups $Y_{2}\to Y_{1}\to Y_{0}$, the singularities of $Y_{2}/G$ become toric (in Kato’s mixed characteristic sense), namely quotients of a regular scheme by $\mu_{3}$. That makes it easy to check that $Y_{0}/G$ is terminal, without having to continue making $G$-equivariant blow-ups.
To put the fixed point at the origin, we change coordinates on $Y_{0}$ by: $x_{0}=(u_{0}+u_{1}+u_{2}-3)/u_{1}$ and $x_{1}=(-u_{1}+u_{2})/u_{1}$. Then $G$ acts on
$U:=\{(x_{0},x_{1},e_{2})\in A^{3}_{\text{\bf Z}_{3}}:0=3-3e_{2}^{2}+e_{2}^{3},1+x_{1}\neq 0,1+x_{0}-x_{1}\neq 0\}$
by
[…]
In what follows, we will often not need to keep track of the precise open set on which $G$ acts, because we are only concerned with the $G$-fixed point scheme.
The blow-up $Y_{1}\to Y_{0}$ at the $G$-fixed point is, over the open set $U\subset Y_{0}$:
$\{((x_{0},x_{1},e_{2}),[y_{0},y_{1},y_{2}])\in U\times_{\text{\bf Z}_{3}}\text{\bf P}^{2}_{\text{\bf Z}_{3}}:x_{0}y_{1}=x_{1}y_{0},\;x_{0}y_{2}=e_{2}y_{0},\;x_{1}y_{2}=e_{2}y_{1}\}.$
Clearly the fixed point set $Y_{1}^{G}$ is contained in the exceptional divisor $E_{0}\cong\text{\bf P}^{2}_{\text{\bf F}_{3}}$. It turns out to be a curve isomorphic to $\text{\bf P}^{1}_{\text{\bf F}_{3}}$. We need three coordinate charts to cover $E_{0}$. First look at the open set $U_{0}=\{y_{0}=1\}$ in $Y_{1}$. Then $(x_{0},x_{1},e_{2})=(x_{0},x_{0}y_{1},x_{0}y_{2})$, and so
$U_{0}=\{(x_{0},y_{1},y_{2})\in A^{3}_{\text{\bf Z}_{3}}:0=x_{0}^{3}y_{2}^{3}-3x_{0}^{2}y_{2}^{2}+3,\;1+x_{0}y_{1}\neq 0,\;1+x_{0}-x_{0}y_{1}\neq 0\}.$
The exceptional divisor $E_{0}$ is $\{x_{0}=0\}$ (isomorphic to $A^{2}_{\text{\bf F}_{3}}$, in this open set). Here $G$ acts by
$\tau(x_{0},y_{1},y_{2})=\bigg{(}\frac{x_{0}(1-3y_{1})}{1+x_{0}y_{1}},\;\frac{1-2y_{1}}{1-3y_{1}},\;\frac{y_{2}(1+2x_{0}y_{2}-x_{0}^{2}y_{2}^{2})(1+x_{0}y_{1})}{1-3y_{1}}\bigg{)}.$
We know that the fixed point scheme $Y_{1}^{G}$ is contained in $E_{0}=\{x_{0}=0\}$ as a set, and that $3=0$ on $E_{0}$. The fixed point scheme is defined by: $I(x_{0})=x_{0}^{2}y_{1}(-1+O(x_{0}))$, $I(y_{1})=1+O(x_{0})$, and $I(y_{2})=x_{0}y_{2}(y_{1}-y_{2}+O(x_{0}))$. By the second equation, $Y_{1}^{G}$ is empty in this open set.
Next, consider the open set $U_{1}=\{y_{1}=1\}$ in $Y_{1}$. Then $(x_{0},x_{1},e_{2})=(x_{1}y_{0},x_{1},x_{1}y_{2})$, and $G$ acts on
$U_{1}=\{(y_{0},x_{1},y_{2})\in A^{3}_{\text{\bf Z}_{3}}:0=3-3x_{1}^{2}y_{2}^{2}+x_{1}^{3}y_{2}^{3},\;1+x_{1}\neq 0,\;1-x_{1}+x_{1}y_{2}\neq 0\}.$
Namely, $G$ acts by
$\tau(y_{0},x_{1},y_{2})=\bigg{(}\frac{y_{0}-3}{y_{0}-2},\;\frac{x_{1}(y_{0}-2)}{1+x_{1}},\;\frac{y_{2}(1+x_{1})(1+2x_{1}y_{2}-x_{1}^{2}y_{2}^{2})}{y_{0}-2}\bigg{)}.$
Here $E_{0}=\{x_{1}=0\}$. We know that $Y_{1}^{G}$ is contained in $E_{0}=\{x_{1}=0\}$, as a set (and hence $3=0$ on $(Y_{1})^{G}_{\operatorname{red}}$). More precisely, the fixed point scheme $Y_{1}^{G}$ is defined by: $I(y_{0})=(-y_{0}^{2}+O(x_{1}))/(1+y_{0}+O(x_{1}))$, $I(x_{1})=x_{1}(y_{0}+O(x_{1}))$, and $I(y_{2})=y_{2}(-y_{0}+O(x_{1}))/(1+y_{0}+O(x_{1}))$. So $Y_{1}^{G}$ is the line $\{y_{0}=x_{1}=0\}$, as a set.
Finally, consider the chart $\{y_{2}=1\}$ in $Y_{1}$. Here $(x_{0},x_{1},e_{2})=(e_{2}y_{0},e_{2}y_{1},e_{2})$, and $G$ acts on the open set
$U_{2}=\{(y_{0},y_{1},e_{2})\in A^{3}_{\text{\bf Z}_{3}}:0=3-3e_{2}^{2}+e_{2}^{3},\;1+e_{2}y_{0}\neq 0,\;1+e_{2}y_{0}-e_{2}y_{1}\neq 0\}.$
Namely, $G$ acts by
[…]
The fixed point scheme $Y_{1}^{G}$ is defined by: $I(y_{0})=e_{2}(y_{0}-y_{0}y_{1}+O(e_{2}))$, $I(y_{1})=y_{0}+O(e_{2})$, and $I(e_{2})=e_{2}^{2}(-1+O(e_{2}))$. We know that the fixed point set is contained in $E_{0}=\{e_{2}=0\}$. We read off that the fixed point set is the line $0=y_{0}=e_{2}$, which is the same line seen in the previous chart. Thus we have shown that the fixed point set in all of $Y_{1}$ is a curve isomorphic to $\text{\bf P}^{1}_{\text{\bf F}_{3}}$ in $E_{0}$.
Seeking to make the fixed point set a divisor, we let $Y_{2}$ be the blow-up of $Y_{1}$ along the (reduced) $G$-fixed curve, with exceptional divisor $E_{1}\subset Y_{2}$. It is clear that $Y_{2}^{G}$ is contained in $E_{1}$ as a set. We will see that the fixed point scheme $Y_{2}^{G}$ is equal to the Cartier divisor $E_{1}$ except at six points on $E_{1}$. These correspond exactly to the six bad points that occur in the example over $\text{\bf F}_{3}$ (Figure 2). Since $E_{1}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}_{\text{\bf F}_{3}}$, we will need to look at four affine charts to see all of it.
First work over the open subset $\{y_{1}=1\}$ in $Y_{1}$, with coordinates $(y_{0},x_{1},y_{2})$, where $E_{0}=\{x_{1}=0\}$. Since $Y_{2}$ is the blow-up of $Y_{1}$ along the $G$-fixed curve $\{0=y_{0}=x_{1}\}$, this part of $Y_{2}$ is given by
$\{((y_{0},x_{1},y_{2}),[w_{0},w_{1}])\in A^{3}_{\text{\bf Z}_{3}}\times\text{\bf P}^{1}_{\text{\bf Z}_{3}}:0=3-3x_{1}^{2}y_{2}^{2}+x_{1}^{3}y_{2}^{3},\;y_{0}w_{1}=x_{1}w_{0}\}.$
First consider the open set $\{w_{0}=1\}\subset Y_{2}$ over $\{y_{1}=1\}\subset Y_{1}$. Then $(y_{0},x_{1},y_{2})=(y_{0},y_{0}w_{1},y_{2})$, $E_{0}=\{w_{1}=0\}$, and $E_{1}=\{y_{0}=0\}$. Also, $e_{2}=y_{0}w_{1}y_{2}$, and so $3=3y_{0}^{2}w_{1}^{2}y_{2}^{2}-y_{0}^{3}w_{1}^{3}y_{2}^{3}$. Here $G$ acts by
$\tau(y_{0},w_{1},y_{2})=\bigg{(}\frac{y_{0}-3}{y_{0}-2},\;\frac{w_{1}(y_{0}-2)^{2}}{(1+y_{0}w_{1})(1-3y_{0}w_{1}^{2}y_{2}^{2}+y_{0}^{2}w_{1}^{3}y_{2}^{3})},\\ \frac{y_{2}(1+y_{0}w_{1})(1+2y_{0}w_{1}y_{2}-y_{0}^{2}w_{1}^{2}y_{2}^{2})}{y_{0}-2}\bigg{)}.$
The fixed point scheme $Y_{2}^{G}$ (near $E_{1}$) is defined by: $I(y_{0})=y_{0}^{2}(-1+O(y_{0}))$, $I(w_{1})=y_{0}w_{1}(-1-w_{1}+O(y_{0}))$, and $I(y_{2})=y_{0}y_{2}(-1+w_{1}-w_{1}y_{2}+O(y_{0}))$. So $Y_{2}^{G}$ is the Cartier divisor $E_{1}$ except where $0=y_{2}(1-w_{1}+w_{1}y_{2})$ and $0=w_{1}(1+w_{1})$. So we have found three bad points, $(y_{0},w_{1},y_{2})=(0,0,0)$ in $E_{0}\cap E_{1}$ and $(0,-1,0)$, and $(0,-1,-1)$ in $E_{1}$.
The other chart over $\{y_{1}=1\}$ in $Y_{1}$ is $\{w_{1}=1\}$ in $Y_{2}$. Then $(y_{0},x_{1},y_{2})=(w_{0}x_{1},x_{1},y_{2})$, $E_{0}$ does not appear, and $E_{1}=\{x_{1}=0\}$. Also, $e_{2}=x_{1}y_{2}$, and so $3=3x_{1}^{2}y_{2}^{2}-x_{1}^{3}y_{2}^{3}$. Here $G$ acts by
$\tau(w_{0},x_{1},y_{2})=\bigg{(}\frac{(w_{0}-3x_{1}y_{2}^{2}+x_{1}^{2}y_{2}^{3})(1+x_{1})}{(w_{0}x_{1}-2)^{2}},\;\frac{x_{1}(w_{0}x_{1}-2)}{1+x_{1}},\\ \frac{y_{2}(1+x_{1})(1+2x_{1}y_{2}-x_{1}^{2}y_{2}^{2})}{w_{0}x_{1}-2}\bigg{)}.$
So the fixed point scheme $Y_{2}^{G}$ (near $E_{1}$) is defined by: $I(w_{0})=x_{1}(w_{0}+w_{0}^{2}+O(x_{1}))$, $I(x_{1})=x_{1}^{2}(-1+w_{0}+O(x_{1}))$, and $I(y_{2})=x_{1}y_{2}(1-w_{0}-y_{2}+O(x_{1}))$. So $Y_{2}^{G}$ is the Cartier divisor $E_{1}$ except where $x_{1}=0$ (so $3=0$), $0=w_{0}(1+w_{0})$, and $0=y_{2}(1-w_{0}-y_{2})$. So the bad points are $(w_{0},x_{1},y_{2})=(0,0,0)$, $(0,0,1)$, $(-1,0,0)$, and $(-1,0,-1)$. The points with $w_{0}\neq 0$ appeared in the previous chart, $\{w_{0}=1\}$. So we have two new bad points, $(w_{0},x_{1},y_{2})$ equal to $(0,0,0)$ or $(0,0,1)$, for a total of five so far.
To see all of $E_{1}$ in $Y_{2}$, we also have to work over the open set $\{y_{2}=1\}\subset Y_{1}$, with coordinates $(y_{0},y_{1},x_{2})$, where $E_{0}=\{x_{2}=0\}$. The corresponding open subset of $Y_{2}$ is an open subset of
$\{((y_{0},y_{1},e_{2}),[z_{0},z_{2}])\in A^{3}_{\text{\bf Z}_{3}}\times_{\text{\bf Z}_{3}}\text{\bf P}^{1}_{\text{\bf Z}_{3}}:0=3-3e_{2}^{2}+e_{2}^{3},\;y_{0}z_{2}=e_{2}z_{0}\}.$
First consider the chart $\{z_{0}=1\}\subset Y_{2}$. Then $(y_{0},y_{1},e_{2})=(y_{0},y_{1},y_{0}z_{2})$, $E_{0}=\{z_{2}=0\}$, and $E_{1}=\{y_{1}=0\}$. Also, $e_{2}=y_{0}z_{2}$, and so $0=3-3y_{0}^{2}z_{2}^{2}+y_{0}^{3}z_{2}^{3}$. Here $G$ acts by
$\tau(y_{0},y_{1},z_{2})=\bigg{(}\frac{(y_{0}-3y_{1})(1+y_{0}z_{2}-y_{0}^{2}z_{2}^{2})}{1+y_{0}y_{1}z_{2}},\\ \frac{(y_{0}-2y_{1})(1+y_{0}z_{2}-y_{0}^{2}z_{2}^{2})}{1+y_{0}y_{1}z_{2}},\;\frac{z_{2}(4+y_{0}z_{2}-y_{0}^{2}z_{2}^{2})(1+y_{0}y_{1}z_{2})}{1-3y_{0}y_{1}z_{2}^{2}+y_{0}^{2}y_{1}z_{2}^{3}}\bigg{)}.$
The fixed point scheme (near $E_{1}$) is defined by: $I(y_{0})=y_{0}^{2}z_{2}(1-y_{1}+O(y_{0}))$, $I(y_{1})=y_{0}(1+y_{1}z_{2}-y_{1}^{2}z_{2}+O(y_{0}))$, and $I(z_{2})=y_{0}z_{2}^{2}(1+y_{1}+O(y_{0}))$. So $Y_{2}^{G}$ is the Cartier divisor $E_{1}$ except where $y_{0}=0$ (so $3=0$), $0=1+y_{1}z_{2}-y_{1}^{2}z_{2}$, and $0=z_{2}^{2}(1+y_{1})$. So we have one bad point in this open set, $(y_{0},y_{1},z_{2})=(0,-1,-1)$. This already appeared in the chart $\{w_{0}=1\}$ over $\{y_{1}=1\}$ (as the point $(y_{0},w_{1},y_{2})=(0,-1,-1)$).
The other chart is $\{z_{2}=1\}\subset Y_{2}$ over $\{y_{2}=1\}$ in $Y_{1}$, with coordinates $(z_{0},y_{1},e_{2})$. Here $y_{0}=e_{2}z_{0}$, $E_{0}$ does not appear, and $E_{1}=\{e_{2}=0\}$. Here $G$ acts by
[…]
We know that the fixed point scheme $Y_{2}^{G}$ is contained in $E_{1}$ as a set. (Also, $3=O(e_{2})$ by the equation for $Y_{2}$.) Explicitly, the fixed point scheme (near $E_{2}$) is defined by: $I(z_{0})=e_{2}(-z_{0}-z_{0}y_{1}+O(e_{2}))$, $I(y_{1})=e_{2}(z_{0}+y_{1}-y_{1}^{2}+O(e_{2}))$, and $I(e_{2})=e_{2}^{2}(-1+O(e_{2}))$. So $Y_{2}^{G}$ is the Cartier divisor $E_{1}$ except where $e_{2}=0$ (so $3=0$), $0=z_{0}(1+y_{1})$, and $0=z_{0}+y_{1}-y_{1}^{2}$. So we see three bad points in $E_{1}$: $(z_{0},y_{1},x_{2})=(0,1,0)$, $(-1,-1,0)$, or $(0,0,0)$. The first two points appeared in the chart $\{w_{1}=1\}$ in $Y_{2}$ over $\{y_{1}=1\}$ in $Y_{1}$ (as $(w_{0},x_{1},y_{2})=(0,0,1)$ or $(-1,0,-1)$, respectively). On the other hand, the origin in this chart is new; so we have a total of six bad points.
Thus the fixed point scheme $Y_{2}^{G}$ is $E_{1}$ with multiplicity 1 except at six points on $E_{1}$. As in Theorem 6.1, Theorem 2.2 shows that five of the singular points of $Y_{2}/G$ are toric singularities of the form $\frac{1}{3}(1,1,2)$ (hence terminal), while the sixth is a toric singularity of the form $\frac{1}{3}(1,1,1)$ (hence canonical). In fact, our calculations of $I=\sigma-1$ on the coordinates in this section are identical to those in section 6, to the accuracy we state. We also need to check the assumption in Theorem 2.2 that $p\in e^{p-1}\operatorname{\mathfrak{m}}$, that is, that $3\in e^{2}\operatorname{\mathfrak{m}}$. This is true because $3=e_{2}^{3}(\text{unit})$ on $Y_{2}$, and $e_{2}$ is a multiple of the function $e$ defining $E_{1}$ in each coordinate chart; so 3 is in the ideal $(e^{3})$, hence in $e^{2}\operatorname{\mathfrak{m}}$ at each of the bad points. As a result, Theorem 2.2 gives the conclusions above about the 6 singular points of $Y_{2}/G$.
The calculation of the discrepancies of $Y_{0}/G$ is likewise identical to the calculation in section 6. Therefore, $Y_{0}/G$ is terminal. We showed earlier that it is not Cohen-Macaulay. Theorem 7.1 is proved. ∎
Remark 7.2.
In Theorem 7.1, the canonical class of $Y_{0}/G$ is not Cartier. I expect that there is also a 3-dimensional scheme $X$, flat over $\text{\bf Z}_{3}$, that is terminal and non-Cohen-Macaulay with $K_{X}$ Cartier. Namely, one should replace the $p$-adic integer ring $R=\text{\bf Z}_{3}[\zeta_{9}]^{\text{\bf Z}/2}$ in Theorem 7.1 by $S=\text{\bf Z}_{3}[\zeta_{9}]$, with the action of $G=\text{\bf Z}/3\subset(\text{\bf Z}/9)^{*}$. The point is that the canonical sheaf of $S$ over $\text{\bf Z}_{3}$ has a $G$-equivariant trivialization.
8 Characteristic 5
Theorem 8.1.
Let the group $G=\text{\bf Z}/5$ with generator $\tau$ act on the quintic del Pezzo surface $S_{5}$ over $\text{\bf F}_{5}$ by an embedding of $G$ into the symmetric group $\Sigma_{5}=\operatorname{Aut}(S_{5})$, and let $G$ act on $\text{\bf P}^{1}$ by
$\tau([y_{0},y_{1}])=[y_{0},y_{0}+y_{1}].$
Then $(S_{5}\times\text{\bf P}^{1})/G$ is terminal, not Cohen-Macaulay, and of dimension 3 over $\text{\bf F}_{5}$.
We define the quintic del Pezzo surface (over any field) as the moduli space $\overline{M_{0,5}}$ of 5-pointed curves of genus 0. That makes it clear that the symmetric group $\Sigma_{5}$ acts on this surface.
Proof.
We work throughout over $k=\text{\bf F}_{5}$. Write $G=\text{\bf Z}/5=\langle\sigma:\sigma^{5}=1\rangle$, with $\tau:=\sigma^{-1}$. Let $Y_{0}=S_{5}\times\text{\bf P}^{1}$ and $X=Y_{0}/G$. In characteristic 5, $G$ has only one fixed point in $S_{5}$, and so $G$ has only one fixed point $P$ in $Y_{0}$. So $X$ is normal of dimension 3, and $X$ is smooth over $k$ outside the image of $P$ (which we also call $P$). Also, $5K_{X}$ is Cartier. By Fogarty, since $A$ is an isolated fixed point of $G=\text{\bf Z}/p$ on a smooth 3-fold in characteristic $p$, $X$ is not Cohen-Macaulay at $P$ [12, Proposition 4].
It remains to show that $X$ is terminal. This example is more complicated than those in characteristics 2 and 3, and it may be impossible to resolve the singularities of $X$ by performing $G$-equivariant blow-ups of $Y_{0}$. (Indeed, in the simpler situation of actions of $G=\text{\bf Z}/p$ in characteristic zero, one cannot always resolve the singularities of a quotient $Y/G$ via $G$-equivariant blow-ups of $Y$ when $p\geq 5$ [17, Claim 2.29.2].) Fortunately, as in earlier sections, we can reach toric singularities after some $G$-equivariant blow-ups. It will then be easy to check that $Y_{0}/G$ is terminal.
The $G$-action on $S_{5}$ over $k$ is given on an open subset isomorphic to an open subset of $A^{2}$ by:
$\tau(s_{0},s_{1})=\bigg{(}\frac{s_{0}-s_{1}+s_{0}^{2}+s_{0}s_{1}}{(1-2s_{0})(1-s_{0}-s_{1})},\;\frac{s_{1}-2s_{0}^{2}-2s_{0}s_{1}}{(1-2s_{0})(1-s_{0}-s_{1})}\bigg{)}.$
Here the fixed point is at the origin. (Section 9 explains where this formula comes from.) So the $G$-action on an open subset $U\subset Y_{0}$ is given (on an open neighborhood of the origin in $A^{3}$) by:
$\tau(s_{0},s_{1},s_{2})=\bigg{(}\frac{s_{0}-s_{1}+s_{0}^{2}+s_{0}s_{1}}{(1-2s_{0})(1-s_{0}-s_{1})},\;\frac{s_{1}-2s_{0}^{2}-2s_{0}s_{1}}{(1-2s_{0})(1-s_{0}-s_{1})},\frac{s_{2}}{1+s_{2}}\bigg{)}.$
As we blow up, we will not need to keep track of the precise affine open set on which $G$ acts, since we are only concerned with the action near the fixed point set.
Let $Y_{1}$ be the blow-up of $Y_{0}$ at the $G$-fixed point, which is the origin in these coordinates. Then the open subset of $Y_{1}$ over $U\subset Y_{0}$ is
$\{((s_{0},s_{1},s_{2}),[y_{0},y_{1},y_{2}])\in U\times\text{\bf P}^{2}:s_{0}y_{1}=s_{1}y_{0},\;s_{0}y_{2}=s_{2}y_{0},s_{1}y_{2}=s_{2}y_{1}\}.$
It turns out that the fixed point set of $G$ on $Y_{1}$ is a curve isomorphic to $\text{\bf P}^{1}$. To check that, first work in the open subset $\{y_{0}=1\}$ in $Y_{1}$. Here $(s_{0},s_{1},s_{2})=(s_{0},s_{0}y_{1},s_{0}y_{2})$, and $G$ acts by
$\tau(s_{0},y_{1},y_{2})=\bigg{(}\frac{s_{0}(1+s_{0}-y_{1}+s_{0}y_{1})}{(1-2s_{0})(1-s_{0}-s_{0}y_{1})},\;\frac{-2s_{0}+y_{1}-2s_{0}y_{1}}{1+s_{0}-y_{1}+s_{0}y_{1}},\;\\ \frac{y_{2}(1-2s_{0})(1-s_{0}-s_{0}y_{1})}{(1+s_{0}y_{2})(1+s_{0}-y_{1}+s_{0}y_{1})}\bigg{)}.$
The exceptional divisor $E_{0}$ is $\{s_{0}=0\}$, in this chart. The fixed point scheme $Y_{1}^{G}$ is defined by the vanishing of: $I(s_{0})=s_{0}(-y_{1}+O(s_{0}))$, $I(y_{1})=(y_{1}^{2}+O(s_{0}))/(1-y_{1}+O(s_{0}))$, and $I(y_{2})=y_{2}(y_{1}+O(s_{0}))/(1-y_{1}+O(s_{0}))$. We know that $Y_{1}^{G}$ is contained (as a set) in $E_{0}$ (since $Y_{0}^{G}$ is only the origin). So the fixed point set is the line $\{0=s_{0}=y_{1}\}$, in this chart.
In the chart $\{y_{1}=1\}$ in $Y_{1}$, we have $s_{0}=s_{1}y_{0}$ and $s_{2}=s_{1}y_{2}$, so we have coordinates $(y_{0},s_{1},y_{2})$. Here $E_{0}=\{s_{1}=0\}$. We can write the action of $G$ in these coordinates (for example using Magma). We find that the fixed point scheme $Y_{1}^{G}$ is defined by the vanishing of: $I(y_{0})=-1+O(s_{1})$, $I(s_{1})=s_{1}^{2}(1+y_{0}-2y_{0}^{2})$, and $I(y_{2})=s_{1}y_{2}(-1-y_{0}-y_{2}+2y_{0}^{2})$. Since $Y_{1}^{G}$ is contained (as a set) in $E_{0}$, the first equation shows that $Y_{1}^{G}$ is empty, in this chart. In the last chart $\{y_{2}=1\}$ in $Y_{1}$, we have coordinates $(y_{0},y_{1},s_{2})$, and $E_{0}=\{s_{2}=0\}$. The fixed point scheme is defined by: $I(y_{0})=-y_{1}+O(s_{2})$, $I(y_{1})=s_{2}(y_{1}+y_{1}^{2}+y_{0}y_{1}-2y_{0}^{2}+O(s_{2}))$, and $I(s_{2})=s_{2}^{2}(-1+O(s_{2}))$. Since $Y_{1}^{G}$ is contained (as a set) in $E_{0}$, the fixed point set is the line $\{0=y_{1}=s_{2}\}$, the same line seen in an earlier chart.
Thus $(Y_{1}^{G})_{\operatorname{red}}$ is isomorphic to $\text{\bf P}^{1}$. Our criterion for a quotient by $G$ to have toric singularities (Theorem 2.2) requires the $G$-fixed locus to have codimension 1; so let $Y_{2}$ be the blow-up of $Y_{1}$ along this $\text{\bf P}^{1}$. Clearly $G$ continues to act on $Y_{2}$. The exceptional divisor $E_{1}$ in $Y_{2}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}$, and so the natural way to cover $E_{1}$ by affine charts involves 4 charts, as follows.
[…]
Figure 3: For $j=0,1,2$, $G$ acts freely on $Y_{j}$ outside the red loci.
Over the open set $\{y_{0}=1\}$ in $Y_{1}$, $Y_{2}$ is the blow-up along the $G$-fixed curve $\{0=s_{0}=y_{1}\}$, so $Y_{2}$ has coordinates $((s_{0},y_{1},y_{2}),[w_{0},w_{1}])$. First take $\{w_{0}=1\}$, so $y_{1}=s_{0}w_{1}$, and we have coordinates $(s_{0},w_{1},y_{2})$. Here $E_{0}$ does not appear, and $E_{1}=\{s_{0}=0\}$. The fixed point scheme $Y_{2}^{G}$ is defined by: $I(s_{0})=s_{0}^{2}(-1-w_{1}+O(s_{0}))$, $I(w_{1})=-2+O(s_{0})$, and $I(y_{2})=s_{0}y_{2}(1+w_{1}-y_{2}+O(s_{0}))$. We know that the fixed point set is contained in $E_{1}$, and so the formula for $I(w_{1})$ implies that $Y_{2}^{G}$ is empty, in this chart. In the other chart $\{w_{1}=1\}$ in $Y_{2}$ over the same open set in $Y_{1}$, we have $s_{0}=y_{1}w_{0}$, and so $Y_{2}$ has coordinates $(w_{0},y_{1},y_{2})$. Here $E_{0}=\{w_{0}=0\}$ and $E_{1}=\{y_{1}=0\}$. The fixed point scheme is defined by $I(w_{0})=w_{0}(2w_{0}+O(y_{1}))/(1-2w_{0}+O(y_{1}))$, $I(y_{1})=y_{1}(-2w_{0}+O(y_{1}))$, and $I(y_{2})=y_{1}y_{2}(1+w_{0}-w_{0}y_{2}+O(y_{1}))$. So $Y_{2}^{G}$ is the line $\{0=w_{0}=y_{1}\}=E_{0}\cap E_{1}$, in this chart.
To see the rest of $E_{1}\subset Y_{2}$, work over the open set $\{y_{2}=1\}$ in $Y_{1}$. Here $Y_{2}$ is the blow-up along the $G$-fixed curve $\{0=y_{1}=s_{2}\}$, so $Y_{2}$ has coordinates $((y_{0},y_{1},s_{2}),[r_{1},r_{2}])$. First take $\{r_{1}=1\}$ in $Y_{2}$, so $s_{2}=y_{1}r_{2}$, and we have coordinates $(y_{0},y_{1},r_{2})$. Here $E_{0}=\{r_{2}=0\}$ and $E_{1}=\{y_{1}=0\}$. Here $Y_{2}^{G}$ is given by $I(y_{0})=y_{1}(-1+y_{0}r_{2}-y_{0}^{2}r_{2}+O(y_{1}))$, $I(y_{1})=y_{1}r_{2}(-2y_{0}^{2}+O(y_{1}))$, and $I(r_{2})=r_{2}^{2}(2y_{0}^{2}+O(y_{1}))/(1-2y_{0}^{2}r_{2}+O(y_{1}))$. We know that the fixed point set is contained in $E_{1}$, and we read off that it is the union of the two lines $\{0=y_{1}=r_{2}\}=E_{0}\cap E_{1}$ and $\{0=y_{0}=y_{1}\}$ in $E_{1}$. The first curve appeared in an earlier chart, and the second is new. Finally, the other open set is $\{r_{2}=1\}$ in $Y_{2}$, so $y_{1}=s_{2}r_{1}$, and we have coordinates $(y_{0},r_{1},s_{2})$. Here $E_{0}$ does not appear, and $E_{1}=\{s_{2}=0\}$. Here $Y_{2}^{G}$ is given by $I(y_{0})=s_{2}(y_{0}-r_{1}-y_{0}^{2}+O(s_{2}))$, $I(r_{1})=-2y_{0}^{2}+O(s_{2})$, and $I(s_{2})=s_{2}^{2}(-1+O(s_{2}))$. We read off that the fixed point set is the curve $\{0=y_{0}=s_{2}\}$, which is the second curve in the previous chart.
Thus $(Y_{2})^{G}$ as a set is the union of two $\text{\bf P}^{1}$’s meeting at a point. We are trying to make the fixed locus have codimension 1, and so our next step is to blow up one of those curves. Namely, let $Y_{3}$ be the blow-up of $Y_{2}$ along the $G$-fixed curve $E_{0}\cap E_{1}$. The exceptional divisor $E_{2}$ in $Y_{3}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}$, and so we need to look at four affine charts to see all of it.
Figure 4: For $j=3,4$, $G$ acts freely on $Y_{j}$ outside the shaded or marked loci; $Y_{j}/G$ is regular outside the marked loci; and $Y_{j}/G$ has toric singularities outside the red loci.
First, work over the open set $\{r_{1}=1\}$ in $Y_{2}$ over $\{y_{2}=1\}$ in $Y_{1}$. Then $Y_{3}$ is the blow-up along the curve $\{0=y_{1}=r_{2}\}=E_{0}\cap E_{1}$, and so $Y_{3}$ has coordinates $(y_{0},y_{1},r_{2}),[z_{1},z_{2}]$. First take $\{z_{1}=1\}$, so $r_{2}=y_{1}z_{2}$, and we have coordinates $(y_{0},y_{1},z_{2})$. Here $E_{0}=\{z_{2}=0\}$, $E_{1}$ does not appear, and $E_{2}=\{y_{1}=0\}$. The fixed point scheme $Y_{3}^{G}$ is defined by: $I(y_{0})=y_{1}(-1+O(y_{1}))$, $I(y_{1})=y_{1}^{2}z_{2}(-2y_{0}^{2}+O(y_{1}))$, and $I(z_{2})=y_{1}z_{2}^{2}(-y_{0}^{2}+O(y_{1}))$. These equations are equivalent to $y_{1}=0$, near $E_{2}$; so the fixed point scheme $Y_{2}^{G}$ is the Cartier divisor $E_{2}$, in this chart. (Thus, by Theorem 2.1, $Y_{2}/G$ is smooth over $k=\text{\bf F}_{5}$, in this open set.)
The other chart is $\{z_{2}=1\}$ in $Y_{3}$, so $y_{1}=r_{2}z_{1}$, and we have coordinates $(y_{0},z_{1},r_{2})$. Here $E_{0}$ does not appear, $E_{1}=\{z_{1}=0\}$, and $E_{2}=\{r_{2}=0\}$. The fixed point scheme $Y_{3}^{G}$ is given by $I(y_{0})=z_{1}r_{2}(-1+O(r_{2}))$, $I(z_{1})=z_{1}r_{2}(y_{0}^{2}+O(r_{2}))$, and $I(r_{2})=r_{2}^{2}(2y_{0}^{2}+O(r_{2}))$. The fixed point scheme is generically $E_{2}$ with multiplicity 1, together with the other fixed curve we knew from $Y_{2}$, here given by $\{0=y_{0}=z_{1}\}\subset E_{1}$. In more detail, the “bad locus” where the scheme $Y_{3}^{G}$ is not just $E_{2}$ as a Cartier divisor is given by removing a factor of $r_{2}$ from these equations, yielding: $0=z_{1}(-1+O(r_{2}))$, $0=z_{1}(y_{0}^{2}+O(r_{2}))$, and $0=r_{2}(2y_{0}^{2}+O(r_{2}))$. We know the fixed locus away from $E_{2}$, so assume that $r_{2}=0$; then these equations show that the bad locus inside $E_{2}$ is the curve $\{0=z_{1}=r_{2}\}=E_{1}\cap E_{2}$.
Fortunately, Theorem 2.2 implies that $Y_{3}/G$ has toric singularities at points of $E_{1}\cap E_{2}$ outside the origin. Namely, let $e=r_{2}$ and $s=z_{1}$; then $I(s)=es(\text{unit})$ near $E_{1}\cap E_{2}=\{0=z_{1}=r_{2}\}$ outside the origin. The theorem gives that $Y_{3}/G$ has singularity $\frac{1}{5}(0,1,2)$ at points of $E_{1}\cap E_{2}$ outside the origin.
To see all of $E_{2}$, we also have to work over $\{w_{1}=1\}$ in $Y_{2}$, with coordinates $(w_{0},y_{1},y_{2})$, over $\{y_{0}=1\}$ in $Y_{1}$. Here $Y_{3}$ is the blow-up along the $G$-fixed curve $\{0=w_{0}=y_{1}\}=E_{0}\cap E_{1}$, so $Y_{3}$ has coordinates $(w_{0},y_{1},y_{2}),[v_{0},v_{1}]$. First take $\{v_{0}=1\}$, so $y_{1}=w_{0}v_{1}$, and we have coordinates $(w_{0},v_{1},y_{2})$ on $Y_{3}$. Here $E_{0}$ does not appear, $E_{1}=\{v_{1}=0\}$, and $E_{2}=\{w_{0}=0\}$. The fixed point scheme is defined by: $I(w_{0})=w_{0}^{2}(2-2v_{1}+O(w_{0}))$, $I(v_{1})=w_{0}v_{1}(1-2v_{1}+O(w_{0}))$, and $I(y_{2})=w_{0}v_{1}y_{2}(1+O(w_{0}))$. In the chart we are working over in $Y_{2}$, the fixed set $Y_{2}^{G}$ is only the curve $E_{0}\cap E_{1}$ we are blowing up, and so $Y_{3}^{G}$ (in this chart) is contained in $E_{2}$ as a set. By the equations, $Y_{3}^{G}$ is generically the Cartier divisor $E_{2}$, and the bad locus (where that fails) is given by $0=w_{0}$, $0=v_{1}(1-2v_{1})$, and $0=v_{1}y_{2}$. So the bad locus is the union of the curve $\{0=w_{0}=v_{1}\}=E_{1}\cap E_{2}$ and the point $(w_{0},v_{1},y_{2})=(0,-2,0)$ in $E_{2}$. By Theorem 2.2 (using $e=s=w_{0}$), $Y_{3}/G$ has singularity $\frac{1}{5}(2,1,0)$ everywhere on the curve $E_{1}\cap E_{2}$ (in this chart), in agreement with an earlier calculation.
To analyze the bad point above, change coordinates temporarily by $t_{1}=v_{1}+2$; then the bad point becomes the origin in coordinates $(w_{0},t_{1},y_{2})$. In these coordinates, we have $I(w_{0})=w_{0}^{2}(1-2t_{1}+O(w_{0}))$, $I(t_{1})=I(v_{1})=(-t_{1}-2t_{1}^{2}+O(w_{0}))$, and $I(y_{2})=w_{0}y_{2}(-2+O(w_{0}))$. Theorem 2.2 applies, with $s=e=w_{0}$, and we read off that $Y_{3}/G$ has singularity $\frac{1}{5}(1,-1,-2)$ at this point. That is terminal by the Reid-Tai criterion (Theorem 1.1, which will be convenient.
The last chart we need to consider in $Y_{3}$ is the other open set $\{v_{1}=1\}$ over the open set above in $Y_{2}$, $\{w_{1}=1\}\subset Y_{2}$ over $\{y_{0}=1\}\subset Y_{1}$. So $w_{0}=y_{1}v_{0}$, and we have coordinates $(v_{0},y_{1},y_{2})$. Here $E_{0}=\{v_{0}=0\}$, $E_{1}$ does not appear, and $E_{2}=\{y_{1}=0\}$. Here $Y_{3}^{G}$ is defined by: $I(v_{0})=v_{0}y_{1}(2-v_{0}+O(y_{1}))$, $I(y_{1})=y_{1}^{2}(1-2v_{0}+O(y_{1}))$, and $I(y_{2})=y_{1}y_{2}(1+O(y_{1}))$. As in the previous chart, we know that $Y_{3}^{G}$ is contained in $E_{2}$ as a set. By the equations, $Y_{3}^{G}$ is generically the Cartier divisor $E_{2}$, and the bad locus (where that fails) is given by $0=y_{1}$, $0=v_{0}(2-v_{0})$, and $0=y_{2}$. Thus there are two bad points in this chart, $(v_{0},y_{1},y_{2})$ equal to $(2,0,0)\in E_{2}$ or $(0,0,0)\in E_{0}\cap E_{2}$. The first is the bad point from the previous chart, but the second one is new. Theorem 2.2 works to analyze the second point (the origin), with $e=s=y_{1}$. We read off that $Y_{3}/G$ has singularity $\frac{1}{5}(2,1,1)$ at this point.
That finishes the analysis of $Y_{3}$. In particular, as a set, $Y_{3}^{G}$ is the union of the divisor $E_{2}$ and a curve in $E_{1}$. It is tempting to blow up the $G$-fixed curve next, but that leads to a large number of blow-ups over one point of the curve, where the fixed point scheme is especially complicated. We therefore define $Y_{4}$ as the blow-up at that point, and only later blow up the whole curve. This leads more efficiently to toric singularities.
Namely, let $Y_{4}$ be the blow-up of $Y_{3}$ at the origin in the chart $\{r_{2}=1\}$ in $Y_{2}$ (unchanged in $Y_{3}$), with coordinates $(y_{0},r_{1},s_{2})$. So $Y_{4}$ has coordinates $(y_{0},r_{1},s_{2}),[q_{0},q_{1},q_{2}]$. The exceptional divisor $E_{3}$ is isomorphic to $\text{\bf P}^{2}$, and so it is covered by 3 affine charts. First take $\{q_{0}=1\}$ in $Y_{4}$, so $r_{1}=y_{0}q_{1}$ and $s_{2}=y_{0}q_{2}$, and we have coordinates $(y_{0},q_{1},q_{2})$. Here $E_{1}=\{q_{2}=0\}$ and $E_{3}=\{y_{0}=0\}$. The fixed point scheme $Y_{4}^{G}$ is defined by: $I(y_{0})=y_{0}^{2}q_{2}(1-q_{1}+O(y_{0}))$, $I(q_{1})=y_{0}(-2+q_{1}q_{2}+q_{1}^{2}q_{2}+O(y_{0}))$, and $I(q_{2})=y_{0}q_{2}^{2}(-2+q_{1}+O(y_{0}))$. So $Y_{4}^{G}$ is generically the Cartier divisor $E_{3}$; the $G$-fixed curve in $E_{1}$ does not appear in this chart. The bad locus (where the scheme $Y_{4}^{G}$ is not just $E_{3}$) is given by $0=y_{0}$, $0=-2+q_{1}q_{2}+q_{1}^{2}q_{2}$, and $0=q_{2}^{2}(-2+q_{1})$. By the second equation, $q_{2}\neq 0$, and so the third equation gives that $q_{1}=2$. Then the second equation gives that $0=-2+2q_{2}-q_{2}=-2+q_{2}$, so $q_{2}=2$. That is, there is only one bad point in this chart, $(y_{0},q_{1},q_{2})=(0,2,2)\in E_{3}$. To analyze that point, change coordinates temporarily by $s_{1}=q_{1}-2$ and $s_{2}=q_{2}-2$. In these coordinates, $I(y_{0})=y_{0}^{2}(-2-s_{1}-s_{2}-s_{1}s_{2}+O(y_{0}))$, $I(s_{1})=I(q_{1})=y_{0}(s_{2}+2s_{1}^{2}+s_{1}^{2}s_{2}+O(y_{0}))$, and $I(s_{2})=I(q_{2})=y_{0}(-s_{1}-s_{1}s_{2}+s_{1}s_{2}^{2}+O(y_{0}))$. By Theorem 2.2, with $e=s=y_{0}$, $Y_{4}/G$ has a $\mu_{5}$-quotient singularity. Explicitly, the linear map $\varphi$ over $k$ in the theorem is $\varphi(y_{0})=-2y_{0}$, $\varphi(s_{1})=s_{2}$, and $\varphi(s_{2})=-s_{1}$, which has eigenvalues $-2,2,-2$. So $Y_{4}/G$ has singularity $\frac{1}{5}(-2,2,-2)$ at this point. This is terminal, by the Reid-Tai criterion.
Next, take the open set $\{q_{1}=1\}$ in $Y_{4}$, so $y_{0}=r_{1}q_{0}$ and $s_{2}=r_{1}q_{2}$, and $Y_{4}$ has coordinates $(q_{0},r_{1},q_{2})$. Here $E_{1}=\{q_{2}=0\}$ and $E_{3}=\{r_{1}=0\}$. The fixed point scheme $Y_{4}^{G}$ is defined by: $I(q_{0})=r_{1}(-q_{2}-q_{0}q_{2}+2q_{0}^{3}+O(r_{1}))$, $I(r_{1})=r_{1}^{2}(2q_{2}-2q_{0}^{2}+O(r_{1}))$, and $I(q_{2})=r_{1}q_{2}(2q_{2}+2q_{0}^{2}+O(r_{1}))$. So $Y_{4}^{G}$ is generically the Cartier divisor $E_{3}$, together with the $G$-fixed curve $\{0=q_{0}=q_{2}\}$ in $E_{1}$. The bad locus in $E_{3}$ is given by $0=r_{1}$, $0=-q_{2}-q_{0}q_{2}+2q_{0}^{3}$, and $0=q_{2}(2q_{2}+2q_{0}^{2})$. This yields two bad points, $(q_{0},r_{1},q_{2})$ equal to $(-2,0,1)$ or $(0,0,0)$. The first one is the bad point from the previous chart, and the second is not surprising, as it is the intersection point of $E_{3}$ with the $G$-fixed curve.
Finally, take the open set $\{q_{2}=1\}$ in $Y_{4}$, so $y_{0}=s_{2}q_{0}$ and $r_{1}=s_{2}q_{1}$, and we have coordinates $(q_{0},q_{1},s_{2})$. Here $E_{1}$ does not appear, and $E_{3}=\{s_{2}=0\}$. The fixed point scheme $Y_{4}^{G}$ is defined by: $I(q_{0})=s_{2}(2q_{0}-q_{1}+O(s_{2}))$, $I(q_{1})=s_{2}(-2q_{1}-2q_{0}^{2}+O(s_{2}))$, and $I(s_{2})=s_{2}^{2}(-1+O(s_{2}))$. So $Y_{4}^{G}$ is generically $E_{3}$. The bad locus in $E_{3}$ is given by: $0=s_{2}$, $0=2q_{0}-q_{1}$, and $0=-2q_{1}-2q_{0}^{2}$. This yields two bad points, $(q_{0},q_{1},s_{2})$ equal to $(-2,1,0)$ (seen in the previous two charts) or $(0,0,0)$, which is new. Theorem 2.2 applies at this new point, with $e=s=s_{2}$. The $k$-linear map $\varphi$ is given by $\varphi(q_{0})=2q_{0}-q_{1}$, $\varphi(q_{1})=-2q_{1}$, and $\varphi(s_{2})=-s_{2}$. So $\varphi$ has eigenvalues $(2,-2,-1)$, and hence $Y_{4}/G$ has singularity $\frac{1}{5}(2,-2,-1)$ at this point. This is terminal, by the Reid-Tai criterion.
That completes our description of $Y_{4}$. Next, let $Y_{5}$ be the blow-up of $Y_{4}$ along the $G$-fixed curve in $E_{1}$. The exceptional divisor $E_{4}$ in $Y_{5}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}$, and so it is covered by four affine charts. First work over the open set $\{z_{2}=1\}$ in $Y_{3}$ (unchanged in $Y_{4}$), with coordinates $(y_{0},z_{1},r_{2})$; this contains the point where the $G$-fixed curve in $E_{1}=\{z_{1}=0\}$ meets $E_{2}=\{r_{2}=0\}$. Here $Y_{5}$ is the blow-up along the $G$-fixed curve $\{0=y_{0}=z_{1}\}$, and so $Y_{5}$ has coordinates $(y_{0},z_{1},r_{2}),[n_{0},n_{1}]$. First take the open set $\{n_{0}=1\}$ in $Y_{5}$, so $z_{1}=y_{0}n_{1}$, and we have coordinates $(y_{0},n_{1},r_{2})$. Here $E_{1}=\{n_{1}=0\}$, $E_{2}=\{r_{2}=0\}$, and $E_{4}=\{y_{0}=0\}$. The fixed point scheme $Y_{5}^{G}$ is defined by: $I(y_{0})=y_{0}n_{1}r_{2}(-1+O(y_{0}))$, $I(n_{1})=n_{1}r_{2}(n_{1}+O(y_{0}))/(1-n_{1}r_{2}+O(y_{0}))$, and $I(r_{2})=y_{0}r_{2}^{2}(-2n_{1}r_{2}+O(y_{0}))$. So $Y_{5}^{G}$, as a set, is the union of the divisor $E_{2}$ and the curve $\{0=y_{0}=n_{1}\}=E_{1}\cap E_{4}$. (In particular, the fixed point set is still not all of codimension 1.) We have analyzed the bad locus of $E_{2}$ in previous steps, but we have to add here that the bad locus of $E_{2}$ is disjoint from $E_{2}\cap E_{4}$ except for the point where $E_{2}$ meets the $G$-fixed curve, by the formula for $I(n_{1})$.
Figure 5: For $j=5,6$, $G$ acts freely on $Y_{j}$ outside the shaded or marked loci; $Y_{j}/G$ is regular outside the marked loci; and $Y_{j}/G$ has toric singularities outside the red loci.
The other chart is $\{n_{1}=1\}$ in $Y_{5}$, so $y_{0}=z_{1}n_{0}$, and we have coordinates $(n_{0},z_{1},r_{2})$. Here $E_{1}$ does not appear, $E_{2}=\{r_{2}=0\}$, and $E_{4}=\{z_{1}=0\}$. The fixed point scheme $Y_{5}^{G}$ is defined by: $I(n_{0})=r_{2}(-1+O(z_{1}))$, $I(z_{1})=z_{1}^{2}r_{2}(-2r_{2}+O(z_{1}))$, and $I(r_{2})=z_{1}r_{2}^{2}(-2r_{2}+O(z_{1}))$. These equations reduce to $r_{2}=0$ near $E_{4}$, and so $Y_{5}^{G}$ is the Cartier divisor $E_{2}$, in this chart.
To finish our description of $E_{4}$ in $Y_{5}$, we work over the open set where the $G$-fixed curve in $Y_{4}$ meets $E_{3}$, namely $\{q_{1}=1\}$ in $Y_{4}$. Here $Y_{4}$ has coordinates $(q_{0},r_{1},q_{2})$, $E_{1}=\{q_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, and the $G$-fixed curve is $\{0=q_{0}=q_{2}\}$ in $E_{1}$. So the blow-up $Y_{5}$ along the $G$-fixed curve has coordinates $(q_{0},r_{1},q_{2}),[u_{0},u_{2}]$. First take $\{u_{0}=1\}$ in $Y_{5}$, so $q_{2}=q_{0}u_{2}$, and we have coordinates $(q_{0},r_{1},u_{2})$. Here $E_{1}=\{u_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, and $E_{4}=\{q_{0}=0\}$. The fixed point scheme $Y_{5}^{G}$ is defined by: $I(q_{0})=q_{0}r_{1}(-u_{2}+O(q_{0}))$, $I(r_{1})=q_{0}r_{1}^{2}(2u_{2}+O(q_{0}))$, and $I(u_{2})=r_{1}u_{2}^{2}(1+O(q_{0}))/(1-r_{1}u_{2}+O(q_{0}))$. We know the fixed set outside $E_{4}$, and so we read off that the fixed set is the divisor $E_{3}$ together with the $G$-fixed curve $E_{1}\cap E_{4}$ found earlier. We have analyzed the bad set of $E_{3}$ away from $E_{4}$ in earlier blow-ups, and we see from the formula for $I(u_{2})$ that the bad set of $E_{3}$ near $E_{3}\cap E_{4}$ is only the point $E_{1}\cap E_{3}\cap E_{4}$ where the $G$-fixed curve meets $E_{3}$.
The other chart is $\{u_{2}=1\}$ in $Y_{5}$. Here $q_{0}=q_{2}u_{0}$, and so we have coordinates $(u_{0},r_{1},q_{2})$. Here $E_{1}$ does not appear, $E_{3}=\{r_{1}=0\}$, and $E_{4}=\{q_{2}=0\}$. The fixed point scheme $Y_{5}^{G}$ is defined by: $I(u_{0})=r_{1}(-1+O(q_{2}))$, $I(r_{1})=r_{1}^{2}q_{2}(2+O(q_{2}))$, and $I(q_{2})=r_{1}q_{2}^{2}(2+O(q_{2}))$. These equations reduce to $r_{1}=0$ near $E_{4}$, and so the fixed point scheme $Y_{5}^{G}$ is the Cartier divisor $E_{3}$, in this chart.
That completes our description of $Y_{5}$. Let $Y_{6}$ be the blow-up of $Y_{5}$ along the $G$-fixed curve $E_{1}\cap E_{4}$. The exceptional divisor $E_{5}$ in $Y_{6}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}$, covered by four affine charts. First take the open set $\{n_{0}=1\}$ in $Y_{5}$, which contains the point where the $G$-fixed curve meets $E_{2}$. Here $Y_{5}$ has coordinates $(y_{0},n_{1},r_{2})$, with $E_{1}=\{n_{1}=0\}$, $E_{2}=\{r_{2}=0\}$, and $E_{4}=\{y_{0}=0\}$. Since $Y_{6}$ is the blow-up along the $G$-fixed curve $\{0=y_{0}=n_{1}\}=E_{1}\cap E_{4}$, $Y_{6}$ has coordinates $(y_{0},n_{1},r_{2}),[m_{0},m_{1}]$. First take $\{m_{0}=1\}$ in $Y_{6}$, so $n_{1}=y_{0}m_{1}$, and we have coordinates $(y_{0},m_{1},r_{2})$. Here $E_{1}=\{m_{1}=0\}$, $E_{2}=\{r_{2}=0\}$, $E_{4}$ does not appear, and $E_{5}=\{y_{0}=0\}$. The fixed point scheme $Y_{6}^{G}$ is defined by: $I(y_{0})=y_{0}^{2}m_{1}r_{2}(-1+O(y_{0}))$, $I(m_{1})=y_{0}m_{1}r_{2}(2m_{1}+O(y_{0}))$, and $I(r_{2})=y_{0}^{2}r_{2}^{2}(2-2m_{1}r_{2}+O(y_{0}))$. We know the fixed point set away from $E_{5}$, and so we read off that the fixed point scheme is generically the Cartier divisor $E_{2}+E_{5}$. (Since $E_{5}$ is fixed by $G$, we have finally made the fixed point set of codimension 1.) Let $e=y_{0}r_{2}$. The bad locus (where the scheme $Y_{6}^{G}$ is more than the Cartier divisor $E_{2}+E_{5}$), on $E_{5}$, is given by factoring out $e$ from the equations and setting $y_{0}=0$, so we get: $0=y_{0}$ and $0=2m_{1}^{2}$. So, as a set, the bad locus is the curve $\{0=y_{0}=m_{1}\}=E_{1}\cap E_{5}$. Theorem 2.2 does not seem to apply to this curve, and so $Y_{6}/G$ might not have toric singularities there; we will have to blow up one more time.
For now, look at the other open set, $\{m_{1}=1\}$ in $Y_{6}$. So $s_{0}=n_{1}m_{0}$, and we have coordinates $(m_{0},n_{1},r_{2})$. Here $E_{1}$ does not appear, $E_{2}=\{r_{2}=0\}$, $E_{4}=\{m_{0}=0\}$, and $E_{5}=\{n_{1}=0\}$. The fixed point scheme $Y_{6}^{G}$ is defined by: $I(m_{0})=m_{0}n_{1}r_{2}(-2+O(n_{1}))$, $I(n_{1})=n_{1}^{2}r_{2}(1+O(n_{1}))$, and $I(r_{2})=m_{0}n_{1}^{2}r_{2}^{2}(2m_{0}-2r_{2}+O(n_{1}))$. Since we know the fixed point set outside $E_{5}$, we read off that the fixed point scheme is generically the Cartier divisor $E_{2}+E_{5}$. Let $e=n_{1}r_{2}$. The bad locus (where the scheme $Y_{6}^{G}$ is more than the Cartier divisor $E_{2}+E_{5}$) is the curve $\{0=m_{0}=n_{1}\}=E_{4}\cap E_{5}$. Fortunately, Theorem 2.2 applies, with $s=m_{0}$. We read off that $Y_{6}/G$ has singularity $\frac{1}{5}(-2,1,0)$ along the whole curve $E_{4}\cap E_{5}$, in this chart.
To finish describing $E_{5}\subset Y_{6}$, we have to work over the open set $\{u_{0}=1\}$ in $Y_{5}$, where the $G$-fixed curve $E_{1}\cap E_{4}$ in $Y_{5}$ meets $E_{3}$. Here $Y_{5}$ has coordinates $(q_{0},r_{1},u_{2})$, with $E_{1}=\{u_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, and $E_{4}=\{q_{0}=0\}$. Then $Y_{6}$ is the blow-up along the $G$-fixed curve $\{0=q_{0}=u_{2}\}=E_{1}\cap E_{4}$, so $Y_{6}$ has coordinates $(q_{0},r_{1},u_{2}),[t_{0},t_{2}]$. First take $\{t_{0}=1\}$ in $Y_{6}$, so $u_{2}=q_{0}u_{2}$ and we have coordinates $(q_{0},r_{1},t_{2})$. Here $E_{1}=\{t_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, $E_{4}$ does not appear, and $E_{5}=\{q_{0}=0\}$. The fixed point scheme is defined by: $I(q_{0})=q_{0}^{2}r_{1}(-t_{2}+O(q_{0}))$, $I(r_{1})=q_{0}^{2}r_{1}^{2}(-2+2t_{2}+O(q_{0}))$, and $I(t_{2})=q_{0}r_{1}t_{2}(2t_{2}+O(q_{0}))$. So the fixed point scheme is generically $E_{3}+E_{5}$. Let $e=q_{0}r_{1}$. The bad locus (where the scheme $Y_{6}^{G}$ is more than the Cartier divisor $E_{3}+E_{5}$), in $E_{5}$, is given by $0=q_{0}$ and $0=2t_{2}^{2}$, so (as a set) it is the curve $\{0=q_{0}=t_{2}\}=E_{1}\cap E_{5}$, which we met in an earlier chart.
The other chart is $\{t_{2}=1\}$ in $Y_{6}$, so $q_{0}=u_{2}t_{0}$, and we have coordinates $(t_{0},r_{1},u_{2})$. Here $E_{1}$ does not appear, $E_{3}=\{r_{1}=0\}$, $E_{4}=\{t_{0}=0\}$, and $E_{5}=\{u_{2}=0\}$. The fixed point scheme $Y_{6}^{G}$ is defined by: $I(t_{0})=t_{0}r_{1}u_{2}(-2+O(u_{2}))$, $I(r_{1})=t_{0}r_{1}^{2}u_{2}^{2}(2-2t_{0}+O(u_{2}))$, and $I(u_{2})=r_{1}u_{2}^{2}(1+O(u_{2}))$. So $Y_{6}^{G}$ is generically $E_{3}+E_{5}$. Let $e=r_{1}u_{2}$. The bad locus (where the scheme $Y_{6}^{G}$ is more than the Cartier divisor $E_{3}+E_{5}$), in $E_{5}$, is the curve $\{0=t_{0}=u_{2}\}=E_{4}\cap E_{5}$, which we met in an earlier chart. Theorem 2.2 applies, with $s=t_{0}$. Namely, $Y_{6}/G$ has singularity $\frac{1}{5}(-2,0,1)$ everywhere on the curve $E_{4}\cap E_{5}$ in this chart (including the origin, which did not appear in the earlier chart).
That completes our description of $Y_{6}$. In particular, the $G$-fixed locus has codimension 1 in $Y_{6}$, and $Y_{6}/G$ has toric singularities outside the image of the curve $E_{1}\cap E_{5}$. Let $Y_{7}$ be the blow-up of $Y_{6}$ along that curve. The exceptional divisor $E_{6}$ in $Y_{7}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}$, and so we will cover $E_{6}$ with four affine charts. First, work over the open set $\{m_{0}=1\}$ in $Y_{6}$, where the bad curve $E_{1}\cap E_{5}$ meets $E_{2}$. Here $Y_{6}$ has coordinates $(y_{0},m_{1},r_{2})$, with $E_{1}=\{m_{1}=0\}$, $E_{2}=\{r_{2}=0\}$, and $E_{5}=\{y_{0}=0\}$. Since $Y_{7}$ is the blow-up along the curve $\{0=y_{0}=m_{1}\}=E_{1}\cap E_{5}$, $Y_{7}$ has coordinates $(y_{0},m_{1},r_{2}),[j_{0},j_{1}]$.
Figure 6: $G$ acts freely on $Y_{7}$ outside the shaded or marked loci, and $Y_{7}/G$ is regular outside the marked loci. Here $Y_{7}/G$ has toric singularities.
First take $\{j_{0}=1\}$ in $Y_{7}$, so $m_{1}=y_{0}j_{1}$, and we have coordinates $(y_{0},j_{1},r_{2})$. Here $E_{1}=\{j_{1}=0\}$, $E_{2}=\{r_{2}=0\}$, $E_{5}$ does not appear, and $E_{6}=\{y_{0}=0\}$. The fixed point scheme $Y_{7}^{G}$ is defined by: $I(y_{0})=y_{0}^{3}j_{1}r_{2}(-1+O(y_{0}))$, $I(j_{1})=y_{0}^{2}j_{1}r_{2}(1-2j_{1}+O(y_{0}))$, and $I(r_{2})=y_{0}^{2}r_{2}^{2}(2+O(y_{0}))$. So $Y_{7}^{G}$ is generically the Cartier divisor $E_{2}+2E_{6}$. Let $e=y_{0}^{2}r_{2}$. The bad locus (where the scheme $Y_{7}^{G}$ is more than the Cartier divisor $E_{2}+2E_{6}$), in $E_{6}$, is given by $0=y_{0}$, $0=j_{1}(1-2j_{1})$, and $0=r_{2}$, so it consists of the two points $(y_{0},j_{1},r_{2})$ equal to $(0,0,0)=E_{1}\cap E_{2}\cap E_{6}$ or $(0,-2,0)\in E_{2}\cap E_{6}$. At the first point, Theorem 2.2 applies, with $s=r_{2}$. We read off that $Y_{7}/G$ has singularity $(0,1,2)$ everywhere on the curve $E_{1}\cap E_{2}$ (including the origin, which did not appear when we saw $E_{1}\cap E_{2}$ in an earlier chart). To analyze the second point, change coordinates temporarily by $s_{1}=j_{1}+2$; then that point becomes the origin in coordinates $(y_{0},s_{1},r_{2})$. We have $I(y_{0})=y_{0}^{3}r_{2}(2-s_{1}+O(y_{0}))$, $I(s_{1})=I(j_{1})=y_{0}^{2}r_{2}(-s_{1}-2s_{1}^{2}+O(y_{0}))$, and $I(r_{2})=y_{0}^{2}r_{2}^{2}(2+O(y_{0}))$. Theorem 2.2 applies, with $s=r_{2}$. We read off that $Y_{7}/G$ has singularity $\frac{1}{5}(2,-1,2)$ at this point.
The other open set is $\{j_{1}=1\}$ in $Y_{7}$, so $y_{0}=m_{1}j_{0}$, and we have coordinates $(j_{0},m_{1},r_{2})$. Here $E_{1}$ does not appear, $E_{2}=\{r_{2}=0\}$, $E_{5}=\{j_{0}=0\}$, and $E_{6}=\{m_{1}=0\}$. The fixed point scheme $Y_{7}^{G}$ is defined by: $I(j_{0})=j_{0}^{2}m_{1}^{2}r_{2}(2-j_{0}+O(m_{1}))$, $I(m_{1})=j_{0}m_{1}^{3}r_{2}(2+j_{0}+O(m_{1}))$, and $I(r_{2})=j_{0}^{2}m_{1}^{2}r_{2}^{2}(2+O(m_{1}))$. So $Y_{7}^{G}$ is generically $E_{2}+E_{5}+2E_{6}$. Let $e=j_{0}m_{1}^{2}r_{2}$. The bad locus (where the scheme $Y_{7}^{G}$ is more than the Cartier divisor $E_{2}+E_{5}+2E_{6}$), on $E_{6}$, is given by: $0=m_{1}$, $0=j_{0}(2-j_{0})$, and $0=j_{0}r_{2}$. So the bad locus is the union of the curve $\{0=j_{0}=m_{1}\}=E_{5}\cap E_{6}$ and the point $(j_{0},m_{1},r_{2})=(2,0,0)$ in $E_{2}\cap E_{6}$. That point is the one we analyzed in the previous chart. For the curve, Theorem 2.2 applies, using $s=j_{0}$. We read off that $Y_{7}/G$ has singularity $\frac{1}{5}(2,2,0)$ everywhere on the curve $E_{5}\cap E_{6}$, in this chart.
Last, work over the open set $\{t_{0}=1\}$ in $Y_{6}$, where the bad curve $E_{1}\cap E_{5}$ meets $E_{3}$. Here $Y_{6}$ has coordinates $(q_{0},r_{1},t_{2})$, with $E_{1}=\{t_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, and $E_{5}=\{q_{0}=0\}$. We obtain $Y_{7}$ by blowing up along the curve $\{0=q_{0}=t_{2}\}=E_{1}\cap E_{5}$, so $Y_{7}$ has coordinates $(q_{0},r_{1},t_{2}),[x_{0},x_{2}]$. First take $\{x_{0}=1\}$, so $t_{2}=q_{0}x_{2}$, and we have coordinates $(q_{0},r_{1},x_{2})$. Here $E_{1}=\{x_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, $E_{5}$ does not appear, and $E_{6}=\{q_{0}=0\}$. The fixed point scheme $Y_{7}^{G}$ is defined by: $I(q_{0})=q_{0}^{3}r_{1}(2-x_{2}+O(q_{0}))$, $I(r_{1})=q_{0}^{2}r_{1}^{2}(-2+O(q_{0}))$, and $I(x_{2})=q_{0}^{2}r_{1}x_{2}(1-2x_{2}+O(q_{0}))$. So $Y_{7}^{G}$ is generically $E_{3}+2E_{6}$. Let $e=q_{0}^{2}r_{1}$. The bad locus (where the scheme $Y_{7}^{G}$ is more than the Cartier divisor $E_{3}+2E_{6}$), in $E_{6}$, is given by: $0=q_{0}$, $0=r_{1}$, and $0=x_{2}(1-2x_{2})$, so it consists of the two points $(q_{0},r_{1},x_{2})$ equal to $(0,0,0)=E_{1}\cap E_{3}\cap E_{6}$ or $(0,0,-2)$ in $E_{3}\cap E_{6}$. Since $I(r_{1})=er_{1}(\text{unit})$, Theorem 2.2 applies at both points. At the origin, the theorem gives that $Y_{7}/G$ has singularity $\frac{1}{5}(2,-2,1)$, which is terminal. For the other point, change coordinates temporarily by $y_{2}=x_{2}+2$, so that the point becomes the origin in coordinates $(q_{0},r_{1},y_{2})$. We have $I(q_{0})=q_{0}^{3}r_{1}(-1-y_{2}+O(q_{0}))$, $I(r_{1})=q_{0}^{2}r_{1}^{2}(-2+O(q_{0}))$, and $I(y_{2})=I(x_{2})=q_{0}^{2}r_{1}(-y_{2}-2y_{2}^{2}+O(q_{0}))$. So Theorem 2.2 gives that $Y_{7}/G$ has singularity $\frac{1}{5}(-1,-2,-1)$ at this point.
The other chart is $\{x_{2}=1\}$ in $Y_{7}$, so $q_{0}=t_{2}x_{0}$, and we have coordinates $(x_{0},r_{1},t_{2})$. Here $E_{1}$ does not appear, $E_{3}=\{r_{1}=0\}$, $E_{5}=\{x_{0}=0\}$, and $E_{6}=\{t_{2}=0\}$. The fixed point scheme $Y_{7}^{G}$ is defined by: $I(x_{0})=x_{0}^{2}r_{1}t_{2}^{2}(2-x_{0}+O(t_{2}))$, $I(r_{1})=x_{0}^{2}r_{1}^{2}t_{2}^{2}(-2+O(t_{2}))$, and $I(t_{2})=x_{0}r_{1}t_{2}^{3}(2-2x_{0}+O(t_{2}))$. So $Y_{7}^{G}$ is generically $E_{3}+E_{5}+2E_{6}$. Let $e=x_{0}r_{1}t_{2}^{2}$. The bad locus (where $Y_{7}^{G}$ is more than the Cartier divisor $E_{3}+E_{5}+2E_{6}$), in $E_{6}$, is given by: $0=t_{2}$, $0=x_{0}(2-x_{0})$, and $0=x_{0}r_{1}$, which is the union of the curve $\{0=x_{0}=t_{2}\}=E_{5}\cap E_{6}$ and the point $(x_{0},r_{1},t_{2})=(2,0,0)$ in $E_{3}\cap E_{6}$. We analyzed that point in the previous chart. Theorem 2.2 applies to the curve, using $s=x_{0}$. We read off that $Y_{7}/G$ has singularity $\frac{1}{5}(2,0,2)$ everywhere on the curve $E_{5}\cap E_{6}$ (including the origin, which did not appear in the earlier chart where we met this curve).
That completes our analysis of $Y_{7}$; we have shown that $Y_{7}/G$ has toric singularities. It will now be straightforward to show that $Y_{0}/G$ is terminal.
First, we can compute the canonical class of $Y_{7}$, since $Y_{7}$ is obtained from $Y_{0}$ by a sequence of blow-ups along points and smooth curves. Write $E_{j}$ for the strict transform of the exceptional divisor in $Y_{j+1}$ to any higher model. Write $\pi_{ij}$ for the morphism $Y_{i}\to Y_{j}$ (with $i>j$), and also for the resulting morphism $Y_{i}/G\to Y_{j}/G$. First, $K_{Y_{1}}=\pi_{10}^{*}K_{Y_{0}}+2E_{0}$, since $Y_{1}\to Y_{0}$ is the blow-up of a smooth 3-fold at a point. Next, $K_{Y_{2}}=\pi_{21}^{*}K_{Y_{1}}+E_{1}$, since $Y_{2}\to Y_{1}$ is the blow-up along a smooth curve, and we have $\pi_{21}^{*}E_{0}=E_{0}+E_{1}$ because the curve being blown up is contained in $E_{0}$. Likewise, we have:
$K_{Y_{3}}=\pi_{32}^{*}K_{Y_{2}}+E_{2},\;\pi_{32}^{*}E_{0}=E_{0}+E_{2},\;$ $\pi_{32}^{*}E_{1}=E_{1}+E_{2},$
$K_{Y_{4}}=\pi_{43}^{*}K_{Y_{3}}+2E_{3},\;\pi_{43}^{*}E_{0}=E_{0},\;$ $\pi_{43}^{*}E_{1}=E_{1}+E_{3},\;\pi_{43}^{*}E_{2}=E_{2},$
$K_{Y_{5}}=\pi_{54}^{*}K_{Y_{4}}+E_{4},\;\pi_{54}^{*}E_{0}=E_{0},$ $\pi_{54}^{*}E_{1}=E_{1}+E_{4},\;\pi_{54}^{*}E_{2}=E_{2},\;\pi_{54}^{*}E_{3}=E_{3},$
[…]
$\;\pi_{65}^{*}E_{1}=E_{1}+E_{5},\;$ $\pi_{65}^{*}E_{4}=E_{4}+E_{5},$
[…]
$\;\pi_{76}^{*}E_{1}=E_{1}+E_{6},\;$ $\pi_{76}^{*}E_{5}=E_{5}+E_{6}.$
Combining these equations gives that
$K_{Y_{7}}=\pi_{70}^{*}K_{Y_{0}}+2E_{0}+3E_{1}+6E_{2}+5E_{3}+4E_{4}+8E_{5}+12E_{6}.$
Write $f$ for any of the quotient maps $Y_{j}\to Y_{j}/G$. First, $K_{Y_{0}/G}$ is Q-Cartier since $Y_{0}$ is smooth, and $K_{Y_{0}}=f^{*}K_{Y_{0}/G}$ because $f\colon Y_{0}\to Y_{0}/G$ is étale in codimension 1. Next, we computed that the fixed point scheme $Y_{7}^{G}$ is the Cartier divisor $E_{2}+E_{3}+E_{5}+2E_{6}$ outside a codimension-2 subset of $Y_{7}$. By section 3, it follows that
$K_{Y_{7}}$ $=f^{*}K_{Y_{7}/G}+(p-1)(E_{2}+E_{3}+E_{5}+2E_{6})$
$=f^{*}K_{Y_{7}/G}+4E_{2}+4E_{3}+4E_{5}+8E_{6}.$
For each $j\in\{0,\ldots,7\}$, let $F_{j}$ be the image of $E_{j}$ in $Y_{7}/G$, as an irreducible divisor. For $j\in\{0,1,4\}$, $G$ acts nontrivially on $E_{j}$ (so $f$ is unramified along $E_{j}$), and hence $E_{j}=f^{*}F_{j}$. For the other $j$’s, in $\{2,3,5,6\}$, section 3 and our calculations imply that $f$ is fiercely ramified along $E_{j}$, and so again we have $E_{j}=f^{*}F_{j}$. (For example, for $E_{2}$, use the first chart where $E_{2}$ appeared, $\{z_{1}=1\}$ in $Y_{3}$. There $E_{2}$ is the divisor $\{y_{1}=0\}$, and $Y_{3}^{G}$ has multiplicity 1 along $E_{2}$, but $I(y_{1})=y_{1}^{2}z_{2}(-2y_{0}^{2}+O(y_{1}))$ vanishes to order $2>1$ along $E_{2}$; so section 3 gives that $f$ is fiercely ramified along $E_{2}$.)
We can combine these results to compute the discrepancies of the morphism $Y_{7}/G\to Y_{0}/G$. Namely, we have
$f^{*}(K_{Y_{7}/G}-\pi_{70}^{*}K_{Y_{0}/G})$ $=f^{*}K_{Y_{7}/G}-\pi_{70}^{*}f^{*}K_{Y_{0}/G}$
$=K_{Y_{7}}-4E_{2}-4E_{3}-4E_{5}-8E_{6}-\pi_{70}^{*}K_{Y_{0}}$
$=2E_{0}+3E_{1}+2E_{2}+E_{3}+4E_{4}+4E_{5}+4E_{6}$
$=f^{*}(2F_{0}+3F_{1}+2F_{2}+F_{3}+4F_{4}+4F_{5}+4F_{6}).$
Therefore, $K_{Y_{7}/G}=\pi_{70}^{*}(K_{Y_{0}/G})+2F_{0}+3F_{1}+2F_{2}+F_{3}+4F_{4}+4F_{5}+4F_{6}$. In particular, these coefficients are all positive, which is part of showing that $Y_{0}/G$ is terminal. (That would be all we need if $Y_{7}/G$ were smooth.)
To show that $Y_{0}/G$ is terminal, it now suffices to show that the pair $(Y_{7}/G,D)$ is terminal, where $D:=-2F_{0}-3F_{1}-2F_{2}-F_{3}-4F_{4}-4F_{5}-4F_{6}$. Because the coefficients of $D$ are negative (which works to our advantage), this is clear at points where $Y_{7}/G$ is terminal. There are 6 subvarieties (points or curves) where $Y_{7}/G$ is not terminal, as we now address.
(1) Along the curve $E_{1}\cap E_{2}$, $Y_{7}/G$ has singularity $\frac{1}{5}(0,1,2)$, with $F_{1}$ a toric divisor of weight 1 and $F_{2}$ a toric divisor of weight 2, using Theorem 4.1. To show that $(Y_{7}/G,D)=(Y_{7}/G,-3F_{1}-2F_{2}-\cdots)$ is terminal along this curve, we need to show that $54(i\bmod 5)+3(2i\bmod 5)>$ 5 for $i=1,\ldots 4$, by Theorem 1.2. This is clear, since the left side is $\geq 4+3=7>5$.
(2) At a point in $E_{0}\cap E_{2}$, $Y_{7}/G$ has singularity $\frac{1}{5}(2,1,1)$, with $F_{0}$ of weight 2 and $F_{2}$ of weight 1. To show that $(Y_{7}/G,D)=(Y_{7}/G,-2F_{0}-2F_{2}-\cdots)$ is terminal, we need that $53(2i\bmod 5)+3(i\bmod 5)+(i\bmod 5)>$ 5 for $i=1,\ldots 4$. Indeed, the left side is $\geq 3+3+1=7>5$.
(3) Along the curve $E_{4}\cap E_{5}$, $Y_{7}/G$ has singularity $\frac{1}{5}(-2,1,0)$, with $E_{4}$ of weight $-2$ and $E_{5}$ of weight 1. To show that $(Y_{7}/G,D)=(Y_{7}/G,-4F_{4}-4F_{5}-\cdots)$ is terminal, we need that $55(-2i\bmod 5)+5(i\bmod 5)>$ 5 for $i=1,\ldots 4$. Indeed, the left side is $\geq 5+5=10>5$.
(4) At a point in $E_{2}\cap E_{6}$, $Y_{7}/G$ has singularity $\frac{1}{5}(2,-1,2)$, with $E_{2}$ and $E_{6}$ both of weight 2. To show that $(Y_{7}/G,D)=(Y_{7}/G,-2F_{2}-4F_{6}-\cdots)$ is terminal, we need that $53(2i\bmod 5)+5(2i\bmod 5)+(-i\bmod 5)>$ 5 for $i=1,\ldots 4$. Indeed, the left side is $\geq 3+5+1=9>5$.
(5) Along the curve $E_{5}\cap E_{6}$, $Y_{7}/G$ has singularity $\frac{1}{5}(2,2,0)$, with $E_{5}$ and $E_{6}$ both of weight 2. To show that $(Y_{7}/G,D)=(Y_{7}/G,-4F_{5}-4F_{6}-\cdots)$ is terminal, we need that $55(2i\bmod 5)+5(2i\bmod 5)>$ 5 for $i=1,\ldots 4$. Indeed, the left side is $\geq 5+5=10>5$.
(6) At a point in $E_{3}\cap E_{6}$, $Y_{7}/G$ has singularity $\frac{1}{5}(-1,-2,-1)$, with $E_{3}$ of weight $-2$ and $E_{6}$ of weight $-1$. To show that $(Y_{7}/G,D)=(Y_{7}/G,-F_{3}-4F_{6}-\cdots)$ is terminal, we need that $52(-2i\bmod 5)+5(-i\bmod 5)+(-i\bmod 5)>$ 5 for $i=1,\ldots 4$. Indeed, the left side is $\geq 2+5+1=8>5$.
That completes the proof that $Y_{0}/G$ is terminal. Theorem 8 is proved. ∎
Remark 8.2.
The divisor class $\pi^{*}K_{Y_{0}/G}=K_{Y_{7}/G}+D$ happens to be Cartier on the loci (1)–(6), above. However, it is not Cartier at the terminal singularity $\frac{1}{5}(2,-2,-1)$ in $E_{3}$; one can compute that some discrepancies at divisors over that point are not integers. As a result, $K_{Y_{0}/G}$ is not Cartier (as one can also check directly). I expect that there is also a 3-fold $X$ over $\text{\bf F}_{5}$ that is terminal and non-Cohen-Macaulay with $K_{X}$ Cartier. Namely, one should replace $\text{\bf P}^{1}$ in Theorem 8.1 by the Harbater-Katz-Gabber curve of Remark 6.2, now with $p=5$.
9 The example over the 5-adic integers
Theorem 9.1.
Let the group $G=\text{\bf Z}/5$ with generator $\tau$ act on the quintic del Pezzo surface $S_{5}$ over $\text{\bf Z}_{5}$ by an embedding of $G$ into the symmetric group $\Sigma_{5}=\operatorname{Aut}(S_{5})$. Let $R=\text{\bf Z}_{5}[e]/(e^{5}-5e^{4}+25e^{2}-25e+5)$, which is the ring of integers in a Galois extension of $\text{\bf Q}_{5}$ with group $G=\text{\bf Z}/5$. Let $G$ act on the scheme $(S_{5})_{R}$ by the diagonal action on $S_{5}$ and on $R$. Then the scheme $(S_{5})_{R}/G$ is terminal, not Cohen-Macaulay, of dimension 3, and flat over $\text{\bf Z}_{5}$.
We define the quintic del Pezzo surface $S_{5}$ (over any commutative ring) as the moduli space $\overline{M_{0,5}}$ of 5-pointed curves of genus 0. That makes it clear that the symmetric group $\Sigma_{5}$ acts on $S_{5}$.
This example behaves much like the example over $\text{\bf F}_{5}$, Theorem 8.1. In particular, the figures in section 8 accurately depict the blow-ups we make in mixed characteristic $(0,5)$, just as in characteristic 5. We can view $R$ as the subring of the cyclotomic ring $\text{\bf Z}_{5}[\zeta_{25}]$ fixed by the automorphism $\zeta_{25}\mapsto\zeta_{25}^{7}$ of order 4, with $e=1+\zeta_{25}+\zeta_{25}^{-1}+\zeta_{25}^{7}+\zeta_{25}^{-7}$. Informally, $R$ is the simplest ramified $\text{\bf Z}/5$-extension of $\text{\bf Z}_{5}$. More broadly, this action of $G$ on $(S_{2})_{R}$ was chosen as possibly the simplest action of $\text{\bf Z}/5$ on a 5-fold in mixed characteristic $(0,5)$ with an isolated fixed point. The simplicity helps to ensure that the quotient scheme is terminal.
Proof.
We work throughout over $\text{\bf Z}_{5}$. Write $G=\text{\bf Z}/5=\langle\sigma:\sigma^{5}=1\rangle$, with $\tau:=\sigma^{-1}$. By de Fernex [10], the action of $G$ on $S_{5}$ is conjugate to the birational action of $G$ on $\text{\bf P}^{2}$ by
$\tau([x,y,z])=[x(z-y),z(x-y),xz].$
The fixed point over $\text{\bf F}_{5}$ is $[-2,1,-1]$. Let us change variables over $\text{\bf Z}_{5}$ to move that point to $[0,0,1]$ (although it is only fixed over $\text{\bf F}_{5}$). Namely, let $v_{0}=x+2y$, $v_{1}=z-x-y$, and $v_{2}=y$. In these coordinates, the action of $G$ becomes
$\tau[v_{0},v_{1},v_{2}]=[3v_{0}^{2}+3v_{0}v_{1}-12v_{0}v_{2}-8v_{1}v_{2}+10v_{2}^{2},\\ -v_{0}^{2}-v_{0}v_{1}+5v_{0}v_{2}+3v_{1}v_{2}-5v_{2}^{2},(v_{0}+v_{1}-v_{2})(v_{0}-3v_{2})].$
Therefore, in affine coordinates $(s_{0},s_{1}):=(v_{0}/v_{2},v_{1}/v_{2})$, $G$ acts by
$\tau(s_{0},s_{1})=\bigg{(}\frac{10-12s_{0}-8s_{1}+3s_{0}^{2}+3s_{0}s_{1}}{(3-s_{0})(1-s_{0}-s_{1})},\;\frac{-5+5s_{0}+3s_{1}-s_{0}^{2}-s_{0}s_{1}}{(3-s_{0})(1-s_{0}-s_{1})}\bigg{)}.$
This reduces modulo 5 to the formula for the action of $G$ on $S_{5}$ over $\text{\bf F}_{5}$ in section 8.
Let $Y_{0}=(S_{5})_{R}$, with the diagonal action of $G$ on $S_{5}$ and on $R$. Write $e_{2}$ for the generator $e$ of $R$, to fit with our numbering of coordinates on $Y_{0}$; so we have
$50=e_{2}^{5}-5e_{2}^{4}+25e_{2}^{2}-25e_{2}+$ 5 .
Then $G$ acts on an affine neighborhood $U$ of the origin by:
$\tau(s_{0},s_{1},e_{2})=\bigg{(}\frac{10-12s_{0}-8s_{1}+3s_{0}^{2}+3s_{0}s_{1}}{(3-s_{0})(1-s_{0}-s_{1})},\;\frac{-5+5s_{0}+3s_{1}-s_{0}^{2}-s_{0}s_{1}}{(3-s_{0})(1-s_{0}-s_{1})},\\ \frac{1}{7}(20-53e_{2}+8e_{2}^{2}+9e_{2}^{3}-2e_{2}^{4})\bigg{)}.$
(The last expression is a generator of the Galois group of $R$ over $\text{\bf Z}_{5}$, as one can check via Magma. The denominator 7 occurs because the ring of integers of $\text{\bf Q}(\zeta_{25})^{\text{\bf Z}/4}$ is not monogenic, hence not generated over Z by $e_{2}$. This causes no difficulties, because we are working over $\text{\bf Z}_{5}$.) Note that $U$ is written with three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e_{2}$:
$50=e_{2}^{5}-5e_{2}^{4}+25e_{2}^{2}-25e_{2}+$ 5 .
We will apply Theorem 2.2 repeatedly to recognize the singularities of $Y_{j}/G$, for various blow-ups $Y_{j}$ of $Y_{0}$. We remark now that the assumption in Theorem 2.2 that $p\in e^{p-1}\operatorname{\mathfrak{m}}$ will be valid in each case, that is, that $5\in e^{4}\operatorname{\mathfrak{m}}$. Indeed, we have $5=e_{2}^{5}(\text{unit})$ on $Y_{0}$, hence on each blow-up $Y_{j}$, and $e_{2}$ is a multiple of the function $e$ defining the Weil divisor $[Y_{j}^{G}]$ in each case, that being the function $e$ we will use for Theorem 2.2. So 5 is in the ideal $(e^{5})$, hence in $e^{4}\operatorname{\mathfrak{m}}$ at each of the bad points.
Let $X=Y_{0}/G$. Since $G$ acts freely on $\operatorname{Spec}R$ outside its closed point, the only fixed point of $G$ on $Y_{0}$ is the closed point $P\cong\operatorname{Spec}\text{\bf F}_{5}$ given by $(s_{0},s_{1},e_{2})=(0,0,0)$. So $X$ is normal of dimension 3, and $X$ is regular outside the image of $P$, which we also call $P$. Also, $5K_{X}$ is Cartier.
It is not automatic from Fogarty’s results [12], but we can use his methods to show that $X$ is not Cohen-Macaulay at $P$. As in the proof of Theorem 5, using that $G$ has an isolated fixed point on the 3-fold $Y_{0}$, it suffices to show that $H^{1}(G,O(Y_{0}))$ is not zero. This cohomology group is $\ker(\operatorname{tr})/\operatorname{im}(1-\sigma)$ on $O(Y_{0})$, where the trace is $1+\sigma+\cdots+\sigma^{4}$. The equation $50=e_{2}^{5}-5e_{2}^{4}+25e_{2}^{2}-25e_{2}+$ 5 (specifically, the coefficient of $e_{2}^{4}$) implies that $e_{2}$ has trace 5. So $\operatorname{tr}(1-e_{2})=0$, and hence $1-e_{2}$ defines an element of $H^{1}(G,O(Y_{0}))$. Note that $1-e_{2}$ restricts to $1\in O(P)=\text{\bf F}_{5}$ on the fixed point $P$. Therefore, $1-e_{2}$ has nonzero image under the restriction map $H^{1}(G,O(Y_{0}))\to H^{1}(G,O(P)\cong\text{\bf F}_{5}$. So $H^{1}(G,O(Y_{0}))$ is not zero, and hence $Y_{0}/G$ is not Cohen-Macaulay.
It remains to show that $Y_{0}/G$ is terminal. This example is complicated, and it may be impossible to resolve the singularities of $X$ by performing $G$-equivariant blow-ups of $Y_{0}$. Fortunately, as in earlier sections, we can make $Y_{7}/G$ have toric singularities after some $G$-equivariant blow-ups $Y_{7}\to\cdots\to Y_{0}$, exactly parallel to those in the characteristic 5 example (section 8). In fact, all the formulas we write for the fixed point loci will look identical to those in the characteristic 5 example, because we only need to write those formulas modulo suitable error terms. It will then be easy to check that $Y_{0}/G$ is terminal.
The blow-up $Y_{1}\to Y_{0}$ at the $G$-fixed point is, over the open set $U\subset Y_{0}$:
$\{((x_{0},x_{1},e_{2}),[y_{0},y_{1},y_{2}])\in U\times_{\text{\bf Z}_{5}}\text{\bf P}^{2}_{\text{\bf Z}_{5}}:x_{0}y_{1}=x_{1}y_{0},\;x_{0}y_{2}=e_{2}y_{0},\;x_{1}y_{2}=e_{2}y_{1}\}.$
We will see that the fixed point set in $Y_{1}$ is a curve isomorphic to $\text{\bf P}^{1}_{\text{\bf F}_{5}}$. To check that, first work in the open subset $\{y_{0}=1\}$ in $Y_{1}$, with coordinates $(s_{0},y_{1},y_{2})$; here $(s_{0},s_{1},e_{2})=(s_{0},s_{0}y_{1},s_{0}y_{2})$. This is an open neighborhood of the origin in
$\operatorname{Spec}\text{\bf Z}_{5}[s_{0},y_{1},y_{2}]/((s_{0}y_{2})^{5}-5(s_{0}y_{2})^{4}+25(s_{0}y_{2})^{2}-25(s_{0}y_{2})+5),$
by the equation for $e_{2}$. Since $e_{2}=s_{0}y_{2}$, $e_{2}$ is in the ideal $(s_{0})$, and hence 5 is also in $(s_{0})$ (which lets us simplify formulas written modulo $(s_{0})$). It is straightforward to compute how $G$ acts in this chart, but we do not write it out, for brevity. The exceptional divisor $E_{0}$ is $\{s_{0}=0\}$, in this chart (and so $E_{0}$ is isomorphic to $\text{\bf P}^{2}$ over $\text{\bf F}_{5}$). The fixed point scheme $Y_{1}^{G}$ is defined by the vanishing of: $I(s_{0})=s_{0}(-y_{1}+O(s_{0}))$, $I(y_{1})=(y_{1}^{2}+O(s_{0}))/(1-y_{1}+O(s_{0}))$, and $I(y_{2})=y_{2}(y_{1}+O(s_{0}))/(1-y_{1}+O(s_{0}))$. We know that $Y_{1}^{G}$ is contained (as a set) in $E_{0}$ (since $Y_{0}^{G}$ is only the origin in characteristic $5$). So the fixed point set is the line $\{0=s_{0}=y_{1}\}$, in this chart.
In the chart $\{y_{1}=1\}$ in $Y_{1}$, we have $s_{0}=s_{1}y_{0}$ and $e_{2}=s_{1}y_{2}$, so we have coordinates $(y_{0},s_{1},y_{2})$. Here $E_{0}=\{s_{1}=0\}$. We can write the action of $G$ in these coordinates (for example using Magma). We find that the fixed point scheme $Y_{1}^{G}$ is defined by the vanishing of: $I(y_{0})=-1+O(s_{1})$, $I(s_{1})=s_{1}^{2}(1+y_{0}-2y_{0}^{2}+O(s_{1}))$, and $I(y_{2})=s_{1}y_{2}(-1-y_{0}-y_{2}+2y_{0}^{2}+O(s_{1}))$. Since $Y_{1}^{G}$ is contained (as a set) in $E_{0}$, the first equation shows that $Y_{1}^{G}$ is empty, in this chart. In the last chart $\{y_{2}=1\}$ in $Y_{1}$, we have coordinates $(y_{0},y_{1},e_{2})$, and $E_{0}=\{e_{2}=0\}$. The fixed point scheme is defined by: $I(y_{0})=-y_{1}+O(e_{2})$, $I(y_{1})=e(y_{1}+y_{1}^{2}+y_{0}y_{1}-2y_{0}^{2}+O(e_{2}))$, and $I(e_{2})=e_{2}^{2}(-1+O(e_{2}))$. Since $Y_{1}^{G}$ is contained (as a set) in $E_{0}$, the fixed point set is the line $\{0=y_{1}=e_{2}\}$, the same line seen in an earlier chart.
Thus $(Y_{1}^{G})_{\operatorname{red}}$ is isomorphic to $\text{\bf P}^{1}_{\text{\bf F}_{5}}$. Our criterion for a quotient by $G$ to have toric singularities (Theorem 2.2) requires the $G$-fixed locus to have codimension 1; so let $Y_{2}$ be the blow-up of $Y_{1}$ along this $\text{\bf P}^{1}$. Clearly $G$ continues to act on $Y_{2}$. The exceptional divisor $E_{1}$ in $Y_{2}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}_{\text{\bf F}_{5}}$, and so the natural way to cover $E_{1}$ by affine charts involves 4 charts, as follows. (See Figure 3, which applies to the current example as well.)
Over the open set $\{y_{0}=1\}$ in $Y_{1}$, $Y_{2}$ is the blow-up along the $G$-fixed curve $\{0=s_{0}=y_{1}\}$, so $Y_{2}$ has coordinates $((s_{0},y_{1},y_{2}),[w_{0},w_{1}])$. First take $\{w_{0}=1\}$, so $y_{1}=s_{0}w_{1}$, and we have coordinates $(s_{0},w_{1},y_{2})$. As in every other chart, there are three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e$. In this case, we have $e_{2}=s_{0}y_{2}$, and so
$50=(s_{0}y_{2})^{5}-5(s_{0}y_{2})^{4}+25(s_{0}y_{2})^{2}-25(s_{0}y_{2})+$ 5 .
In this chart, $E_{0}$ does not appear, and $E_{1}=\{s_{0}=0\}$. The fixed point scheme $Y_{2}^{G}$ is defined by: $I(s_{0})=s_{0}^{2}(-1-w_{1}+O(s_{0}))$, $I(w_{1})=-2+O(s_{0})$, and $I(y_{2})=s_{0}y_{2}(1+w_{1}-y_{2}+O(s_{0}))$. We know that the fixed point set is contained in $E_{1}$, and so the formula for $I(w_{1})$ implies that $Y_{2}^{G}$ is empty, in this chart.
In the other chart $\{w_{1}=1\}$ in $Y_{2}$ over the same open set in $Y_{1}$, we have $s_{0}=y_{1}w_{0}$, and so $Y_{2}$ has coordinates $(w_{0},y_{1},y_{2})$. Here $E_{0}=\{w_{0}=0\}$, $E_{1}=\{y_{1}=0\}$. Also, $e_{2}=w_{0}y_{1}y_{2}$. The fixed point scheme is defined by $I(w_{0})=w_{0}(2w_{0}+O(y_{1}))/(1-2w_{0}+O(y_{1}))$, $I(y_{1})=y_{1}(-2w_{0}+O(y_{1}))$, and $I(y_{2})=y_{1}y_{2}(1+w_{0}-w_{0}y_{2}+O(y_{1}))$. So $Y_{2}^{G}$ is the line $\{0=w_{0}=y_{1}\}=E_{0}\cap E_{1}$ over $\text{\bf F}_{5}$, in this chart.
To see the rest of $E_{1}\subset Y_{2}$, work over the open set $\{y_{2}=1\}$ in $Y_{1}$. Here $Y_{2}$ is the blow-up along the $G$-fixed curve $\{0=y_{1}=e_{2}\}$, so $Y_{2}$ has coordinates $((y_{0},y_{1},e_{2}),[r_{1},r_{2}])$. First take $\{r_{1}=1\}$ in $Y_{2}$, so $e_{2}=y_{1}r_{2}$, and we have coordinates $(y_{0},y_{1},r_{2})$. Here $E_{0}=\{r_{2}=0\}$ and $E_{1}=\{y_{1}=0\}$. Here $Y_{2}^{G}$ is given by $I(y_{0})=y_{1}(-1+y_{0}r_{2}-y_{0}^{2}r_{2}+O(y_{1}))$, $I(y_{1})=y_{1}r_{2}(-2y_{0}^{2}+O(y_{1}))$, and $I(r_{2})=r_{2}^{2}(2y_{0}^{2}+O(y_{1}))/(1-2y_{0}^{2}r_{2}+O(y_{1}))$. We know that the fixed point set is contained in $E_{1}$, and we read off that it is the union of the two lines $\{0=y_{1}=r_{2}\}=E_{0}\cap E_{1}$ and $\{0=y_{0}=y_{1}\}$ in $E_{1}$. The first curve appeared in an earlier chart, and the second is new. Finally, the other open set is $\{r_{2}=1\}$ in $Y_{2}$, so $y_{1}=e_{2}r_{1}$, and we have coordinates $(y_{0},r_{1},e_{2})$. Here $E_{0}$ does not appear, and $E_{1}=\{e_{2}=0\}$. Here $Y_{2}^{G}$ is given by $I(y_{0})=e_{2}(y_{0}-r_{1}-y_{0}^{2}+O(e_{2}))$, $I(r_{1})=-2y_{0}^{2}+O(e_{2})$, and $I(e_{2})=e_{2}^{2}(-1+O(e_{2}))$. We read off that the fixed point set is the curve $\{0=y_{0}=e_{2}\}$, which is the second curve in the previous chart.
Thus $(Y_{2})^{G}$ as a set is the union of two $\text{\bf P}^{1}$’s over $\text{\bf F}_{5}$ meeting at a point. We are trying to make the fixed locus have codimension 1, and so our next step is to blow up one of those curves. Namely, let $Y_{3}$ be the blow-up of $Y_{2}$ along the $G$-fixed curve $E_{0}\cap E_{1}$. The exceptional divisor $E_{2}$ in $Y_{3}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}_{\text{\bf F}_{5}}$, and so we need to look at four affine charts to see all of it. (See Figure 4, which applies to the current example as well.)
First, work over the open set $\{r_{1}=1\}$ in $Y_{2}$ over $\{y_{2}=1\}$ in $Y_{1}$. Then $Y_{3}$ is the blow-up along the curve $\{0=y_{1}=r_{2}\}=E_{0}\cap E_{1}$, and so $Y_{3}$ has coordinates $(y_{0},y_{1},r_{2}),[z_{1},z_{2}]$. First take $\{z_{1}=1\}$, so $r_{2}=y_{1}z_{2}$, and we have coordinates $(y_{0},y_{1},z_{2})$. As in every other chart, there are three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e_{2}$. In this case, we have $e_{2}=y_{1}^{2}z_{2}$, and so
$50=(y_{1}^{2}z_{2})^{5}-5(y_{1}^{2}z_{2})^{4}+25(y_{1}^{2}z_{2})^{2}-25(y_{1}^{2}z_{2})+$ 5 .
In this chart, $E_{0}=\{z_{2}=0\}$, $E_{1}$ does not appear, and $E_{2}=\{y_{1}=0\}$. The fixed point scheme $Y_{3}^{G}$ is defined by: $I(y_{0})=y_{1}(-1+O(y_{1}))$, $I(y_{1})=y_{1}^{2}z_{2}(-2y_{0}^{2}+O(y_{1}))$, and $I(z_{2})=y_{1}z_{2}^{2}(-y_{0}^{2}+O(y_{1}))$. These equations are equivalent to $y_{1}=0$, near $E_{2}$; so the fixed point scheme $Y_{2}^{G}$ is the Cartier divisor $E_{2}$, in this chart. (Thus, by Theorem 2.1, $Y_{2}/G$ is regular, in this open set.)
The other chart is $\{z_{2}=1\}$ in $Y_{3}$, so $y_{1}=r_{2}z_{1}$, and we have coordinates $(y_{0},z_{1},r_{2})$. Here $E_{0}$ does not appear, $E_{1}=\{z_{1}=0\}$, and $E_{2}=\{r_{2}=0\}$. Also, $e_{2}=z_{1}r_{2}^{2}$. The fixed point scheme $Y_{3}^{G}$ is given by $I(y_{0})=z_{1}r_{2}(-1+O(r_{2}))$, $I(z_{1})=z_{1}r_{2}(y_{0}^{2}+O(r_{2}))$, and $I(r_{2})=r_{2}^{2}(2y_{0}^{2}+O(r_{2}))$. The fixed point scheme is generically $E_{2}$ with multiplicity 1, together with the other fixed curve we knew from $Y_{2}$, here given by $\{0=y_{0}=z_{1}\}\subset E_{1}$. In more detail, the “bad locus” where the scheme $Y_{3}^{G}$ is not just $E_{2}$ as a Cartier divisor is given by removing a factor of $r_{2}$ from these equations, yielding: $0=z_{1}(-1+O(r_{2}))$, $0=z_{1}(y_{0}^{2}+O(r_{2}))$, and $0=r_{2}(2y_{0}^{2}+O(r_{2}))$. We know the fixed locus away from $E_{2}$, so assume that $r_{2}=0$; then these equations show that the bad locus inside $E_{2}$ is the curve $\{0=z_{1}=r_{2}\}=E_{1}\cap E_{2}$.
Fortunately, Theorem 2.2 implies that $Y_{3}/G$ has toric singularities at points of $E_{1}\cap E_{2}$ outside the origin. Namely, let $e=r_{2}$ and $s=z_{1}$; then $I(s)=es(\text{unit})$ near $E_{1}\cap E_{2}=\{0=z_{1}=r_{2}\}$ outside the origin. The theorem gives that $Y_{3}/G$ has singularity $\frac{1}{5}(0,1,2)$ at points of $E_{1}\cap E_{2}$ outside the origin.
To see all of $E_{2}\subset Y_{3}$, we also have to work over $\{w_{1}=1\}$ in $Y_{2}$, with coordinates $(w_{0},y_{1},y_{2})$, over $\{y_{0}=1\}$ in $Y_{1}$. Here $Y_{3}$ is the blow-up along the $G$-fixed curve $\{0=w_{0}=y_{1}\}=E_{0}\cap E_{1}$, so $Y_{3}$ has coordinates $(w_{0},y_{1},y_{2}),[v_{0},v_{1}]$. First take $\{v_{0}=1\}$, so $y_{1}=w_{0}v_{1}$, and we have coordinates $(w_{0},v_{1},y_{2})$ on $Y_{3}$. Here $E_{0}$ does not appear, $E_{1}=\{v_{1}=0\}$, and $E_{2}=\{w_{0}=0\}$. Also, $e_{2}=w_{0}^{2}v_{1}y_{2}$. The fixed point scheme is defined by: $I(w_{0})=w_{0}^{2}(2-2v_{1}+O(w_{0}))$, $I(v_{1})=w_{0}v_{1}(1-2v_{1}+O(w_{0}))$, and $I(y_{2})=w_{0}v_{1}y_{2}(1+O(w_{0}))$. In the chart we are working over in $Y_{2}$, the fixed set $Y_{2}^{G}$ is only the curve $E_{0}\cap E_{1}$ we are blowing up, and so $Y_{3}^{G}$ (in this chart) is contained in $E_{2}$ as a set. By the equations, $Y_{3}^{G}$ is generically the Cartier divisor $E_{2}$, and the bad locus (where that fails) is given by $0=w_{0}$, $0=v_{1}(1-2v_{1})$, and $0=v_{1}y_{2}$. So the bad locus is the union of the curve $\{0=w_{0}=v_{1}\}=E_{1}\cap E_{2}$ and the point $(w_{0},v_{1},y_{2})=(0,-2,0)$ in $E_{2}$. By Theorem 2.2 (using $e=s=w_{0}$), $Y_{3}/G$ has singularity $\frac{1}{5}(2,1,0)$ everywhere on the curve $E_{1}\cap E_{2}$ (in this chart), in agreement with an earlier calculation.
To analyze the bad point above, change coordinates temporarily by $t_{1}=v_{1}+2$; then the bad point becomes the origin in coordinates $(w_{0},t_{1},y_{2})$. In these coordinates, we have $I(w_{0})=w_{0}^{2}(1-2t_{1}+O(w_{0}))$, $I(t_{1})=I(v_{1})=(-t_{1}-2t_{1}^{2}+O(w_{0}))$, and $I(y_{2})=w_{0}y_{2}(-2+O(w_{0}))$. Theorem 2.2 applies, with $s=e=w_{0}$, and we read off that $Y_{3}/G$ has singularity $\frac{1}{5}(1,-1,-2)$ at this point. That is terminal by the Reid-Tai criterion (Theorem 1.1, which will be convenient.
The last chart we need to consider in $Y_{3}$ is the other open set $\{v_{1}=1\}$ over the open set above in $Y_{2}$, $\{w_{1}=1\}\subset Y_{2}$ over $\{y_{0}=1\}\subset Y_{1}$. So $w_{0}=y_{1}v_{0}$, and we have coordinates $(v_{0},y_{1},y_{2})$. Here $E_{0}=\{v_{0}=0\}$, $E_{1}$ does not appear, and $E_{2}=\{y_{1}=0\}$. Also, $e_{2}=v_{0}y_{1}^{2}y_{2}$. Here $Y_{3}^{G}$ is defined by: $I(v_{0})=v_{0}y_{1}(2-v_{0}+O(y_{1}))$, $I(y_{1})=y_{1}^{2}(1-2v_{0}+O(y_{1}))$, and $I(y_{2})=y_{1}y_{2}(1+O(y_{1}))$. As in the previous chart, we know that $Y_{3}^{G}$ is contained in $E_{2}$ as a set. By the equations, $Y_{3}^{G}$ is generically the Cartier divisor $E_{2}$, and the bad locus (where that fails) is given by $0=y_{1}$, $0=v_{0}(2-v_{0})$, and $0=y_{2}$. Thus there are two bad points in this chart, $(v_{0},y_{1},y_{2})$ equal to $(2,0,0)\in E_{2}$ or $(0,0,0)\in E_{0}\cap E_{2}$. The first is the bad point from the previous chart, but the second one is new. Theorem 2.2 works to analyze the second point (the origin), with $e=s=y_{1}$. We read off that $Y_{3}/G$ has singularity $\frac{1}{5}(2,1,1)$ at this point.
That finishes the analysis of $Y_{3}$. In particular, as a set, $Y_{3}^{G}$ is the union of the divisor $E_{2}$ and a curve in $E_{1}$. It is tempting to blow up the $G$-fixed curve next, but that leads to a large number of blow-ups over one point of the curve, where the fixed point scheme is especially complicated. We therefore define $Y_{4}$ as the blow-up at that point, and only later blow up the whole curve. This leads more efficiently to toric singularities.
Namely, let $Y_{4}$ be the blow-up of $Y_{3}$ at the origin in the chart $\{r_{2}=1\}$ in $Y_{2}$ (unchanged in $Y_{3}$), with coordinates $(y_{0},r_{1},e_{2})$. So $Y_{4}$ has coordinates $(y_{0},r_{1},e_{2}),[q_{0},q_{1},q_{2}]$. The exceptional divisor $E_{3}$ is isomorphic to $\text{\bf P}^{2}_{\text{\bf F}_{5}}$, and so it is covered by 3 affine charts. First take $\{q_{0}=1\}$ in $Y_{4}$, so $r_{1}=y_{0}q_{1}$ and $e_{2}=y_{0}q_{2}$, and we have coordinates $(y_{0},q_{1},q_{2})$. As in every other chart, there are three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e_{2}$. In this case, we have $e_{2}=y_{0}q_{2}$, and so
$50=(y_{0}q_{2})^{5}-5(y_{0}q_{2})^{4}+25(y_{0}q_{2})^{2}-25(y_{0}q_{2})+$ 5 .
Here $E_{1}=\{q_{2}=0\}$ and $E_{3}=\{y_{0}=0\}$. The fixed point scheme $Y_{4}^{G}$ is defined by: $I(y_{0})=y_{0}^{2}q_{2}(1-q_{1}+O(y_{0}))$, $I(q_{1})=y_{0}(-2+q_{1}q_{2}+q_{1}^{2}q_{2}+O(y_{0}))$, and $I(q_{2})=y_{0}q_{2}^{2}(-2+q_{1}+O(y_{0}))$. So $Y_{4}^{G}$ is generically the Cartier divisor $E_{3}$; the $G$-fixed curve in $E_{1}$ does not appear in this chart. The bad locus (where the scheme $Y_{4}^{G}$ is not just $E_{3}$) is given by $0=y_{0}$, $0=-2+q_{1}q_{2}+q_{1}^{2}q_{2}$, and $0=q_{2}^{2}(-2+q_{1})$. By the second equation, $q_{2}\neq 0$, and so the third equation gives that $q_{1}=2$. Then the second equation gives that $0=-2+2q_{2}-q_{2}=-2+q_{2}$, so $q_{2}=2$. That is, there is only one bad point in this chart, $(y_{0},q_{1},q_{2})=(0,2,2)\in E_{3}$. To analyze that point, change coordinates temporarily by $s_{1}=q_{1}-2$ and $s_{2}=q_{2}-2$. In these coordinates, $I(y_{0})=y_{0}^{2}(-2-s_{1}-s_{2}-s_{1}s_{2}+O(y_{0}))$, $I(s_{1})=I(q_{1})=y_{0}(s_{2}+2s_{1}^{2}+s_{1}^{2}s_{2}+O(y_{0}))$, and $I(s_{2})=I(q_{2})=y_{0}(-s_{1}-s_{1}s_{2}+s_{1}s_{2}^{2}+O(y_{0}))$. By Theorem 2.2, with $e=s=y_{0}$, $Y_{4}/G$ has a $\mu_{5}$-quotient singularity. Explicitly, the linear map $\varphi$ over $\text{\bf F}_{5}$ in the theorem is $\varphi(y_{0})=-2y_{0}$, $\varphi(s_{1})=s_{2}$, and $\varphi(s_{2})=-s_{1}$, which has eigenvalues $-2,2,-2$. So $Y_{4}/G$ has singularity $\frac{1}{5}(-2,2,-2)$ at this point. This is terminal, by the Reid-Tai criterion.
Next, take the open set $\{q_{1}=1\}$ in $Y_{4}$, so $y_{0}=r_{1}q_{0}$ and $e_{2}=r_{1}q_{2}$, and $Y_{4}$ has coordinates $(q_{0},r_{1},q_{2})$. Here $E_{1}=\{q_{2}=0\}$ and $E_{3}=\{r_{1}=0\}$. The fixed point scheme $Y_{4}^{G}$ is defined by: $I(q_{0})=r_{1}(-q_{2}-q_{0}q_{2}+2q_{0}^{3}+O(r_{1}))$, $I(r_{1})=r_{1}^{2}(2q_{2}-2q_{0}^{2}+O(r_{1}))$, and $I(q_{2})=r_{1}q_{2}(2q_{2}+2q_{0}^{2}+O(r_{1}))$. So $Y_{4}^{G}$ is generically the Cartier divisor $E_{3}$, together with the $G$-fixed curve $\{0=q_{0}=q_{2}\}$ in $E_{1}$. The bad locus in $E_{3}$ is given by $0=r_{1}$, $0=-q_{2}-q_{0}q_{2}+2q_{0}^{3}$, and $0=q_{2}(2q_{2}+2q_{0}^{2})$. This yields two bad points, $(q_{0},r_{1},q_{2})$ equal to $(-2,0,1)$ or $(0,0,0)$. The first one is the bad point from the previous chart, and the second is not surprising, as it is the intersection point of $E_{3}$ with the $G$-fixed curve.
Finally, take the open set $\{q_{2}=1\}$ in $Y_{4}$, so $y_{0}=e_{2}q_{0}$ and $r_{1}=e_{2}q_{1}$, and we have coordinates $(q_{0},q_{1},e_{2})$. Here $E_{1}$ does not appear, and $E_{3}=\{e_{2}=0\}$. The fixed point scheme $Y_{4}^{G}$ is defined by: $I(q_{0})=e(2q_{0}-q_{1}+O(e_{2}))$, $I(q_{1})=e(-2q_{1}-2q_{0}^{2}+O(e_{2}))$, and $I(e_{2})=e_{2}^{2}(-1+O(e_{2}))$. So $Y_{4}^{G}$ is generically $E_{3}$. The bad locus in $E_{3}$ is given by: $0=e_{2}$, $0=2q_{0}-q_{1}$, and $0=-2q_{1}-2q_{0}^{2}$. This yields two bad points, $(q_{0},q_{1},e_{2})$ equal to $(-2,1,0)$ (seen in the previous two charts) or $(0,0,0)$, which is new. Theorem 2.2 applies at this new point, with $e=s:=e_{2}$. The $\text{\bf F}_{5}$-linear map $\varphi$ is given by $\varphi(q_{0})=2q_{0}-q_{1}$, $\varphi(q_{1})=-2q_{1}$, and $\varphi(e_{2})=-e_{2}$. So $\varphi$ has eigenvalues $(2,-2,-1)$, and hence $Y_{4}/G$ has singularity $\frac{1}{5}(2,-2,-1)$ at this point. This is terminal, by the Reid-Tai criterion.
That completes our description of $Y_{4}$. Next, let $Y_{5}$ be the blow-up of $Y_{4}$ along the $G$-fixed curve in $E_{1}$. The exceptional divisor $E_{4}$ in $Y_{5}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}_{\text{\bf F}_{5}}$, and so it is covered by four affine charts. First work over the open set $\{z_{2}=1\}$ in $Y_{3}$ (unchanged in $Y_{4}$), with coordinates $(y_{0},z_{1},r_{2})$; this contains the point where the $G$-fixed curve in $E_{1}=\{z_{1}=0\}$ meets $E_{2}=\{r_{2}=0\}$. Here $Y_{5}$ is the blow-up along the $G$-fixed curve $\{0=y_{0}=z_{1}\}$, and so $Y_{5}$ has coordinates $(y_{0},z_{1},r_{2}),[n_{0},n_{1}]$. First take the open set $\{n_{0}=1\}$ in $Y_{5}$, so $z_{1}=y_{0}n_{1}$, and we have coordinates $(y_{0},n_{1},r_{2})$. As in every other chart, there are three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e_{2}$. In this case, we have $e_{2}=y_{0}n_{1}r_{2}^{2}$, and so
$50=(y_{0}n_{1}r_{2}^{2})^{5}-5(y_{0}n_{1}r_{2}^{2})^{4}+25(y_{0}n_{1}r_{2}^{2})^{2}-25(y_{0}n_{1}r_{2}^{2})+$ 5 .
In this chart, $E_{1}=\{n_{1}=0\}$, $E_{2}=\{r_{2}=0\}$, and $E_{4}=\{y_{0}=0\}$. The fixed point scheme $Y_{5}^{G}$ is defined by: $I(y_{0})=y_{0}n_{1}r_{2}(-1+O(y_{0}))$, $I(n_{1})=n_{1}r_{2}(n_{1}+O(y_{0}))/(1-n_{1}r_{2}+O(y_{0}))$, and $I(r_{2})=y_{0}r_{2}^{2}(-2n_{1}r_{2}+O(y_{0}))$. So $Y_{5}^{G}$, as a set, is the union of the divisor $E_{2}$ and the curve $\{0=y_{0}=n_{1}\}=E_{1}\cap E_{4}$. (In particular, the fixed point set is still not all of codimension 1.) We have analyzed the bad locus of $E_{2}$ in previous steps, but we have to add here that the bad locus of $E_{2}$ is disjoint from $E_{2}\cap E_{4}$ except for the point where $E_{2}$ meets the $G$-fixed curve, by the formula for $I(n_{1})$. (See Figure 5, which applies to the current example as well.)
The other chart is $\{n_{1}=1\}$ in $Y_{5}$, so $y_{0}=z_{1}n_{0}$, and we have coordinates $(n_{0},z_{1},r_{2})$. Here $E_{1}$ does not appear, $E_{2}=\{r_{2}=0\}$, and $E_{4}=\{z_{1}=0\}$. Also, $e_{2}=z_{1}r_{2}^{2}$. The fixed point scheme $Y_{5}^{G}$ is defined by: $I(n_{0})=r_{2}(-1+O(z_{1}))$, $I(z_{1})=z_{1}^{2}r_{2}(-2r_{2}+O(z_{1}))$, and $I(r_{2})=z_{1}r_{2}^{2}(-2r_{2}+O(z_{1}))$. These equations reduce to $r_{2}=0$ near $E_{4}$, and so $Y_{5}^{G}$ is the Cartier divisor $E_{2}$, in this chart.
To finish our description of $E_{4}$ in $Y_{5}$, we work over the open set where the $G$-fixed curve in $Y_{4}$ meets $E_{3}$, namely $\{q_{1}=1\}$ in $Y_{4}$. Here $Y_{4}$ has coordinates $(q_{0},r_{1},q_{2})$, $E_{1}=\{q_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, and the $G$-fixed curve is $\{0=q_{0}=q_{2}\}$ in $E_{1}$. So the blow-up $Y_{5}$ along the $G$-fixed curve has coordinates $(q_{0},r_{1},q_{2}),[u_{0},u_{2}]$. First take $\{u_{0}=1\}$ in $Y_{5}$, so $q_{2}=q_{0}u_{2}$, and we have coordinates $(q_{0},r_{1},u_{2})$. Here $E_{1}=\{u_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, and $E_{4}=\{q_{0}=0\}$. Also, $e_{2}=q_{0}r_{1}u_{2}$. The fixed point scheme $Y_{5}^{G}$ is defined by: $I(q_{0})=q_{0}r_{1}(-u_{2}+O(q_{0}))$, $I(r_{1})=q_{0}r_{1}^{2}(2u_{2}+O(q_{0}))$, and $I(u_{2})=r_{1}u_{2}^{2}(1+O(q_{0}))/(1-r_{1}u_{2}+O(q_{0}))$. We know the fixed set outside $E_{4}$, and so we read off that the fixed set is the divisor $E_{3}$ together with the $G$-fixed curve $E_{1}\cap E_{4}$ found earlier. We have analyzed the bad set of $E_{3}$ away from $E_{4}$ in earlier blow-ups, and we see from the formula for $I(u_{2})$ that the bad set of $E_{3}$ near $E_{3}\cap E_{4}$ is only the point $E_{1}\cap E_{3}\cap E_{4}$ where the $G$-fixed curve meets $E_{3}$.
The other chart is $\{u_{2}=1\}$ in $Y_{5}$. Here $q_{0}=q_{2}u_{0}$, and so we have coordinates $(u_{0},r_{1},q_{2})$. Here $E_{1}$ does not appear, $E_{3}=\{r_{1}=0\}$, and $E_{4}=\{q_{2}=0\}$. Also, $e_{2}=r_{1}q_{2}$. The fixed point scheme $Y_{5}^{G}$ is defined by: $I(u_{0})=r_{1}(-1+O(q_{2}))$, $I(r_{1})=r_{1}^{2}q_{2}(2+O(q_{2}))$, and $I(q_{2})=r_{1}q_{2}^{2}(2+O(q_{2}))$. These equations reduce to $r_{1}=0$ near $E_{4}$, and so the fixed point scheme $Y_{5}^{G}$ is the Cartier divisor $E_{3}$, in this chart.
That completes our description of $Y_{5}$. Let $Y_{6}$ be the blow-up of $Y_{5}$ along the $G$-fixed curve $E_{1}\cap E_{4}$. The exceptional divisor $E_{5}$ in $Y_{6}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}_{\text{\bf F}_{5}}$, covered by four affine charts. First take the open set $\{n_{0}=1\}$ in $Y_{5}$, which contains the point where the $G$-fixed curve meets $E_{2}$. Here $Y_{5}$ has coordinates $(y_{0},n_{1},r_{2})$, with $E_{1}=\{n_{1}=0\}$, $E_{2}=\{r_{2}=0\}$, and $E_{4}=\{y_{0}=0\}$. Since $Y_{6}$ is the blow-up along the $G$-fixed curve $\{0=y_{0}=n_{1}\}=E_{1}\cap E_{4}$, $Y_{6}$ has coordinates $(y_{0},n_{1},r_{2}),[m_{0},m_{1}]$. First take $\{m_{0}=1\}$ in $Y_{6}$, so $n_{1}=y_{0}m_{1}$, and we have coordinates $(y_{0},m_{1},r_{2})$. As in every other chart, there are three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e_{2}$. In this case, we have $e_{2}=y_{0}^{2}m_{1}r_{2}^{2}$, and so
$50=(y_{0}^{2}m_{1}r_{2}^{2})^{5}-5(y_{0}^{2}m_{1}r_{2}^{2})^{4}+25(y_{0}^{2}m_{1}r_{2}^{2})^{2}-25(y_{0}^{2}m_{1}r_{2}^{2})+$ 5 .
In this chart, $E_{1}=\{m_{1}=0\}$, $E_{2}=\{r_{2}=0\}$, $E_{4}$ does not appear, and $E_{5}=\{y_{0}=0\}$. The fixed point scheme $Y_{6}^{G}$ is defined by: $I(y_{0})=y_{0}^{2}m_{1}r_{2}(-1+O(y_{0}))$, $I(m_{1})=y_{0}m_{1}r_{2}(2m_{1}+O(y_{0}))$, and $I(r_{2})=y_{0}^{2}r_{2}^{2}(2-2m_{1}r_{2}+O(y_{0}))$. We know the fixed point set away from $E_{5}$, and so we read off that the fixed point scheme is generically the Cartier divisor $E_{2}+E_{5}$. (Since $E_{5}$ is fixed by $G$, we have finally made the fixed point set of codimension 1.) Let $e=y_{0}r_{2}$. The bad locus (where the scheme $Y_{6}^{G}$ is more than the Cartier divisor $E_{2}+E_{5}$), on $E_{5}$, is given by factoring out $e$ from the equations and setting $y_{0}=0$, so we get: $0=y_{0}$ and $0=2m_{1}^{2}$. So, as a set, the bad locus is the curve $\{0=y_{0}=m_{1}\}=E_{1}\cap E_{5}$. Theorem 2.2 does not seem to apply to this curve, and so $Y_{6}/G$ might not have toric singularities there; we will have to blow up one more time.
For now, look at the other open set, $\{m_{1}=1\}$ in $Y_{6}$. So $s_{0}=n_{1}m_{0}$, and we have coordinates $(m_{0},n_{1},r_{2})$. Here $E_{1}$ does not appear, $E_{2}=\{r_{2}=0\}$, $E_{4}=\{m_{0}=0\}$, and $E_{5}=\{n_{1}=0\}$. Also, $e_{2}=y_{0}m_{1}^{2}r_{2}^{2}$. The fixed point scheme $Y_{6}^{G}$ is defined by: $I(m_{0})=m_{0}n_{1}r_{2}(-2+O(n_{1}))$, $I(n_{1})=n_{1}^{2}r_{2}(1+O(n_{1}))$, and $I(r_{2})=m_{0}n_{1}^{2}r_{2}^{2}(2m_{0}-2r_{2}+O(n_{1}))$. Since we know the fixed point set outside $E_{5}$, we read off that the fixed point scheme is generically the Cartier divisor $E_{2}+E_{5}$. Let $e=n_{1}r_{2}$. The bad locus (where the scheme $Y_{6}^{G}$ is more than the Cartier divisor $E_{2}+E_{5}$) is the curve $\{0=m_{0}=n_{1}\}=E_{4}\cap E_{5}$. Fortunately, Theorem 2.2 applies, with $s=m_{0}$. We read off that $Y_{6}/G$ has singularity $\frac{1}{5}(-2,1,0)$ along the whole curve $E_{4}\cap E_{5}$, in this chart.
To finish describing $E_{5}\subset Y_{6}$, we have to work over the open set $\{u_{0}=1\}$ in $Y_{5}$, where the $G$-fixed curve $E_{1}\cap E_{4}$ in $Y_{5}$ meets $E_{3}$. Here $Y_{5}$ has coordinates $(q_{0},r_{1},u_{2})$, with $E_{1}=\{u_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, and $E_{4}=\{q_{0}=0\}$. Then $Y_{6}$ is the blow-up along the $G$-fixed curve $\{0=q_{0}=u_{2}\}=E_{1}\cap E_{4}$, so $Y_{6}$ has coordinates $(q_{0},r_{1},u_{2}),[t_{0},t_{2}]$. First take $\{t_{0}=1\}$ in $Y_{6}$, so $u_{2}=q_{0}u_{2}$ and we have coordinates $(q_{0},r_{1},t_{2})$. Here $E_{1}=\{t_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, $E_{4}$ does not appear, and $E_{5}=\{q_{0}=0\}$. Also, $e_{2}=q_{0}^{2}r_{1}t_{2}$. The fixed point scheme is defined by: $I(q_{0})=q_{0}^{2}r_{1}(-t_{2}+O(q_{0}))$, $I(r_{1})=q_{0}^{2}r_{1}^{2}(-2+2t_{2}+O(q_{0}))$, and $I(t_{2})=q_{0}r_{1}t_{2}(2t_{2}+O(q_{0}))$. So the fixed point scheme is generically $E_{3}+E_{5}$. Let $e=q_{0}r_{1}$. The bad locus (where the scheme $Y_{6}^{G}$ is more than the Cartier divisor $E_{3}+E_{5}$), in $E_{5}$, is given by $0=q_{0}$ and $0=2t_{2}^{2}$, so (as a set) it is the curve $\{0=q_{0}=t_{2}\}=E_{1}\cap E_{5}$, which we met in an earlier chart.
The other chart is $\{t_{2}=1\}$ in $Y_{6}$, so $q_{0}=u_{2}t_{0}$, and we have coordinates $(t_{0},r_{1},u_{2})$. Here $E_{1}$ does not appear, $E_{3}=\{r_{1}=0\}$, $E_{4}=\{t_{0}=0\}$, and $E_{5}=\{u_{2}=0\}$. Also, $e_{2}=q_{0}^{2}r_{1}t_{2}$. The fixed point scheme $Y_{6}^{G}$ is defined by: $I(t_{0})=t_{0}r_{1}u_{2}(-2+O(u_{2}))$, $I(r_{1})=t_{0}r_{1}^{2}u_{2}^{2}(2-2t_{0}+O(u_{2}))$, and $I(u_{2})=r_{1}u_{2}^{2}(1+O(u_{2}))$. So $Y_{6}^{G}$ is generically $E_{3}+E_{5}$. Let $e=r_{1}u_{2}$. The bad locus (where the scheme $Y_{6}^{G}$ is more than the Cartier divisor $E_{3}+E_{5}$), in $E_{5}$, is the curve $\{0=t_{0}=u_{2}\}=E_{4}\cap E_{5}$, which we met in an earlier chart. Theorem 2.2 applies, with $s=t_{0}$. Namely, $Y_{6}/G$ has singularity $\frac{1}{5}(-2,0,1)$ everywhere on the curve $E_{4}\cap E_{5}$ in this chart (including the origin, which did not appear in the earlier chart).
That completes our description of $Y_{6}$. In particular, the $G$-fixed locus has codimension 1 in $Y_{6}$, and $Y_{6}/G$ has toric singularities outside the image of the curve $E_{1}\cap E_{5}$. Let $Y_{7}$ be the blow-up of $Y_{6}$ along that curve. The exceptional divisor $E_{6}$ in $Y_{7}$ is a $\text{\bf P}^{1}$-bundle over $\text{\bf P}^{1}_{\text{\bf F}_{5}}$, and so we will cover $E_{6}$ with four affine charts. First, work over the open set $\{m_{0}=1\}$ in $Y_{6}$, where the bad curve $E_{1}\cap E_{5}$ meets $E_{2}$. Here $Y_{6}$ has coordinates $(y_{0},m_{1},r_{2})$, with $E_{1}=\{m_{1}=0\}$, $E_{2}=\{r_{2}=0\}$, and $E_{5}=\{y_{0}=0\}$. Since $Y_{7}$ is the blow-up along the curve $\{0=y_{0}=m_{1}\}=E_{1}\cap E_{5}$, $Y_{7}$ has coordinates $(y_{0},m_{1},r_{2}),[j_{0},j_{1}]$. (See Figure 6, which applies to the current example as well.)
First take $\{j_{0}=1\}$ in $Y_{7}$, so $m_{1}=y_{0}j_{1}$, and we have coordinates $(y_{0},j_{1},r_{2})$. As in every other chart, there are three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e_{2}$. In this case, we have $e_{2}=y_{0}^{3}j_{1}r_{2}^{2}$, and so
$50=(y_{0}^{3}j_{1}r_{2}^{2})^{5}-5(y_{0}^{3}j_{1}r_{2}^{2})^{4}+25(y_{0}^{3}j_{1}r_{2}^{2})^{2}-25(y_{0}^{3}j_{1}r_{2}^{2})+$ 5 .
In this chart, $E_{1}=\{j_{1}=0\}$, $E_{2}=\{r_{2}=0\}$, $E_{5}$ does not appear, and $E_{6}=\{y_{0}=0\}$. The fixed point scheme $Y_{7}^{G}$ is defined by: $I(y_{0})=y_{0}^{3}j_{1}r_{2}(-1+O(y_{0}))$, $I(j_{1})=y_{0}^{2}j_{1}r_{2}(1-2j_{1}+O(y_{0}))$, and $I(r_{2})=y_{0}^{2}r_{2}^{2}(2+O(y_{0}))$. So $Y_{7}^{G}$ is generically the Cartier divisor $E_{2}+2E_{6}$. Let $e=y_{0}^{2}r_{2}$. The bad locus (where the scheme $Y_{7}^{G}$ is more than the Cartier divisor $E_{2}+2E_{6}$), in $E_{6}$, is given by $0=y_{0}$, $0=j_{1}(1-2j_{1})$, and $0=r_{2}$, so it consists of the two points $(y_{0},j_{1},r_{2})$ equal to $(0,0,0)=E_{1}\cap E_{2}\cap E_{6}$ or $(0,-2,0)\in E_{2}\cap E_{6}$. At the first point, Theorem 2.2 applies, with $s=r_{2}$. We read off that $Y_{7}/G$ has singularity $(0,1,2)$ everywhere on the curve $E_{1}\cap E_{2}$ (including the origin, which did not appear when we saw $E_{1}\cap E_{2}$ in an earlier chart). To analyze the second point, change coordinates temporarily by $s_{1}=j_{1}+2$; then that point becomes the origin in coordinates $(y_{0},s_{1},r_{2})$. We have $I(y_{0})=y_{0}^{3}r_{2}(2-s_{1}+O(y_{0}))$, $I(s_{1})=I(j_{1})=y_{0}^{2}r_{2}(-s_{1}-2s_{1}^{2}+O(y_{0}))$, and $I(r_{2})=y_{0}^{2}r_{2}^{2}(2+O(y_{0}))$. Theorem 2.2 applies, with $s=r_{2}$. We read off that $Y_{7}/G$ has singularity $\frac{1}{5}(2,-1,2)$ at this point.
The other open set is $\{j_{1}=1\}$ in $Y_{7}$, so $y_{0}=m_{1}j_{0}$, and we have coordinates $(j_{0},m_{1},r_{2})$. Here $E_{1}$ does not appear, $E_{2}=\{r_{2}=0\}$, $E_{5}=\{j_{0}=0\}$, and $E_{6}=\{m_{1}=0\}$. Also, $e_{2}=j_{0}^{2}m_{1}^{3}r_{2}^{2}$. The fixed point scheme $Y_{7}^{G}$ is defined by: $I(j_{0})=j_{0}^{2}m_{1}^{2}r_{2}(2-j_{0}+O(m_{1}))$, $I(m_{1})=j_{0}m_{1}^{3}r_{2}(2+j_{0}+O(m_{1}))$, and $I(r_{2})=j_{0}^{2}m_{1}^{2}r_{2}^{2}(2+O(m_{1}))$. So $Y_{7}^{G}$ is generically $E_{2}+E_{5}+2E_{6}$. Let $e=j_{0}m_{1}^{2}r_{2}$. The bad locus (where the scheme $Y_{7}^{G}$ is more than the Cartier divisor $E_{2}+E_{5}+2E_{6}$), on $E_{6}$, is given by: $0=m_{1}$, $0=j_{0}(2-j_{0})$, and $0=j_{0}r_{2}$. So the bad locus is the union of the curve $\{0=j_{0}=m_{1}\}=E_{5}\cap E_{6}$ and the point $(j_{0},m_{1},r_{2})=(2,0,0)$ in $E_{2}\cap E_{6}$. That point is the one we analyzed in the previous chart. For the curve, Theorem 2.2 applies, using $s=j_{0}$. We read off that $Y_{7}/G$ has singularity $\frac{1}{5}(2,2,0)$ everywhere on the curve $E_{5}\cap E_{6}$, in this chart.
Last, work over the open set $\{t_{0}=1\}$ in $Y_{6}$, where the bad curve $E_{1}\cap E_{5}$ meets $E_{3}$. Here $Y_{6}$ has coordinates $(q_{0},r_{1},t_{2})$, with $E_{1}=\{t_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, and $E_{5}=\{q_{0}=0\}$. We obtain $Y_{7}$ by blowing up along the curve $\{0=q_{0}=t_{2}\}=E_{1}\cap E_{5}$, so $Y_{7}$ has coordinates $(q_{0},r_{1},t_{2}),[x_{0},x_{2}]$. First take $\{x_{0}=1\}$, so $t_{2}=q_{0}x_{2}$, and we have coordinates $(q_{0},r_{1},x_{2})$. Here $E_{1}=\{x_{2}=0\}$, $E_{3}=\{r_{1}=0\}$, $E_{5}$ does not appear, and $E_{6}=\{q_{0}=0\}$. Also, $e_{2}=q_{0}^{3}r_{1}x_{2}$. The fixed point scheme $Y_{7}^{G}$ is defined by: $I(q_{0})=q_{0}^{3}r_{1}(2-x_{2}+O(q_{0}))$, $I(r_{1})=q_{0}^{2}r_{1}^{2}(-2+O(q_{0}))$, and $I(x_{2})=q_{0}^{2}r_{1}x_{2}(1-2x_{2}+O(q_{0}))$. So $Y_{7}^{G}$ is generically $E_{3}+2E_{6}$. Let $e=q_{0}^{2}r_{1}$. The bad locus (where the scheme $Y_{7}^{G}$ is more than the Cartier divisor $E_{3}+2E_{6}$), in $E_{6}$, is given by: $0=q_{0}$, $0=r_{1}$, and $0=x_{2}(1-2x_{2})$, so it consists of the two points $(q_{0},r_{1},x_{2})$ equal to $(0,0,0)=E_{1}\cap E_{3}\cap E_{6}$ or $(0,0,-2)$ in $E_{3}\cap E_{6}$. Since $I(r_{1})=er_{1}(\text{unit})$, Theorem 2.2 applies at both points. At the origin, the theorem gives that $Y_{7}/G$ has singularity $\frac{1}{5}(2,-2,1)$, which is terminal. For the other point, change coordinates temporarily by $y_{2}=x_{2}+2$, so that the point becomes the origin in coordinates $(q_{0},r_{1},y_{2})$. We have $I(q_{0})=q_{0}^{3}r_{1}(-1-y_{2}+O(q_{0}))$, $I(r_{1})=q_{0}^{2}r_{1}^{2}(-2+O(q_{0}))$, and $I(y_{2})=I(x_{2})=q_{0}^{2}r_{1}(-y_{2}-2y_{2}^{2}+O(q_{0}))$. So Theorem 2.2 gives that $Y_{7}/G$ has singularity $\frac{1}{5}(-1,-2,-1)$ at this point.
The other chart is $\{x_{2}=1\}$ in $Y_{7}$, so $q_{0}=t_{2}x_{0}$, and we have coordinates $(x_{0},r_{1},t_{2})$. Here $E_{1}$ does not appear, $E_{3}=\{r_{1}=0\}$, $E_{5}=\{x_{0}=0\}$, and $E_{6}=\{t_{2}=0\}$. Also, $e_{2}=x_{0}^{2}r_{1}t_{2}^{3}$. The fixed point scheme $Y_{7}^{G}$ is defined by: $I(x_{0})=x_{0}^{2}r_{1}t_{2}^{2}(2-x_{0}+O(t_{2}))$, $I(r_{1})=x_{0}^{2}r_{1}^{2}t_{2}^{2}(-2+O(t_{2}))$, and $I(t_{2})=x_{0}r_{1}t_{2}^{3}(2-2x_{0}+O(t_{2}))$. So $Y_{7}^{G}$ is generically $E_{3}+E_{5}+2E_{6}$. Let $e=x_{0}r_{1}t_{2}^{2}$. The bad locus (where $Y_{7}^{G}$ is more than the Cartier divisor $E_{3}+E_{5}+2E_{6}$), in $E_{6}$, is given by: $0=t_{2}$, $0=x_{0}(2-x_{0})$, and $0=x_{0}r_{1}$, which is the union of the curve $\{0=x_{0}=t_{2}\}=E_{5}\cap E_{6}$ and the point $(x_{0},r_{1},t_{2})=(2,0,0)$ in $E_{3}\cap E_{6}$. We analyzed that point in the previous chart. Theorem 2.2 applies to the curve, using $s=x_{0}$. We read off that $Y_{7}/G$ has singularity $\frac{1}{5}(2,0,2)$ everywhere on the curve $E_{5}\cap E_{6}$ (including the origin, which did not appear in the earlier chart where we met this curve).
That completes our analysis of $Y_{7}$; we have shown that $Y_{7}/G$ has toric singularities. The sequence of blow-ups and the descriptions of the singularities of $Y_{7}/G$ are identical to those in the characteristic 5 example, Theorem 8.1. Given that, the proof that $Y_{0}/G$ is terminal is unchanged from that of the characteristic 5 example. Theorem 9.1 is proved. ∎
Remark 9.2.
As in Remark 7.2, I expect that there is also a 3-dimensional scheme $X$, flat over $\text{\bf Z}_{5}$, that is terminal and non-Cohen-Macaulay with $K_{X}$ Cartier. Namely, one should replace the $p$-adic integer ring $R=\text{\bf Z}_{5}[\zeta_{25}]^{\text{\bf Z}/4}$ in Theorem 9.1 by $S=\text{\bf Z}_{5}[\zeta_{25}]$, with the action of $\text{\bf Z}/5\subset(\text{\bf Z}/25)^{*}$.

### v2

@@
 It turns out that Kodaira vanishing fails even for Fano varieties, in every characteristic $p>0$ [29].
-This led to examples of klt, and even terminal, singularities that are not Cohen-Macaulay [26, 28].
+This led to examples of klt, and even terminal, singularities that are not Cohen-Macaulay [21, 29, 31].
 (Terminal singularities are the smallest class of singularities that can be allowed on minimal models.)
@@
 $\sigma(x_{1},x_{2},x_{3})=\bigg{(}\frac{1}{x_{1}},\;\frac{1}{x_{2}},\;\frac{1}{x_{3}}\bigg{)}.$
-Then $X$ is terminal but not Cohen-Macaulay [26, Theorem 5.1].
+Then $X$ is terminal but not Cohen-Macaulay [29, Theorem 5.1].
 This is the lowest possible dimension, because every terminal (or just normal) surface is Cohen-Macaulay.
 Cohen-Macaulayness and stronger properties such as $F$-regularity help to construct contractions of varieties.
-Partly for this reason, the MMP for 3-folds is known only in characteristics at least 5 [13].
+Partly for this reason, the MMP for 3-folds is known only in characteristics at least 5 [15, 8, 14].
 By Arvidsson–Bernasconi–Lacini, klt singularities in characteristic greater than 5 are Cohen-Macaulay, whereas there are klt singularities that are not Cohen-Macaulay in characteristics 2, 3, and 5 [4, 6, 10].
-In this paper, we construct terminal 3-fold singularities that are not Cohen-Macaulay in five new cases: mixed characteristic $(0,2)$, characteristic 3, mixed characteristic $(0,3)$, characteristic 5, and mixed characteristic $(0,5)$ (Theorems 0.1, 6.1, 7.1, 8.1, and 9.1.
+In this paper, we construct terminal 3-fold singularities that are not Cohen-Macaulay in five new cases: mixed characteristic $(0,2)$, characteristic 3, mixed characteristic $(0,3)$, characteristic 5, and mixed characteristic $(0,5)$ (Theorems 0.1, 6.1, 7.1, 8.1, and 9.1).
 This is optimal, in view of the result of Arvidsson–Bernasconi–Lacini.
@@
 Then the scheme $Y/G$ is terminal, not Cohen-Macaulay, of dimension 3, and flat over $\text{\bf Z}_{2}$.
-Note that an action of a $p$-group on a smooth variety in characteristic $p$ with an isolated fixed point is never formally isomorphic to a linear action, since a nonzero representation $V$ of a $p$-group $G$ in characteristic $p$ has nonzero $G$-fixed subspace.
+Note that an action of a $p$-group with an isolated fixed point on a positive-dimensional smooth variety in characteristic $p$ is never formally isomorphic to a linear action, because a nonzero representation $V$ of a $p$-group $G$ in characteristic $p$ has nonzero $G$-fixed subspace.
 In fact, there are continuous families of non-equivalent actions on smooth varieties in characteristic $p$, and likewise on regular schemes of mixed characteristic.
@@
 These examples should lead to other failures of vanishing theorems.
-In particular, by Baudin, Bernasconi, and Kawakami, these examples imply that Frobenius-stable Grauert-Riemenschneider vanishing fails in characteristic 2, 3, and 5 [5, Remark 1.2].
+In particular, by Baudin, Bernasconi, and Kawakami, these examples imply that Frobenius-stable Grauert-Riemenschneider vanishing fails for terminal 3-folds in characteristic 2, 3, or 5 [5, Remark 1.2].
 This work was supported by NSF grant DMS-2054553, Simons Foundation grant SFI-MPS-SFM-00005512, and the Charles Simonyi Endowment at the Institute for Advanced Study.
@@
 Another variant (modeled on big-O notation in analysis) is to write “$x=y+O(z)$” for “$x=y+(z)$”.
-We write $R\{x_{1},\ldots,x_{n}\}$ to mean the free module over a ring $R$ with basis elements $x_{1},\ldots,x_{n}$.
+We write $R\{x_{1},\ldots,x_{n}\}$ for the free module over a ring $R$ with basis elements $x_{1},\ldots,x_{n}$.
 For a closed point $P$ in a regular scheme $U$ with residue field $k_{U}$, we say that $f_{1},\ldots,f_{n}$ are coordinates for $U$ at $P$ (or a regular system of parameters) if $f_{1},\ldots,f_{n}$ are elements of the maximal ideal $\operatorname{\mathfrak{m}}$ of $O(U)$ (the regular functions vanishing at $P$) that map to a basis for the $k_{U}$-vector space $\operatorname{\mathfrak{m}}/\operatorname{\mathfrak{m}}^{2}$.
@@
 Write $I(f)=\sigma(f)-f$ for a function $f$ on a $G$-scheme.
-See section 5 for the definition of terminal singularities on general schemes, following [18, section 2.1].
+See section 5 for the definition of the canonical divisor and terminal singularities on general schemes, following [20, section 2.1].
 For a positive integer $r$, let $\mu_{r}$ be the group scheme (over any base scheme) of $r$th roots of unity.
-The Reid-Tai criterion is the following description of which cyclic quotient singularities are canonical or terminal [23, Theorem 4.11].
+The Reid-Tai criterion is the following description of which cyclic quotient singularities are canonical or terminal [25, Theorem 4.11].
 This is often stated over a field, but it works even in mixed characteristic for $\mu_{r}$-quotient singularities.
@@
 For such schemes, resolutions of singularities and the canonical divisor can be described in purely combinatorial terms.
-(The proof of Theorem 5.1 explains how to interpret the canonical divisor and terminal singularities in mixed characteristic.) In the following criterion, for $a$ an integer and $b$ a positive integer, consider $a\bmod b$ as an integer in the set $\{0,\ldots,b-1\}$.
+In the following criterion, for $a$ an integer and $b$ a positive integer, consider $a\bmod b$ as an integer in the set $\{0,\ldots,b-1\}$.
 Theorem 1.1.
@@
 Let $c_{1},\ldots,c_{n}$ be real numbers.
-Then the pair $(X,\sum c_{j}D_{j})$ is terminal near the image of $P$ if and only if $c_{j}<1$ for each $j$, $c_{j}+c_{k}<1$ for each $j\neq k$, and
+Then the pair $(X/\mu_{r},\sum c_{j}D_{j})$ is terminal near the image of $P$ if and only if $c_{j}<1$ for each $j$, $c_{j}+c_{k}<1$ for each $j\neq k$, and
 $\sum_{j=1}^{n}(1-c_{j})(ib_{j}\bmod r)>r$
@@
 Király and Lütkebohmert analyzed when a quotient scheme by $\text{\bf Z}/p$ is regular, the hard case being where the residue characteristic is $p$ [18].
-More generally, we now give a sufficient condition for a quotient by $\text{\bf Z}/p$ to be a toric singularity, or in particular when it is isomorphic to a $\mu_{p}$-quotient singularity.
+More generally, we now give a sufficient condition for a quotient by $\text{\bf Z}/p$ to be a toric singularity, or in particular to be a $\mu_{p}$-quotient singularity.
 (Actions of $\mu_{p}$ are far simpler than actions of $\text{\bf Z}/p$: they are linearizable near a fixed point, like finite group actions in characteristic zero, because $\mu_{p}$ is linearly reductive.
@@
 When it applies, the theorem can be described as a $\text{\bf Z}/p$$\mu_{p}$ switch.
-Here is Király and Lütkebohmert’s main result [16, Theorem 2]. (See Theorem 2.6 for a more detailed statement.)
+Here is Király and Lütkebohmert’s main result [18, Theorem 2]. (See Theorem 2.6 for a more detailed statement.)
 Theorem 2.1.
@@
 One might ask whether the assumption that $p\in e^{p-1}\operatorname{\mathfrak{m}}$ in Theorem 2.2 can be omitted.
-Fortunately, this assumption is automatic in characteristic $p$, and it will be easy to check in our mixed characteristic examples.
+Fortunately, this assumption is automatic in characteristic $p$, and it will be easy to check in our mixed-characteristic examples.
 Proof.
@@
 The fixed point scheme $U^{G}$ is defined as the closed subscheme of $U$ cut out by the ideal generated by $I(O(U))$.
-The following formulas hold for any action of $G$ on a commutative ring [16, Remark 3]:
+The following formulas hold for any action of $G$ on a commutative ring [18, Remark 3]:
 Lemma 2.3.
@@
 We are given coordinates $x_{1},\ldots,x_{n}$ for $U$ near $P$.
-Under our assumptions ($U$ regular and $k_{U/G}=k_{U}$), the ideal generated by $I(O_{U,P})$ in $O_{U,P}$ is generated by $I(x_{1}),\ldots,I(x_{n})$ [16, Proposition 6].
-That is, after shrinking $U$ around $P$ if necessary, the fixed point scheme $U_{G}$ is the closed subscheme defined by $I(x_{1}),\ldots,I(x_{n})$.
+Under our assumptions ($U$ regular and $k_{U/G}=k_{U}$), the ideal generated by $I(O_{U,P})$ in $O_{U,P}$ is generated by $I(x_{1}),\ldots,I(x_{n})$ [18, Proposition 6].
+That is, after shrinking $U$ around $P$ if necessary, the fixed point scheme $U^{G}$ is the closed subscheme defined by $I(x_{1}),\ldots,I(x_{n})$.
 In particular, if $I(x_{i})/e$ is a unit for some $i=1,\ldots,n$, then the fixed point scheme $U^{G}$ is the Cartier divisor $\{e=0\}$, and then Theorem 2.1 gives that $U/G$ is regular.
@@
 Proof.
-Since the ring $O_{U,P}$ is regular, it is a domain, and so $I(y)/e$ is well-defined for each element $y$ in $\operatorname{\mathfrak{m}}$.
+Since the local ring $O_{U,P}$ is regular, it is a domain, and so $I(y)/e$ is well-defined for each element $y$ in $\operatorname{\mathfrak{m}}$.
 Since $G$ fixes the point $P$, $G$ maps $\operatorname{\mathfrak{m}}$ into itself.
@@
 That is, we have a homomorphism from the ring of integers $\text{\bf Z}[\mu_{p}]$ in $\text{\bf Q}(\mu_{p})$ to $O(U)$, taking a primitive $p$th root of unity to $\zeta$.
-In $\text{\bf Z}[\mu_{p}]$, $p$ is $(1-\zeta)^{p-1}$ times a unit [19, section IV.1], and so the same is true in $O(U)$.
+In $\text{\bf Z}[\mu_{p}]$, $p$ is $(1-\zeta)^{p-1}$ times a unit [22, section IV.1], and so the same is true in $O(U)$.
 We are given that $p$ is in $e^{p-1}\operatorname{\mathfrak{m}}$.
@@
 So, as an element of $O_{U,P}$, $1-\zeta$ is in $e\operatorname{\mathfrak{m}}$.
-After shrinking $U$ if necessary, this gives the same conclusion in $O(U)$.
+After shrinking $U$ around $P$ if necessary, this gives the same conclusion in $O(U)$.
 ∎
@@
 Then $1/g^{p}=\sigma(f)/f=(\sigma(u)/u)^{p}$, and so $\zeta/g=\sigma(u)/u$ for some $p$th root of unity $\zeta$ in $O(U)$.
-Here $\zeta=1+(\zeta-1)=1+e\operatorname{\mathfrak{m}}$, by Lemma 2.5, and $1/g=1-e(1+\operatorname{\mathfrak{m}})$, so $\sigma(u)/u=1-e(1+\operatorname{\mathfrak{m}})$.
-Here $\sigma(u)/u=1+I(u)/u$, so $I(u)=ue(1+\operatorname{\mathfrak{m}})=e(\text{unit})$.
+Here $\zeta=1+(\zeta-1)=1+e\operatorname{\mathfrak{m}}$ by Lemma 2.5, and $1/g=1-e(1+\operatorname{\mathfrak{m}})$, so $\sigma(u)/u=1-e(1+\operatorname{\mathfrak{m}})$.
+Here $\sigma(u)/u=1+I(u)/u$, so $I(u)=-ue(1+\operatorname{\mathfrak{m}})=e(\text{unit})$.
 This contradicts our assumption that $I(O(U))$ is contained in $e\operatorname{\mathfrak{m}}$.
 So in fact $f$ is not a $p$th power.
-From there, we can show that the scheme $W$ is integral (after shrinking $U$ near $P$, if necessary).
+From there, we can show that the scheme $W$ is integral (after shrinking $U$ around $P$, if necessary).
 Namely, since $f$ is not a $p$th power in $O(U)$, $f$ is also not a $p$th power in the function field $k(U)$, and so $k(U)[f^{1/p}]$ is a degree-$p$ field extension of $k(U)$.
@@
 Since $\alpha^{-1}(V)$ is irreducible, it follows that $W$ is irreducible.
-Since $\alpha^{-1}(V)$ is reduced, $W$ is reduced in codimension 0; since $W$ is Cohen-Macaulay, it follows that $W$ is reduced [24, Tag 031R].
+Since $\alpha^{-1}(V)$ is reduced, $W$ is reduced in codimension 0; since $W$ is Cohen-Macaulay, it follows that $W$ is reduced [27, Tag 031R].
 Since $W$ is reduced and irreducible, it is integral.
@@
 Proof.
-Statement (1) is due to Király and Lütkebohmert for $B$ normal, but their proof works without change for $B$ a domain [16, Theorem 2 and Proposition 5].
+Statement (1) is due to Király and Lütkebohmert for $B$ normal, but their proof works without change for $B$ a domain [18, Theorem 2 and Proposition 5].
 They also prove statement (2).
-They prove statement (3) only when $B$ is regular.
+They prove statement (3) when $B$ is regular.
 We now extend the proof of (3) for $B$ only a domain.
@@
 As a result, even though $W$ is typically not normal, Theorem 2.6 gives that the morphism $W\to W/G=Q$ is finite and faithfully flat of degree $p$.
-It follows that $Q$ is noetherian [24, Tag 033E].
-(Beware that for a general noetherian scheme $X$ with an action of a finite group $G$, $X/G$ need not be noetherian, and the morphism $X\to X/G$ need not be finite [22, Proposition 0.10].
-These properties do hold if $X$ is of finite type over a noetherian ring $A$ and $G$ acts $A$-linearly [11, Theorem and Corollary 4].)
+It follows that $Q$ is noetherian [27, Tag 033E].
+(Beware that for a general noetherian scheme $X$ with an action of a finite group $G$, $X/G$ need not be noetherian, and the morphism $X\to X/G$ need not be finite [24, Proposition 0.10].
+These properties do hold if $X$ is of finite type over a noetherian ring $A$ and $G$ acts $A$-linearly [12, Theorem and Corollary 4].)
 The action of $\mu_{p}$ on the affine scheme $Q$ gives a grading of $O(Q)$ by $\text{\bf Z}/p$.
@@
 So the whole ring $O(Q)$ is finite over $O(Q/\mu_{p})$; that is, $Q\to Q/\mu_{p}$ is finite.
-Also, $O(Q/\mu_{p})$ is a pure subring (because it is a summand) of the noetherian ring $O(Q)$, so it is noetherian [14, Proposition 6.15]; that is, $Q/\mu_{p}=U/G$ is noetherian.
+Also, $O(Q/\mu_{p})$ is a pure subring (because it is a summand) of the noetherian ring $O(Q)$, so it is noetherian [16, Proposition 6.15]; that is, $Q/\mu_{p}=U/G$ is noetherian.
 Finally, the composition $W\to Q\to Q/\mu_{p}=U/G$ is finite, and $O(U)$ is a sub-$O(U/G)$-module of $O(W)$, so $O(U)$ is a finitely generated $O(U/G)$-module; that is, $U\to U/G$ is finite.
@@
 By Nakayama’s lemma, it follows that the extended ideal $(h_{1},\ldots,h_{r})$ is equal to the extended ideal $(h_{1},\ldots,h_{n})$ in $O_{W,P}$.
-Since $W\to Q$ is faithfully flat, extending and contracting an ideal in $O_{Q,P}$ gives the same ideal [20, Theorem 7.5].
+Since $W\to Q$ is faithfully flat, extending and contracting an ideal in $O_{Q,P}$ gives the same ideal [27, Tag 05CK].
 As a result, we have $(h_{1},\ldots,h_{n})=(h_{1},\ldots,h_{r})$ in $O_{Q,P}$.
@@
 Since the function $g_{i}w^{b_{i}}$ is $G$-invariant on $W$, we have $0=I(g_{i}w^{b_{i}})=\sigma(g_{i})I(w^{b_{i}})+I(g_{i})w^{b_{i}}=(g_{i}+I(g_{i}))I(w^{b_{i}})+I(g_{i})w^{b_{i}}$.
-We want to consider this equality modulo $e\operatorname{\mathfrak{m}}_{U}^{2}O(W)$, and so the term $I(g_{i})I(w^{b_{i}})$ can be omitted.
+When we consider this equality modulo $e\operatorname{\mathfrak{m}}_{U}^{2}O(W)$, the term $I(g_{i})I(w^{b_{i}})$ can be omitted.
 Namely, we have
@@
 There are several invariants we want to compute in this situation: the ramification index of the divisor $E$ in $Y$ (the positive integer $e$ such that $f^{*}F=eE$), and the coefficient $c$ of $E$ in the ramification divisor (meaning that $K_{Y}=f^{*}K_{Y/G}+cE$ near the generic point of $E$).
-Another name for $c$ is the valuation of the different $v_{E}(\mathcal{D}_{L/K})$, where $L$ is the function field $k(Y)$, $K=k(Y/G)$, and the valuation $v_{E}$ is the order of vanishing along $E$.
+Another name for $c$ is the valuation of the different $v_{E}(\mathcal{D}_{k(Y)/k(Y/G)})$, where the valuation $v_{E}$ on the function field $k(Y)$ is the order of vanishing along $E$.
 Here $ef=p$, where $f$ is the degree of the field extension $k(E)$ over $k(F)$.
@@
 In this case, $f^{*}F=E$, and $K_{Y}=f^{*}K_{Y/G}$ near the generic point of $E$.
-In the ramified case, the relative canonical class (that is, the valuation of the different) can be computed as follows.
-Define the Artin ramification number $i(E)$ of $Y$ over $Y/G$ along $E$ to be the coefficient of $E$ in the fixed point scheme $Y^{G}$.
-Equivalently, in terms of $I(a)=\sigma(a)-a$:
+Define the Artin ramification number $i(E)$ of $Y$ over $Y/G$ along $E$ to be the coefficient of $E$ in the fixed point scheme $Y^{G}$. Equivalently, in terms of $I(a)=\sigma(a)-a$:
 $i(E)=\min_{a\in O_{Y,E}}v_{E}(I(a)).$
-Then the valuation of the different is $(p-1)i(E)$ [27, section 2.1]. When $Y$ is regular, so that $Y/G$ is Q-factorial, we can equivalently say that
-$K_{Y}=f^{*}K_{Y/G}+(p-1)[Y^{G}],$
-where $[Y^{G}]$ denotes the Weil divisor associated to the fixed point scheme.
 In the ramified case, the field $k(E)$ is purely inseparable over $k(F)$, with degree $f$ equal to 1 or $p$.
@@
 It is clear that $v_{E}(I(t))\geq i(E)$.
-A very convenient criterion is: $Y\to Y/G$ is wildly ramified along $E$ if and only if equality holds [27, section 2.1].
+A very convenient criterion is: $Y\to Y/G$ is wildly ramified along $E$ if and only if equality holds [30, section 2.1].
 Otherwise, it is fiercely ramified.
+Furthermore, we can compute the relative canonical class (that is, the valuation of the different) as follows, correcting a typo in [30, section 2.1].
+Lemma 3.1.
+The valuation of the different is $(p-1)i(E)$. When $Y$ is regular, so that $Y/G$ is Q-factorial, we can equivalently say that
+$K_{Y}=f^{*}K_{Y/G}+(p-1)[Y^{G}],$
+where $[Y^{G}]$ denotes the Weil divisor associated to the fixed point scheme.
+Proof.
+If $G$ acts nontrivially on $E$ (the unramified case), then $i(E)=0$ and the statement is clear.
+So assume that $G$ fixes $E$.
+The local ring $O_{Y,E}$ is a discrete valuation ring.
+The algebra $O_{Y,E}$ is generated by one element $y$ as an algebra over $O_{Y/G,F}$; one can take $y$ to be a uniformizer in $O_{Y,E}$ if $E$ is wildly ramified, and an element of $O_{Y,E}$ whose restriction to $k(E)$ is not in $k(F)$ if $E$ is fiercely ramified [30, section 2.1].
+By Lemma 2.3, we have $i(E)=v_{E}(I(y))$.
+Let $u(X)$ be the minimal polynomial of $y$ over $k(Y/G)$.
+In this situation of a monogenic algebra extension, the different $\mathcal{D}_{k(Y)/k(Y/G)}$ is generated by $u^{\prime}(y)$ [26, III, Corollary 2 to Proposition 11].
+But $u(X)=\prod_{j=0}^{p-1}(X-\sigma^{j}(y))$.
+So
+$u^{\prime}(y)=\prod_{j=1}^{p-1}(y-\sigma^{j}(y)),$
+and hence
+$v_{E}(\mathcal{D}_{k(Y)/k(Y/G)})=v_{E}(u^{\prime}(y))=(p-1)i(E),$
+using that $i(E)$ is unchanged if we replace $\sigma$ by another generator of $G$. ∎
 4 Toric divisors
@@
 Assume moreover that the function $e$ is the greatest common divisor of the functions $I(x_{i})$ for $i=1,\ldots,n$ in the local ring $O_{U,P}$.
-Then, for each $y\in\operatorname{\mathfrak{m}}_{U}$ such that $\{y=0\}$ is an irreducible divisor $D\subset U$ and $I(y)\in(ey)$, $D$ is the pullback of a Weil divisor $F$ in $U/G$.
+Then, for each $y\in\operatorname{\mathfrak{m}}_{U}$ such that $\{y=0\}$ is an irreducible divisor $E\subset U$ and $I(y)\in(ey)$, $E$ is the pullback of a Weil divisor $F$ in $U/G$.
   2. (2)
@@
 That is, these functions are part of a toric coordinate system on $Q$.
-Finally, if $I(s)=es$ in Theorem 2.2 (as we can assume), then the $\mu_{p}$-weight of $h_{j}$ is equal to the eigenvalue of $\varphi$ on $y_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$, which is in $\text{\bf F}_{p}$.
+Finally, if $I(s)=es$ in Theorem 2.2 (as we can assume), then the $\mu_{p}$-weight of $h_{j}$ is equal to the eigenvalue of $\varphi(y):=I(y)/e$ on $y_{j}\in\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$, which is in $\text{\bf F}_{p}$.
 Proof.
-(1) Let $E$ be the irreducible divisor $\{y=0\}$ on $U$.
+(1) Let $E$ be the irreducible divisor $\{y=0\}$ in $U$.
 The assumptions imply that $E$ is mapped into itself by $G$, and that there is an $i\in\{1,\ldots,n\}$ such that $v_{E}(I(y))>v_{E}(I(x_{i}))$.
@@
 Indeed, $F$ is a Cartier divisor outside a codimension-2 subset of $U/G$, by normality of $U/G$.)
-(2) Since $Q$ is regular, the pullback of $F$ to $Q$ is a Cartier divisor, hence (after possibly shrinking $U$ and $Q$) of the form $\{t=0\}$ for some function $t\in\operatorname{\mathfrak{m}}_{Q}-\{0\}$.
+(2) Since $Q$ is regular, the pullback of $F$ to $Q$ is a Cartier divisor, hence (after shrinking $U$ and $Q$ around $P$) of the form $\{t=0\}$ for some function $t\in\operatorname{\mathfrak{m}}_{Q}-\{0\}$.
 Clearly the divisor $\{t=0\}$ is $\mu_{p}$-invariant.
@@
 (3) After multiplying $e$ by a unit, we can assume that $I(s)=es$, in the terminology of Theorem 2.2.
-The assumption that $I(y_{j})\in(ey_{j})$ for $j=1,\ldots,r$ implies that $y_{1},\ldots,y_{r}$ in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$ are eigenfunctions of the map $\varphi$.
+The assumption that $I(y_{j})\in(ey_{j})$ for $j=1,\ldots,r$ implies that $y_{1},\ldots,y_{r}$ in $\operatorname{\mathfrak{m}}_{U}/\operatorname{\mathfrak{m}}_{U}^{2}$ are eigenvectors of the map $\varphi$.
 By the proof of Theorem 2.2, the corresponding eigenvalues are in $\text{\bf F}_{p}\subset k_{U}$.
-For $j=1,\ldots,r$, we know from (1) that the divisor $\{y_{j}=0\}$ on $U$ is pulled back from a divisor $F$ on $U/G$.
-(Here $F$ is a Weil divisor, but it is a Cartier divisor outside a codimension-2 subset of $U/G$, since $U/G$ is normal.) By (2), $F$ pulls back to a divisor $\{h_{j}=0\}$ on $Q$ with $h_{j}$ a $\mu_{p}$-eigenfunction.
+For $j=1,\ldots,r$, we know from (1) that the divisor $\{y_{j}=0\}$ on $U$ is pulled back from a divisor $F_{j}$ on $U/G$.
+(Here $F_{j}$ is a Weil divisor, but it is a Cartier divisor outside a codimension-2 subset of $U/G$, since $U/G$ is normal.) By (2), $F_{j}$ pulls back to a divisor $\{h_{j}=0\}$ on $Q$ with $h_{j}$ a $\mu_{p}$-eigenfunction.
 It follows that the Cartier divisors $\{y_{j}=0\}$ on $U$ and $\{h_{j}=0\}$ on $Q$ have the same pullback to $W$; that is, $h_{j}=y_{j}(\text{unit})$ on $W$.
@@
 Proof.
-The scheme $Y$ is regular, being an open subset of the affine plane over the discrete valuation ring $\text{\bf Z}_{2}[\mu_{4}]=\text{\bf Z}_{2}[i]/(i^{2}+1)$.
+The scheme $Y$ is regular, being an open subset of the affine plane over the discrete valuation ring $\text{\bf Z}_{2}[\zeta_{4}]=\text{\bf Z}_{2}[i]/(i^{2}+1)$.
 Since $Y$ is a normal integral affine scheme of dimension 3, so is $Y/G$.
 The ring $O(Y/G)$ of regular functions is a torsion-free $\text{\bf Z}_{2}$-module, since it is a subring of the torsion-free $\text{\bf Z}_{2}$-module $O(Y)$; so $Y/G$ is flat over $\text{\bf Z}_{2}$.
-The fixed point scheme of $G$ on $Y$ is defined by: $I(x)=(1-x^{2})/x$, and $I(y)=(1-y^{2})/y$, and $I(i)=\sigma(i)-i=-2i$, hence by $\{x^{2}=1,y^{2}=1,2i=0\}$.
+The fixed point scheme of $G$ on $Y$ is defined by: $I(x)=(1-x^{2})/x$, $I(y)=(1-y^{2})/y$, and $I(i)=\sigma(i)-i=-2i$, hence by $\{x^{2}=1,y^{2}=1,2i=0\}$.
 Together with the equation $i^{2}=-1$ on $Y$, these equations imply set-theoretically that $2=0$, $x=1$, $y=1$, and $i=1$; so the fixed point set of $G$ is a single closed point $P$ in $Y$, with residue field $\text{\bf F}_{2}$.
@@
 So Cohen-Macaulayness of $Y/G$ would imply that $H^{1}(Y/G-P,O)=0$.
-Fogarty showed that for $G=\text{\bf Z}/p$ acting with an isolated fixed point on a normal scheme $W$ over $\text{\bf F}_{p}$ of dimension at least 3, $W/G$ is not Cohen-Macaulay [12, Proposition 4].
+Fogarty showed that for $G=\text{\bf Z}/p$ acting with an isolated fixed point on a normal scheme $W$ over $\text{\bf F}_{p}$ of dimension at least 3, $W/G$ is not Cohen-Macaulay [13, Proposition 4].
 When $W$ has mixed characteristic $(0,p)$, he needed $\dim(W)\geq 4$ to get the same conclusion.
@@
 Let us recall the definition.
-For a normal quasi-projective scheme $X$ over a regular base scheme $S$, Hartshorne defined the canonical sheaf $\omega_{X/S}$ [18, Definition 1.6].
+For a normal quasi-projective scheme $X$ over a regular base scheme $S$, Hartshorne defined the canonical sheaf $\omega_{X/S}$ [20, Definition 1.6].
 It is a reflexive sheaf of rank 1, or equivalently the sheaf associated to a Weil divisor.
@@
 with $a_{j}>0$ for every exceptional divisor $E_{j}$ of $\pi$.
-If $X$ has a resolution of singularities, terminality of $X$ is equivalent to positivity of the discrepancies $a_{j}$ on this one resolution [18, Corollary 2.12].
+If $X$ has a resolution of singularities, terminality of $X$ is equivalent to positivity of the discrepancies $a_{j}$ on this one resolution [20, Corollary 2.12].
 Let $Y_{0}=Y$.
 To prove that our example $Y_{0}/G$ is terminal, one approach would be to construct an explicit resolution of singularities.
-As with the analogous example in characteristic 2 [26, Theorem 5.1], this can be done by making $G$-equivariant blow-ups of $Y_{0}$ along regular closed subschemes.
+As with the analogous example in characteristic 2 [29, Theorem 5.1], this can be done by making $G$-equivariant blow-ups of $Y_{0}$ along regular closed subschemes.
 Namely, we can make $G$-equivariant blow-ups $Y_{2}\to Y_{1}\to Y_{0}$ such that $Y_{2}/G$ is regular.
-That resolution of $Y_{0}/G$ has dual complex a star, with one edge from a vertex $F_{0}$ to each of 7 other vertices $F_{1},\ldots,F_{7}$.
+That resolution of $Y_{0}/G$ has dual complex a star, with one edge from a vertex $F_{0}$ to each of seven other vertices $F_{1},\ldots,F_{7}$.
 However, we can simplify the proof that $Y_{0}/G$ is terminal by stopping with a partial resolution with toric singularities.
@@
 To focus on the fixed point scheme near $E_{0}$, we can say (more simply): $I(x_{0})=x_{0}^{2}(1+O(x_{0}))$, $I(y_{1})=x_{0}y_{1}(1+y_{1}+O(x_{0}))$, and $I(y_{2})=x_{0}y_{2}(1+y_{2}+O(x_{0}))$.
-We see that the fixed point scheme $Y_{1}^{G}$ is generically the Cartier divisor $E_{0}=\{x_{1}=0\}$.
+We see that the fixed point scheme $Y_{1}^{G}$ is generically the Cartier divisor $E_{0}=\{x_{0}=0\}$.
 The bad locus in $E_{0}$ (where that fails) is given by removing a factor of $x_{0}$ from the equations and setting $x_{0}=0$, so we have: $0=x_{0}$, $0=y_{1}(1+y_{1})$, and $0=y_{2}(1+y_{2})$.
-(Note that $x_{1}=0$ implies $2=0$, by the equation for $U_{0}$.) So the fixed point scheme $Y_{1}^{G}$ (in this chart) is $E_{0}$ as a Cartier divisor except at the 4 points $(x_{0},y_{1},y_{2})$ equal to $(0,0,0)$, $(0,1,0)$, $(0,0,1)$, or $(0,1,1)$.
+(Note that $x_{0}=0$ implies $2=0$, by the equation for $U_{0}$.) So the fixed point scheme $Y_{1}^{G}$ (in this chart) is $E_{0}$ as a Cartier divisor except at the 4 points $(x_{0},y_{1},y_{2})$ equal to $(0,0,0)$, $(0,1,0)$, $(0,0,1)$, or $(0,1,1)$.
 The open set $U_{1}=\{y_{1}=1\}$ works the same way, by the symmetry switching $x_{0}$ and $x_{1}$, hence also switching $y_{0}$ and $y_{1}$.
@@
 One further $G$-equivariant blow-up at each of these 7 points suffices to resolve $Y_{1}/G$, but the equations for these blow-ups are a bit messy.
-Instead, we will use Theorem 2.2 to show that the 7 singular points of $Y_{1}/G$ are all mixed characteristic analogs of the singularity $A^{3}/\mu_{2}$, the simplest terminal singularity whose canonical class is not Cartier.
+Instead, we will use Theorem 2.2 to show that the 7 singular points of $Y_{1}/G$ are all mixed-characteristic analogs of the singularity $A^{3}/\mu_{2}$, the simplest terminal singularity whose canonical class is not Cartier.
 More precisely, each of these 7 singular points is of the form $\frac{1}{2}(1,1,1)$, meaning that it can be written as $Q/\mu_{2}$ for some regular scheme $Q$ of dimension 3 (at an isolated fixed point of $\mu_{2}$).
@@
 Write $K_{Y_{0}}$ for the canonical sheaf $\omega_{Y_{0}/\text{\bf Z}_{2}}$.
-Since $Y_{0}$ is regular, $K_{Y_{0}}$ is a line bundle, described as follows [18, Definition 1.6].
+Since $Y_{0}$ is regular, $K_{Y_{0}}$ is a line bundle, described as follows [20, Definition 1.6].
 First, let $R=\text{\bf Z}_{2}[i]/(i^{2}+1)$.
@@
 It follows that the divisor class $K_{Y_{0}/G}$ is linearly equivalent to zero, in particular Cartier.
-Here $K_{Y_{0}/G}$ is the canonical sheaf in the sense of [18, Definition 1.6]; $Y_{0}/G$ is not Gorenstein, since (as we have shown) it is not Cohen-Macaulay.
+Here $K_{Y_{0}/G}$ is the canonical sheaf in the sense of [20, Definition 1.6]; $Y_{0}/G$ is not Gorenstein, since (as we have shown) it is not Cohen-Macaulay.
 Since $K_{X}$ is Cartier, we can write
@@
 The norm $N(x_{1})$ is a function on $U/G$ that defines a positive multiple of the divisor $F_{1}$, and so $F_{1}$ is Q-Cartier.
-Since the fixed point scheme $U^{G}$ is generically the Cartier divisor $E_{1}$, the ramification divisor of $f$ is $(p-1)E_{1}$, as mentioned above.
+Since the fixed point scheme $U^{G}$ is generically the Cartier divisor $E_{1}$, the ramification divisor of $f$ is $(p-1)E_{1}$ by Lemma 3.1.
 That is, $K_{U}=f^{*}K_{U/G}+(p-1)E_{1}$.
@@
 Let $Y_{0}=\text{\bf P}^{2}\times\text{\bf P}^{1}$ and $X=Y_{0}/G$.
-The only fixed point of $G$ on $Y_{0}$ is $A=([1,1,1],[0,1])$.
-So $X$ is normal of dimension 3, and $X$ is smooth over $k$ outside the image $P$ of $A$.
+The only fixed point of $G$ on $Y_{0}$ is $P=([1,1,1],[0,1])$.
+So $X$ is normal of dimension 3, and $X$ is smooth over $k$ outside the image of $P$, which we also call $P$.
 Also, $3K_{X}$ is Cartier.
-By Fogarty, since $A$ is an isolated fixed point of $G=\text{\bf Z}/p$ on a smooth 3-fold in characteristic $p$, $X$ is not Cohen-Macaulay at $P$ [12, Proposition 4].
+By Fogarty, since $P$ is an isolated fixed point of $G=\text{\bf Z}/p$ on a smooth 3-fold in characteristic $p$, $X$ is not Cohen-Macaulay at $P$ [13, Proposition 4].
 It remains to show that $X$ is terminal.
@@
 We wrote out the $G$-action on this chart above, with coordinates $(y_{0},x_{1},y_{2})$.
-We defined $Y_{2}$ as the blow-up of $Y_{1}$ along the $G$-fixed curve $\{0=y_{0}=x_{1}\}$.
+Here $E_{0}=\{x_{0}=0\}$.
+We defined $Y_{2}$ as the blow-up of $Y_{1}$ along the $G$-fixed curve $\{0=y_{0}=x_{1}\}$ in $E_{0}$.
 So, over this open subset of $Y_{1}$, $Y_{2}$ is an open subset of
@@
 So we have two new bad points, $(w_{0},x_{1},y_{2})$ equal to $(0,0,0)$ or $(0,0,1)$, for a total of five so far.
-To see all of $E_{1}$ in $Y_{2}$, we also need to work over $\{y_{2}=1\}\subset Y_{1}$, the chart containing the lower left vertex of $E_{1}$ in Figure 2.
+To see all of $E_{1}$ in $Y_{2}$, we also need to work over $\{y_{2}=1\}\subset Y_{1}$.
+The coordinates are $(y_{0},y_{1},x_{2})$, and $E_{0}=\{x_{2}=0\}$.
 The corresponding open subset of $Y_{2}$ is an open subset of
 $\{((y_{0},y_{1},x_{2}),[z_{0},z_{2}])\in A^{3}\times\text{\bf P}^{1}:y_{0}z_{2}=x_{2}z_{0}\}.$
-First consider the chart $\{z_{0}=1\}$ in $Y_{2}$. Then $(y_{0},y_{1},x_{2})=(y_{0},y_{1},y_{0}z_{2})$, and $G$ acts by
+First consider $\{z_{0}=1\}$ in $Y_{2}$, the chart containing the lower left vertex of $E_{1}$ in Figure 2. Then $(y_{0},y_{1},x_{2})=(y_{0},y_{1},y_{0}z_{2})$, and $G$ acts by
 $\tau(y_{0},y_{1},z_{2})=\bigg{(}\frac{y_{0}(1+y_{0}z_{2})}{1+y_{0}y_{1}z_{2}},\;\frac{(y_{0}+y_{1})(1+y_{0}z_{2})}{1+y_{0}y_{1}z_{2}},\;\frac{z_{2}(1+y_{0}y_{1}z_{2})}{(1+y_{0}z_{2})^{2}}\bigg{)}.$
@@
 The divisor class $\pi^{*}K_{Y_{0}/G}=K_{Y_{2}/G}-2F_{0}-F_{1}$ has some non-integer discrepancies, for example over the origin in the chart $\{z_{2}=1\}$ in $Y_{2}$ over $\{y_{2}=1\}$ in $Y_{1}$.
-As a result, $K_{Y_{0}/G}$ is not Cartier (as one can also check directly), in contrast to our examples in residue characteristic 2 ([26, Theorem 5.1] and Theorem 5.1).
+As a result, $K_{Y_{0}/G}$ is not Cartier (as one can also check directly), in contrast to our examples in residue characteristic 2 ([29, Theorem 5.1] and Theorem 5.1).
 I expect that there is also a 3-fold $X$ over $\text{\bf F}_{3}$ that is terminal and non-Cohen-Macaulay with $K_{X}$ Cartier.
-Namely, one should replace $\text{\bf P}^{1}$ in Theorem 6.1 by the “Harbater-Katz-Gabber” curve $C$$\{0=y^{p}-yz^{p-1}-x^{p-1}z\}$ in $\text{\bf P}^{2}_{\text{\bf F}_{p}}$ [8], here with $p=3$, which has a $\text{\bf Z}/p$-action by $\tau([x,y,z])=[x+z,y,z]$ that preserves a nonzero 1-form near the fixed point $[1,0,0]$.
+Namely, one should replace $\text{\bf P}^{1}$ in Theorem 6.1 by the “Harbater-Katz-Gabber” curve $C=\{0=y^{p}-yz^{p-1}-x^{p-1}z\}$ in $\text{\bf P}^{2}_{\text{\bf F}_{p}}$ [9], here with $p=3$, which has a $\text{\bf Z}/p$-action by $\tau([x,y,z])=[x+z,y,z]$ that preserves a nonzero 1-form near the fixed point $[1,0,0]$.
 For $p=3$, $C$ is a supersingular elliptic curve.
@@
 One can resolve the singularities of $X$ by performing $G$-equivariant blow-ups of $Y_{0}$.
-However, as in sections 5 and 6, we will shorten the proof by recognizing that, after two $G$-equivariant blow-ups $Y_{2}\to Y_{1}\to Y_{0}$, the singularities of $Y_{2}/G$ become toric (in Kato’s mixed characteristic sense), namely quotients of a regular scheme by $\mu_{3}$.
+However, as in sections 5 and 6, we will shorten the proof by recognizing that, after two $G$-equivariant blow-ups $Y_{2}\to Y_{1}\to Y_{0}$, the singularities of $Y_{2}/G$ become toric (in Kato’s mixed-characteristic sense), namely quotients of a regular scheme by $\mu_{3}$.
 That makes it easy to check that $Y_{0}/G$ is terminal, without having to continue making $G$-equivariant blow-ups.
@@
 $U_{0}=\{(x_{0},y_{1},y_{2})\in A^{3}_{\text{\bf Z}_{3}}:0=x_{0}^{3}y_{2}^{3}-3x_{0}^{2}y_{2}^{2}+3,\;1+x_{0}y_{1}\neq 0,\;1+x_{0}-x_{0}y_{1}\neq 0\}.$
-The exceptional divisor $E_{0}$ is $\{x_{0}=0\}$ (isomorphic to $A^{2}_{\text{\bf F}_{3}}$, in this open set). Here $G$ acts by
+The exceptional divisor $E_{0}$ is $\{x_{0}=0\}$. Here $G$ acts by
 $\tau(x_{0},y_{1},y_{2})=\bigg{(}\frac{x_{0}(1-3y_{1})}{1+x_{0}y_{1}},\;\frac{1-2y_{1}}{1-3y_{1}},\;\frac{y_{2}(1+2x_{0}y_{2}-x_{0}^{2}y_{2}^{2})(1+x_{0}y_{1})}{1-3y_{1}}\bigg{)}.$
@@
 So $Y_{1}^{G}$ is the line $\{y_{0}=x_{1}=0\}$, as a set.
-Finally, consider the chart $\{y_{2}=1\}$ in $Y_{1}$. Here $(x_{0},x_{1},e_{2})=(e_{2}y_{0},e_{2}y_{1},e_{2})$, and $G$ acts on the open set
+Finally, consider the chart $\{y_{2}=1\}$ in $Y_{1}$.
+We have coordinates $(y_{0},y_{1},e_{2})$, with $(x_{0},x_{1},e_{2})=(e_{2}y_{0},e_{2}y_{1},e_{2})$, and $E_{0}=\{e_{2}=0\}$.
+Here $G$ acts on the open set
 $U_{2}=\{(y_{0},y_{1},e_{2})\in A^{3}_{\text{\bf Z}_{3}}:0=3-3e_{2}^{2}+e_{2}^{3},\;1+e_{2}y_{0}\neq 0,\;1+e_{2}y_{0}-e_{2}y_{1}\neq 0\}.$
@@
 Then $(S_{5}\times\text{\bf P}^{1})/G$ is terminal, not Cohen-Macaulay, and of dimension 3 over $\text{\bf F}_{5}$.
-We define the quintic del Pezzo surface (over any field) as the moduli space $\overline{M_{0,5}}$ of 5-pointed curves of genus 0.
+We define the quintic del Pezzo surface (over any field) as the moduli space $\overline{M_{0,5}}$ of 5-pointed stable curves of genus 0.
 That makes it clear that the symmetric group $\Sigma_{5}$ acts on this surface.
@@
 Also, $5K_{X}$ is Cartier.
-By Fogarty, since $A$ is an isolated fixed point of $G=\text{\bf Z}/p$ on a smooth 3-fold in characteristic $p$, $X$ is not Cohen-Macaulay at $P$ [12, Proposition 4].
+By Fogarty, since $P$ is an isolated fixed point of $G=\text{\bf Z}/p$ on a smooth 3-fold in characteristic $p$, $X$ is not Cohen-Macaulay at $P$ [13, Proposition 4].
 It remains to show that $X$ is terminal.
 This example is more complicated than those in characteristics 2 and 3, and it may be impossible to resolve the singularities of $X$ by performing $G$-equivariant blow-ups of $Y_{0}$.
-(Indeed, in the simpler situation of actions of $G=\text{\bf Z}/p$ in characteristic zero, one cannot always resolve the singularities of a quotient $Y/G$ via $G$-equivariant blow-ups of $Y$ when $p\geq 5$ [17, Claim 2.29.2].) Fortunately, as in earlier sections, we can reach toric singularities after some $G$-equivariant blow-ups.
+(Indeed, in the simpler situation of actions of $G=\text{\bf Z}/p$ in characteristic zero, one cannot always resolve the singularities of a quotient $Y/G$ via $G$-equivariant blow-ups of $Y$ when $p\geq 5$ [19, Claim 2.29.2].) Fortunately, as in earlier sections, we can reach toric singularities after some $G$-equivariant blow-ups.
 It will then be easy to check that $Y_{0}/G$ is terminal.
@@
 $\{((s_{0},s_{1},s_{2}),[y_{0},y_{1},y_{2}])\in U\times\text{\bf P}^{2}:s_{0}y_{1}=s_{1}y_{0},\;s_{0}y_{2}=s_{2}y_{0},s_{1}y_{2}=s_{2}y_{1}\}.$
-It turns out that the fixed point set of $G$ on $Y_{1}$ is a curve isomorphic to $\text{\bf P}^{1}$.
+The exceptional divisor $E_{0}$ is isomorphic to $\text{\bf P}^{2}$.
+It turns out that the fixed point set of $G$ on $Y_{1}$ is a curve isomorphic to $\text{\bf P}^{1}$ in $E_{0}$.
 To check that, first work in the open subset $\{y_{0}=1\}$ in $Y_{1}$.
@@
 $\tau(s_{0},y_{1},y_{2})=\bigg{(}\frac{s_{0}(1+s_{0}-y_{1}+s_{0}y_{1})}{(1-2s_{0})(1-s_{0}-s_{0}y_{1})},\;\frac{-2s_{0}+y_{1}-2s_{0}y_{1}}{1+s_{0}-y_{1}+s_{0}y_{1}},\;\\ \frac{y_{2}(1-2s_{0})(1-s_{0}-s_{0}y_{1})}{(1+s_{0}y_{2})(1+s_{0}-y_{1}+s_{0}y_{1})}\bigg{)}.$
-The exceptional divisor $E_{0}$ is $\{s_{0}=0\}$, in this chart.
+Here $E_{0}=\{s_{0}=0\}$.
 The fixed point scheme $Y_{1}^{G}$ is defined by the vanishing of: $I(s_{0})=s_{0}(-y_{1}+O(s_{0}))$, $I(y_{1})=(y_{1}^{2}+O(s_{0}))/(1-y_{1}+O(s_{0}))$, and $I(y_{2})=y_{2}(y_{1}+O(s_{0}))/(1-y_{1}+O(s_{0}))$.
@@
 We can write the action of $G$ in these coordinates (for example using Magma).
-We find that the fixed point scheme $Y_{1}^{G}$ is defined by the vanishing of: $I(y_{0})=-1+O(s_{1})$, $I(s_{1})=s_{1}^{2}(1+y_{0}-2y_{0}^{2})$, and $I(y_{2})=s_{1}y_{2}(-1-y_{0}-y_{2}+2y_{0}^{2})$.
+We find that the fixed point scheme $Y_{1}^{G}$ is defined by: $I(y_{0})=-1+O(s_{1})$, $I(s_{1})=s_{1}^{2}(1+y_{0}-2y_{0}^{2})$, and $I(y_{2})=s_{1}y_{2}(-1-y_{0}-y_{2}+2y_{0}^{2})$.
 Since $Y_{1}^{G}$ is contained (as a set) in $E_{0}$, the first equation shows that $Y_{1}^{G}$ is empty, in this chart.
@@
 Theorem 2.2 applies, with $s=e=w_{0}$, and we read off that $Y_{3}/G$ has singularity $\frac{1}{5}(1,-1,-2)$ at this point.
-That is terminal by the Reid-Tai criterion (Theorem 1.1, which will be convenient.
+That is terminal, by the Reid-Tai criterion (Theorem 1.1).
 The last chart we need to consider in $Y_{3}$ is the other open set $\{v_{1}=1\}$ over the open set above in $Y_{2}$, $\{w_{1}=1\}\subset Y_{2}$ over $\{y_{0}=1\}\subset Y_{1}$.
@@
 Theorem 9.1.
-Let the group $G=\text{\bf Z}/5$ with generator $\tau$ act on the quintic del Pezzo surface $S_{5}$ over $\text{\bf Z}_{5}$ by an embedding of $G$ into the symmetric group $\Sigma_{5}=\operatorname{Aut}(S_{5})$.
+Let the group $G=\text{\bf Z}/5$ act on the quintic del Pezzo surface $S_{5}$ over $\text{\bf Z}_{5}$ by an embedding of $G$ into the symmetric group $\Sigma_{5}=\operatorname{Aut}(S_{5})$.
 Let $R=\text{\bf Z}_{5}[e]/(e^{5}-5e^{4}+25e^{2}-25e+5)$, which is the ring of integers in a Galois extension of $\text{\bf Q}_{5}$ with group $G=\text{\bf Z}/5$.
@@
 Then the scheme $(S_{5})_{R}/G$ is terminal, not Cohen-Macaulay, of dimension 3, and flat over $\text{\bf Z}_{5}$.
-We define the quintic del Pezzo surface $S_{5}$ (over any commutative ring) as the moduli space $\overline{M_{0,5}}$ of 5-pointed curves of genus 0.
+We define the quintic del Pezzo surface $S_{5}$ (over any commutative ring) as the moduli space $\overline{M_{0,5}}$ of 5-pointed stable curves of genus 0.
 That makes it clear that the symmetric group $\Sigma_{5}$ acts on $S_{5}$.
@@
 First take $\{w_{0}=1\}$, so $y_{1}=s_{0}w_{1}$, and we have coordinates $(s_{0},w_{1},y_{2})$.
-As in every other chart, there are three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e$.
+As in every other chart, there are three variables over $\text{\bf Z}_{5}$, but this is a regular scheme of dimension 3 because of the equation satisfied by $e_{2}$.
 In this case, we have $e_{2}=s_{0}y_{2}$, and so
@@
 Theorem 2.2 applies, with $s=e=w_{0}$, and we read off that $Y_{3}/G$ has singularity $\frac{1}{5}(1,-1,-2)$ at this point.
-That is terminal by the Reid-Tai criterion (Theorem 1.1, which will be convenient.
+That is terminal, by the Reid-Tai criterion (Theorem 1.1).
 The last chart we need to consider in $Y_{3}$ is the other open set $\{v_{1}=1\}$ over the open set above in $Y_{2}$, $\{w_{1}=1\}\subset Y_{2}$ over $\{y_{0}=1\}\subset Y_{1}$.